A Bibliography of Publications about TEX

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254

E-mail: beebe@math.utah.edu, beebe@acm.org,
beebe@computer.org (Internet)
WWW URL: https://www.math.utah.edu/~beebe/

30 September 2025 Version 3.101

Abstract

This bibliography records books and articles on TEX, , WEB, fonts, typography, indexing, and software related to these topics. ISBN and ISSN numbers are given, where available. A few troff-related items are included because of their strong overlap with TEX issues.

Title word cross-reference

3 [Fin03, Fin04, Gon04]. e [Mao94]. GL(2) × GL(3) [Wu08]. GL(6) [Wu08]. ∫ [Ver00]. μoνo2πoλυ [Lik04]. Ω [CO04, PHSB04]. π [Bec77]. SO(3) [SS90]. T^{3} [Mil86b].

-dimensional [Fin03, Fin04]. -Wegweiser [Det97].

/L^{A}TEX [LS96]. /The [IEE92].

’05 [Kin05].

1 [Píš04]. 14 [AH05]. 16 [Hof82]. 1950 [Bri50]. 1976 [Bri76]. 1989 [Bri89]. 1992 [Zla92]. 1992/93 [Sch93]. 1VIPS [CYW^{+}03].

2.5 [Ano97]. 2010 [Bri10]. 22nd [Mil86a]. 23rd [Mil85]. 24th [Mil84, Mil86a]. 25th [Mil85, SBH^{+}04, WH86]. 26th [Mil84]. 2D [Kot23]. 2nd [BMJ88].

3.0 [Har93b]. 30th [WH86]. 3D [Kot23]. 3DLDF [Fin03, Fin04].

4TEX [DFKvdV94, DF99]. 4th [HAB98, HY14].

5.25in [Bec91b]. 594 [Eij12].

690 [ISO87].

7-10 [VC92]. 7th [HAB98, Zla92].

’88 [BMJ88]. 8th [Nad90].

’92 [VC92, Zla92]. ’95 [Lev95]. ’98 [HAB98].

= [KD15].

ABC [DE89]. ABC-XYZapf [DE89]. Abecedarium [Fir93]. Abstracting [Har02a]. abstraction [Wan96]. Academic [Ano15, KN14, Wal82a]. accents [FS92]. access [Day83]. accessibility [Mat05]. Accessible [LV23]. achieve [MR92c]. ACM [Kin05, ACM81]. Actes [And83a]. Active [Lin05, Lin06]. Ada [San89]. adaptation [BB89]. Adapting [RS85]. Adaptive [Atk04, JLS03a, JLS^{+}03b, JLS^{+}04, LA05]. Address [Har02a]. Administrative [Tal15]. Adobe [Ado90, Har93a]. Advanced [Her81, KD93, KD95, KD99, Sal95, Wil04]. Advances [Grä93b, Grä94a, Grä94b, Pie91]. Aesthetic [HNJ^{+}04]. affect [MBHL91]. AFIPS [AFI76]. African [WH86]. Age [Whi88]. Agents [Sch88a]. Ahead [BMJ88]. aid [Gar74a]. akapicie [Ryć93]. Aleph [SW92]. Algebraic [Gut04, Lev95]. Algorithm [BK91, BBS01, CYW^{+}03, HM90b, SKS97, LS85]. Algorithmes [Hég84]. Algorithmic [KD15]. Algorithms [Lin06, Fis91]. alignments [Bei00]. Alltag [DL94]. Alphabet [Bee04, DE89, Fir93, Knu79c, Zap07a]. Alphabetgeschichten [Zap07b]. Alphabetic [Dru95, Gar68, RS85]. Alphabets [Tsc95c]. Alternative [Har93a]. always [Loz13]. American [Ano84]. amont [FPR84]. Ampdes [SvR92]. amplifiers [SvR92]. AMS [Ano90b, Jac94, KZ89]. Analysis [DM91]. Analytical [And83b]. Anatomy [Law90]. Ancient [Fil04]. angular [SS90]. Animation [BK91, Hay77, HS04]. Animations [HS04]. Anleitung [WKS^{+}93a, WKS^{+}93b]. annealing [Lag95]. Annex [AH05]. Annotated [SS91]. Annual [Ano91b, SBH^{+}04]. ans [Kna97]. Anthology [McL95]. Anwender [Sch91b]. APA [Vip96]. APE [San89]. APL [Dic87, GSMN93, Hal87, Hoh88, How86]. APL2TEX [GSMN93]. APL87 [Hal87]. appearance [Kna92]. application [Sie92b]. Applications [BU89, Cla90, GR82, HL87, Vul92, CM12, Doo02, Lin06, Rah87a]. appreciated [Kra96]. approach [Fei88, Won86a, Won90, Gho83, GB83]. April [HAB98, VC92, vV86, vV88]. Arabic [Bec87, BEL04, ELS04, Har92b, Lag92, Mac83]. ArabTEX [Lag92]. arbeiten [ST05]. Arbeitsheft [WKS^{+}93a]. arbitrary [dSdOM^{+}05]. Architects [Tsc95c]. Archiv [LS96]. Archive [BP04, KW90]. archived [PB03]. Area [DS83]. Aren’t [Sch88a]. arithmetic [BBS01]. arrows [Ros92a]. Art [Abi88, Bai70, Bai90, SP92, Tsc95c, TW87]. Article [Hof82, vdLL89]. Articles [NCNS82, Zap87, vH89]. Artifacts [DM91]. artistic [HAB98]. Artists [Tsc95c]. Arts [Arn72]. Asian [WH86, Kor07]. ask [Loz13]. Aspects [Her93, vdVM88]. Assistants [Swa86]. Assisté [Ros84]. Associates [Man87b]. Association [Ano84]. Astronomical [Hec92]. Astronomy [Hec92]. Asymptote [HBP04]. aufgabenorientierte [WKS^{+}93a, WKS^{+}93b]. Aufsätze [Tsc87]. August [Cla90, HY14, ICC88, SBH^{+}04, WH86]. Ausgewählte [Tsc87]. auteurs [Féd89]. Author [Bry88, Doo89, Sto92]. Authors [Ano87, Kel85, Swa86, Gar74a]. Automated [Lag95, MR92c, Sch04, HNJ^{+}04, LFN04]. Automatic [JMPS96, JMPS97, MMH07, Pla81]. Automating [CH87, Fei88, MBO93]. Automation [For89, IEE92, YKT90]. available [KW90]. avez [Loz13].

B [WKS^{+}93b]. Babbage [Bux88]. Back [Cla87]. balance [LFN04]. Band [Kop94, Kop95, Kop97, Sch94]. Barriers [KD15]. Based [BU89, DS83, Lik04, AP83, CYW^{+}03, Di 07, Eij12, Fei88, HBP04, JLS^{+}03b, Wit86]. baseline [Kan91]. Basic [Gho84, Col62, Krö90]. Basics [Cla87, vB92b]. Basque [PLA04]. be [Sha01]. Beam [Cer85]. beautiful [Mat19]. Beauty [FvGGM90]. Beginner [Sam83, SL91, Sno92, Kot11]. Beginners [KD93, KD95, KD99]. Begleiter [GMS94b]. beim [Bin93]. Being [Kra96]. Belgium [IEE92]. Benefits [MBO93, Krö90]. berbère [Har94]. Bertinoro [ABD03]. Best [Tsc95c]. better [PK82]. between [Ves92]. Beyond [Huf04, Wil04]. Bibliographic [Ame78, Hun94, ISO87, AKWC82, Bri10]. Bibliographical [Bri50, Zap87, Bri76]. bibliographies [Vos12]. Bibliography [SS91, And83b, HM89]. Binary [Gri90]. Biochemists [Fuj93]. Biography [Lev19]. Birthday [FvGGM90]. bit [Loz13, Vin92]. Bitmap [Cas93, PB03]. Bitmap-free [Cas93]. Bluebook [Ano91a]. bolts [Ame01]. Book [Ano90a, Bai70, Bai90, CH01, Ker81a, PH92, Pop92, Ric78, SL91, Tsc91b, Tsc95c, Wil83, Woh98]. Book-Like [Woh98]. Bookmaking [Lee79]. Books [Ben91, SB90, Col62, Cro80, Dem89, Hur78, Mul94, Mul05]. Boston [Mil87]. Box [Bil89]. Braille [AJ88]. Branch [BBW18, BBW19]. Breaking [KP81, AH05, Fin00, Hol06, Grä09a]. breaks [PK82]. Brief [Z^{+}94]. Bristol [Kin05]. Brush [Hob86]. Brussels [IEE92]. BS [Bri50, Bri76, Bri89, Bri10]. Buch [Abd92, Abd93, Abd96]. Buches [Tsc87]. Buchsatz [Sch89, Sch91b]. building [RKM88]. built [Gun85]. built-in [Gun85]. Bulletin [BP04]. Business [FvGGM90].

Busy [Kwo91]. Buyers [Lab84]. Byte [Pla06].

C [GP89]. C-like [GP89]. Caching [FK82, FK85]. CaesarCM [Sie92a]. Cahiers [Cah ]. Calligraphy [Zap87]. Calling [Knu89b]. Cambridge [But81, AKWC82, UCBB83]. can [MR92c]. Canada [ICC88, Lev95]. Cantabria [Gut04]. Capabilities [Mil90]. caractères [Har04, Hég84]. Carolina [BS83]. Cartesian [SS90]. Cartography [Her69]. Case [BP04, Gau07, PLA04]. Cassowary [BBS01]. ce [Loz13]. Centre [AKWC82]. Chance [Wis01]. Changing [Kna92]. Character [Her87, Nai91, Ada91, Smi90, Tun80]. Characters [sT81, TM88, GH82, Li85, RS85]. Charles [Bux88]. charts [Kot23]. cheap [Lav92a]. Chel [MF86a]. CHEM [BJK87]. Chemical [BJK87, BU89, Hic89]. Chemistry [RSV16]. Chemists [Fuj93]. CHI [Knu78b]. Chicago [Ano82, Ano87, Ano03, Ano10, Ano17]. Chinese [Ano94, ICC88, Li85, CCL88, GH82, HG83, Liu13, Tun80, sT81, TM88]. Chiwriter [Cor86]. Choix [Féd89]. Choosing [Har02b, PK82]. Chronicle [Zap07a]. Chronik [Zap07b]. cientificos [Val97]. circuits [Sim87]. citations [Bri10, Gar74a, Sha01]. cited [Gar68]. City [AFI76]. CJK [CO04]. Class [CSR10a]. Classic [MZH22]. CMS [AMSV85]. Co [Bie90, Ryć93]. codages [Har04]. Code [Féd89, Sch04]. Collection [BU89, NCNS82, Atk04]. Colophon [Adoxx]. column [MMH07]. Com [Lia83, Lia84]. Com-pu-ter [Lia83, Lia84]. Command [BB85, Voß12]. commands [Vos12]. Comment [Loz13]. Comments [Hof82]. Commercial [Tsc95c, Sey82]. Common [Cra87a]. Commonplace [CH01]. Communication [BP16, Gin94, Den92]. Como [Luc85b]. comp.text.tex [com ].

Companion [GMS94a, GRM97, GMR^{+}08, MGB^{+}04, MFW23, MFB^{+}23, GR99]. Comparison [Ano15, KN14, Píš04, Kon01]. Compatibles [otB86]. Competition [Gau07]. Compiler [Rei80, Gel95]. Complete [Tal12, Zap87]. Completely [Ale92]. Complex [BKKW98, Sou92, Woh98, BKKW03]. Composicion [Val97]. Composing [Ano88, LVC89]. Composition [Ale92, BE92, Rub88, Sou92, IO89, Gho83]. Compositors [Har91]. composizione [Bec91b]. compound [Mag11]. Comprehensive [Pak05]. Compress [GH82]. Computation [Lev95, Gut04]. Computer [AK73, BU89, CSR10a, Eij12, Haf97, Hah91, Hah93, Her72, Her81, ICC88, KSW03, KD04, LK76, AFI76, Neg88, Rub88, Spi80, WH76, AKWC82, Bia92, FN87, GB82, KR85, Knu91, Ueh90, vdVM88, Wit85, Knu80b, Knu86c]. Computerized [AJ88]. Computers [Ano91b, BU89, BS83, Fuj93, Bia92, BK85]. Computing [Nad90, Ano94]. Con [Bec91b, Val97]. Concept [Knu81, Knu82b, Knu85]. concepts [FSS82a, FSS82b]. Concise [Kol24]. Concisely [Joh92]. Concurrency [Mal19]. Conference [Ano91b, ABD03, BU89, BS83, Cla90, Dés86, For89, Fur90, Hal87, HAB98, HY14, ICC88, IEE92, Luc85b, Mil84, Mil85, Mil86a, Mil87, AFI76, SBH^{+}04, VC92, vdVM88, Zla92, vV86, vV88, Ano94]. Congress [Pie91, WH86]. connaît [Loz13]. Constrained [LA05]. Constraint [Mat05, BTM^{+}01, BBS01]. Constraints [Her87]. Constructing [Wal87b]. Contemporary [Zap87]. contenente [Bec91b]. Content [ISO87, CM12, JLS03a, POM12]. Context [EtPR94]. Context-free [EtPR94]. Contribution [WH76]. contributions [DE89]. Controller [Cer85]. Conversion [Lei13, Lik04, LL94]. Cookbook [Ado85b, Kot15]. Coordinates [WH76]. Coping [Smi89]. Copy [But81, Swa86]. Copyrights [Sto92]. Cork [Sie92b]. Correction [Ano15]. Cosmetic [OC90]. Cost [Cer85, FS92, Gun85, Lav92b]. coupling [SS90]. CPU [Doo90]. Craft [Mat19, MR92c, Wil83]. Creating [HB96, Píš04, Vos12]. crosswords [vdL92c]. Crystallographic [AKWC82]. CS [Eij12]. CTAN [LS96]. CTAN-3 [LS96]. Current [Ueh90]. customized [POM12]. CWEB [KL93]. Czech [Olš97]. Czechoslovakia [Nad90, Zla92]. czyli [Paw93].

D [Gon04]. Dallas [Hal87]. dans [Bor90]. dansk [Lar89]. DANTE [LS96, Sch94, Sch93]. DANTE-TEX-Stammtisch [Sch93]. daring [Loz13]. Darstellung [Kar92a]. Data [AKWC82, Ben88, CSR10a, Lin06, MBO93]. database [Myh90, SČS92]. Databases [Har02a]. day [McL95]. Decade [Jac17]. Decimal [Gri90]. decisions [Ros92a]. Decorative [MF87a]. Dedicated [KSW03]. defects [Ada91]. définitive [Sie92b]. Delivery [CDMR85]. demander [Loz13]. Descripción [PSH93]. Description [Kar94a, HW97, GB83]. Deseret [Bee04]. Design [Ben86a, BH86, Col91, Dai67, DE89, EPO , Eng87, HC88, IEE92, Knu79c, Lee79, Nel78, Nel91, Rei88, Ric78, Rub88, Rug83, Tsc95c, Whi88, Wil83, Zap87, Gho84, Hur78, MR92b, MR92a, Ros92a, RKM88, Rug87, Sou85a, Ver00, Gho83, GB83, Tun80]. Design/Production/Editing [Lee79]. Designer [Har92c, RKM88]. designers [Tsc95b]. Designing [Cra71, Sou85b]. Desktop [BBW18, BBW19, Col91, Cra87b, FN87, Hec92, Lun92, YKT90, Bla90, Cra87a]. dessin [Hob01]. Detecting [DM91]. detection [Kan91]. Developers [Gau07]. Development [KN14, Ano15].

Developments [Por92, Zap87, Zap07a]. Device [Utt85, BB89]. device-independent [BB89]. Devices [Hug88, Gun85]. dformat [Ben88]. diacritics [FS92]. diagram [Ros92a]. Diagrams [BJK87, Ker81c, Lev90, Ros92b, Kot23, Ohl95]. Dictionary [AA04, Luc85a]. difference [Pap91]. Different [Ale92, Kon01]. Diffusion [Knu87]. Digital [AH89, BD83, Bry96, DM89, Ika83, Kar87, Kar92a, Kar94a, Knu87, Knu99, LS04, McC20, Rub88, STS03, SBH^{+}04, Uli87, And92, Di 07, GC88, HAB98]. Digitale [Kar86, Kar92a]. Digitized [Hob86]. dijitaru [Kar89]. Dijkstra [FvGGM90]. dimensional [Fin03, Fin04]. dimensions [AKWC82]. Direction [Jac94]. directional [BB89]. Directions [Ker81a, Knu79a]. Directly [HS04]. disco [Bec91b]. Display [Ado88, Tuf83, Fei88]. disquisition [Kra96]. Dissemination [EPO ]. diversions [WBS82, vdL92b]. division [Dés84b, Mañ87a]. dla [Paw93]. Do [Gau07]. Doceng [Kin05]. Document [Ala93, Ano15, Ben86a, BBv88, Bro91, CDMR85, Dil93, Dil99, FSS82a, FSS82b, Fur90, Geh86, Geh87, GL88, KLO78b, KLO78a, Ker84, Ker85, Ker89, Ker90, Kin05, KN14, KD93, KD95, KD99, Lam86, Lam94, NCNS82, Ohn91, PR92, Rei80, Sey82, Sri93, VC92, vV86, vV88, AJ92, De 88, Gel95, HNJ^{+}04, JLS03a, JLS^{+}03b, JLS^{+}04, KL82, Kin06, Krö90, Lin05, MH11, McC88, PC09, WBS82, Won90]. Documentation [CS99, Dés86, ISO87, Knu82e, KL93, Luc85b, Bri10, Knu83c]. Documents [Ano88, BKKW98, BEL04, GRM97, Hah91, Hah93, LV23, LS04, MH92, Shu94, Woh98, AP83, BKKW03, Di 07, Eic90, HB96, LA05, MMH07, Ves92, YF04, And83a]. Does [AMS90]. Doing [HW95a, HW95b, Hec03]. Dokument [Kat97]. dokumentów [Lam04, Wyr92]. Donald [Gho83, Hof82]. done [Wu08]. Don’t [CSR10b, Loz13, Van94]. Dot [Knu87]. draw [Kot23]. Drawing [BKW89, Gur94a, Hob95, Ohl95, Fin03, Fin04, Lev90, LS85]. driven [CM12]. Driver [Bee87, Gun85]. Druku [Tat88]. DTP [Ise90, Mor89, SP92]. Dublin [BMJ88, Mil84, Mil85, Mil86a]. Dutch [vdVM88]. Dvůr [Nad90]. DVI [Bee87, IM86a, IM86b, Ves92]. Dynamic [ELS04]. Dzielenie [Koł87].

e-Documents [BEL04]. E-TEX [Mit90]. e.V [Sch94]. e.V. [LS96]. Ease [Abi88]. easier [Lav92a, Lav92b]. easy [FS92, LU92b]. Editing [Bar86, Lee79, Pad04, Ste81, Swa86, But81, EE93, McK91, Wan96, Z^{+}94]. editions [HB95]. Editor [Bro91, Wal87b, CO91, Doo02, Ver00]. editoria [Bec91a]. Editorial [Swa86, Gar74a]. Editors [Bos86, Sch88a]. Edsger [FvGGM90]. Education [BU89]. Eesti [Ibr92]. Effectiveness [DM91]. Efficiency [Ano15, KN14]. Egyptian [Hug88]. eight [Vin92]. Einführung [Gün96, Kat93, Kat97, Kop90, Kop91b, Kop92, Kop94, Sch88b, Sch90a, Sch91c, Abd92, Abd93, Abd96, Kop88]. Einfürung [Kop91a, Kop95]. Einsatz [Bin93]. Einsteiger [Hei96]. eL^{A}TEX [Ala93]. Electronic [Ano87, BH86, EPO , Fur90, GR82, Gin94, HAB98, Man87b, VC92, Whi88, vV88]. elegant [Shi94]. Elementare [Tsc86]. Elements [Bri92, Bri96, Bri04, Bri12, Tob85, Wit85]. elettronica [Bec91a]. Embedding [HS04]. encodings [HH07]. encyclopedia [Hic89]. end [GSMN93]. Engine [LS04, Lin06]. Engineering [Kin05]. Engineers [Bue90]. English} [And92, App85, AA04, GD10, JHR88, Trz95]. Enseigner [And92]. ensure [Lav92b]. ensuring [Lav92a]. Entwicklungen [Zap07b]. Environment [Lau04, Nad90, Hay87, Lub90, Rug87, Wal82a, AMSV85]. Environments [Grä94a, Sie92a]. Envisioning [Tuf90]. EP [HAB98]. EP92 [VC92]. EPSILON [Knu78b]. Equation [McK91]. equations [Mat19]. equipment [Vin92]. ère [Har04]. Ergänzungen [Kop95]. Eric [Pop92]. Errors [DM91, Knu88, Knu89c]. Erweiterungen [Kop97]. Erweiterungsmöglichkeiten [Kop90, Kop91a]. Esq [Bux88]. Essays [Bar86, KSW03]. Essential [KD13]. established [De 88]. Esthetics [Sch04]. Euler [KZ89, Sie85]. Euro [TEX a]. Europe [BMJ88, Kar94c]. European [Cla90, Dés86, IEE92, Luc85b, Nad90, Zla92]. EuroTEX [Zla92]. EuroWorks [Kar94c]. Evaluate [DM91]. Evaluating [Ada91]. Evaluation [LFN04]. ever [Loz13]. everybody [Jur92]. everyday [DL94]. Everyone [Hah91, Hah93, Pop92]. Everything [Knu93, Loz13]. Evidence [Tuf97]. Evolution [BBW19, Grä09b]. Example [Bor92a, Bor93, Pap90a]. Examples [GR82]. Exeter [Cla90]. experience [KM92, Krö90]. Experiments [AS85]. Explanations [Tuf97]. Explicative [Bar86]. Explicit [Spi80]. Expos [Sch88a]. Expos/’e [Sch88a]. expression [Doo02]. expressions [Mat99, YF04]. extensible [Rug87]. extension [BTM^{+}01]. extensions [Hàn Thế Thành00]. Extracting [YF04]. extraction [Kan91].

F.R.S. [Bux88]. Factors [BM89]. Failure [HL87]. fall [Eij12]. Falling [KD15]. Families [MF87b]. Family [BH86, Knu80b, HB09, MF86a, MF86b, MF87a, Bee87]. fantastic [Dem89]. Faster [IM86a]. Features [Ano90b, Wil04]. February [ABD03]. Federal [VC92]. few [JHR88]. Feynman [Lev90, Ohl95]. FIFO [vdL92a]. Fifty [DE89]. File [CSR10a, Wal87a, Knu82c]. Files [CSR10b, LL94, PB03]. Finally [Grä94b].

First [Gin94, Grä99, Luc85b, Mil84, Pla06, Sam83, Gar74b, PC09]. Flexible [LS04]. float [MMH07]. FO [dSdOM^{+}05]. fomatto [Kar89]. Font [Ado90, Eng87, FK82, Hof82, Kar94b, Knu81, LNN00, Ada91, Dea92, FK85, Gho84, GC88, KKL89, LNN95, HG83, Knu82b]. Fontes [Har04]. Fonts [ELS04, Grä93b, Har93a, HH07, HBBG95, Hoe98, Kar92a, Knu90a, MF86a, MF86b, MF87a, MF87b, Píš04, Wil04, WH76, vB92c, vB93b, GC88, Gun85, Kar94c, Knu90c, PB03, Ueh90, Ven92, Vin92, Vos12]. Form [ISO87, Ika83, Tsc91b]. Formal [GB83]. Format [Ado90, CS99, Koh08, Ric78, San89]. Formate [Kar92a]. Formats [Ben88, CSR10a, Kar87, Kar92a, Kar94a, Wal87a, De 88, Ebe95]. Formatter [Fuj94, KV89, MH92]. Formatting [Ber06, Geh87, Sou92, BB89, Fis91, FSS82a, FSS82b, Geh86, GL88, Kin06, LA05, Noo83, SKS97, Wan96, Grä09c]. Forms [Tsc95c]. formula [HW97]. Formulae [Fuj94, Luc85a]. formulas [vdLCL89]. Fortgeschrittene [App88, App94, Hei96]. Forth [For89, Smi89, Smi90]. FORTRAN [Her69, Her70]. Forward [Kar92a]. Foundation [Grä13]. Four [Mil90]. Fourth [Ano91b, Mil87]. Fragen [Tsc87]. FrameMaker [Dar93]. framework [HM90a]. francais [Dés84b]. France [Dés86, HAB98, vV88]. Free [Grä09a, Cas93, EtPR94, Lub90]. French [And05, Dés84a, Dés85, Loz13]. Friendly [AMSV85]. Friends [vD12]. front [GSMN93]. Full [Sha01]. Fun [Knu90a, Knu90c]. Function [GKK14]. functional [HW97, Krö90]. Functions [HL87]. future [Kin06, Lav92a, Lav92b, Mit90, MR92b, Tay92]. Fuzz [MZH22].

Gaithersburg [Fur90]. Garmisch [BU89]. Garmisch-Partenkirchen [BU89].

Gebruikersgroep [NTG ]. Geleitwort [Kar92a]. gemacht [Haf97, WKS^{+}93a, WKS^{+}93b]. GENDRIV [Gun85]. General [IO89]. Generalised [LS85]. Generalized [Ano88, Bry88, Lag95]. Generalizing [PLA04]. Generated [HS04]. Generation [GKK14, Her87, HBBG95, Li85, Por97, KKL89, Kir05, SS90, Won86a, Hég84]. Genesis [Wit86]. gentle [Sie92a]. German [Kar92a, ST05]. Gestalt [Tsc87]. Gestalten [Sie89]. Gestaltungsbuch [Rud88]. Getting [HB95, IM86a]. Give [Wis01]. given [Sha01]. Global [Hol06]. Glossary [EAE^{+}94]. Glyphes [Har04]. Glyphs [Wil04]. GMD [App90]. GNU [CS99, GvdHPR14]. Goodbye [Kir05]. Gourmet [Spi80, Spi86a, Spi90]. goutte [Loz13]. GRAB [Les88]. Grade [Sam83]. Grafik [Klö02, Voß04, Voß05, Wer92]. Grafik-Import [Wer92]. grammars [EtPR94]. Grand [Knu89b, Knu90a, Knu90c]. grap [GL88, BK86, BK90]. graph [Hob01]. Graphic [Arn72, Tsc95c, Whi88, WH76]. Graphical [GKK14, Mol93, HM92]. Graphics [Den92, GRM97, GMR^{+}08, Hoe98, Ker81b, Ker82, Kol24, Lev90, Mar94, Rah89, Sow94, Wal87b, Whi83, Van80, BTM^{+}01, CO91, HBP04, Kot23, Mat05, Van81, Voß11a, Wer92, Zab82]. Graphik [Sow94]. Graphs [BK90, Hob95, BK86]. gray [Woo87]. Grayscale [HBBG95, Nai91]. Great [Whi90]. Greece [SBH^{+}04]. Greek [Fil04, Lik04]. Grid [Her87, Fei88, Hur78, JLS^{+}03b]. grid-based [Fei88, JLS^{+}03b]. Group [Cor86, SBH^{+}04]. Guida [Bec91a, Bia92]. Guide [Ala93, Ame01, Ano87, Bia92, BR93, Bry88, Des99, DFKvdV94, Fuj93, Hah91, Hah93, Hol89, KS87, Kol24, KD93, KD95, KD99, KD04, Lam86, Lam94, Lee79, LNN00, Spi80, Spi86a, Spi90, Trz95, Bla90, Dat17, Kot11, Kot23, LNN95, PC09, Rol93, Rol95, Vos12, Z^{+}94]. Guidelines [Mit90, SvdL89, Bri10].

GUST [BP04]. GUTenberg [Cah , Kir05].

Hahn [Pop92]. Halftones [Knu87]. Halftoning [Uli87]. Hamburg [WH86]. Handbook [Arn72, BR93, Gib96, Har92c, Hig93, Hig98, Hig20, Kra01, WMS04, But81, Tsc95b]. Handbuch [Lam95, Tsc28, Tsc95a]. handwritten [Mat99]. handy [JR92]. Happen [AMS90]. Happening} [Jac94]. Hart [Har91]. having [Kra96]. Hawking [Mag11]. hax [TEX b]. Hebrew [Por92]. held [HAB98, Mil84, Mil85, Mil86a, SBH^{+}04]. hello [Kir05]. help [DL94, Lav92a]. Here [Har02a]. Hermann [DE89, Kar92a, Zap87]. Hershey’s [WH76]. Herwijnen [Pop92]. HeVeA [Mar98]. Hierarchy [MH92]. Hieroglyphs [Hug88]. High [CDFV12, Cla87, MR92b, SvR92]. high-performance [SvR92]. High-Quality [Cla87, CDFV12]. Highest [MR92c]. Hilfsprogramme [Klö02]. hints [Dea92]. History [Bec77, Dru95]. HITAC [Hay87]. Horstman [Cor86]. Hortatory [Bar86]. hours [Dat17]. house [Vác92]. Houses [Sch88a]. Hrsg [Sch94]. HTML [GR99, Lei13, Mar98]. Hudson [McL80]. Human [BM89, vdVM88]. Human-computer [vdVM88]. Humanities [BS83, Rah87a]. Hungarian [WMS04]. Hy [Lia83, Lia84]. Hy-phen-a-tion [Lia83, Lia84]. Hypermedia [DD94]. hypertext [Gel95]. Hyphenation [App85, Fil04, FS92, Sie92a, Dés85]. hyphens [Mag11].

I. [Man87b]. IBM

[otB86, Doo02, FM85, Pap89, Vin92]. ICL [Lei85]. ICMS [HY14]. ideas [Kra96]. Idee [Kat97]. Ideograms [RSV16]. IFIP [DS83]. IFSGen4L^{A}TEX [GKK14]. II [Grä94a, Grä09b, Mil85, MFB^{+}23, Mor94b]. III [Cra87a, Grä94b, Grä09c, Mil86a].

IKARUS [Ika83]. Illuminating [TW87]. Illustrated [Lee79, McL95]. Illustrating [GRM97]. Illustrations [Spi80, Zap87]. Image [Wu08, Ada91, CO91]. Images [Ros84, Tuf97, Kot23]. Imagination [Dru95]. Imaging [AH89, HAB98]. impact [MH11, Pie91]. Impatient [ABH90]. impaziente [ABH91]. implementation [Di 07]. Implementations [FM85, Grä93b, Grä94a, Grä94b, Grä09b]. Implementing [Fil98, Lei85]. Import [Wer92]. Importance [DM89, Krö90]. Importing [Wer92]. Improved [Nai91]. Improving [Har02a]. inaczej [Wie92]. Inc. [Mil86b]. incognito [vdL92a]. Incorporating [Ven92]. Incremental [Fis91]. Independent [Utt85, BB89]. Index [CH87, CH88, Che90]. Indexes [BK88, Les88, Gar74b]. Indexing [Ame01, Col62, Cro80, Mul94, Mul05]. Indian [Gho83]. Industrial [For89, Wil83]. industries [YKT90]. Information [Bri10, Bee13, Har02a, Rah87a, Tuf83, Tuf90, II88, Myh90]. Informatique [And85]. initial [Krö90]. initials [Gar74b]. Ink [Arn72]. Inleiding [Sch90b]. Inputting [sT81, FPR85]. inquiries [Sim87]. InScript [GP89]. Insider [Sch88a]. Installation [App90]. Instant [DL94]. instead [Gar74b]. Institute [VC92]. instrumental [Tau92]. Intégrant [Ros84]. Integrated [BH86, Grä94a, Lau04, HM89]. integrating [GR99]. Interaction [vdVM88, MH11]. Interactive [AGQ85, GKK14, Pad04, Doo02, Fis91, Hic89, Rug87, AMSV85]. interchange [De 88, Krö90, vH89]. Interface [ASW89, GKK14, Nad90]. interfaces [LVC89]. Interfacing [Zab82]. Interleaf [Ude89]. International [Ano94, ABD03, BU89, BS83, Fur90, Gut04, HAB98, HY14, ICC88, Lev95, Mil84, Mil85, Mil86a, Mil87, Pie91, Sch91a, SBH^{+}04, VC92, WH86, vV86, vV88, Par90]. Interviews [BW09, LVC89]. Introduction [Grä93c, Gre05, Hob92, KR85, KS90, Liu13, Rub88, Urb86, Grä96, GC88]. Introduktion [Mad08]. Inverted [Les88]. iPad [Grä13]. Ireland [BMJ88]. IS&T [Pie91]. ISAAC [Gut04]. ISI [Gar74b]. Isn’t [Knu90b, Woo87]. ISO [Bri10, Krö90, vH85]. ISO/STPL [vH85]. ISSAC [Lev95]. Issues [Ker90, FSS82a, FSS82b, Gho84]. Italian [AP83]. Italy [ABD03, Luc85b]. Iterated [GKK14]. iTEXMac [Lau04]. IV [Grä11, Her69, Mil87]. Izdatel’skaiia [Z^{+}94]. iziaschnogo [Shi94].

JACM [CCL88]. Jahrbuch [Sch93]. jamais [Loz13]. Jane [Pop92]. janvier [And85]. Japanese [Ano90a, Fuj93, ACE^{+}08]. Java [Lik04]. Java-Based [Lik04]. jeden [Haf97]. jest [Bie90]. jointly [HAB98, SBH^{+}04]. Journal [Kor07, SvdL89, Wal82b, Win88, Sha01, POSxx]. Journals [Har02a, LK76, vH89]. Journ´ ees [And83a]. Joy [Spi80, Spi86a, Spi90, Spi93]. July [Gut04, Lev95]. June [ACM81, Ano94, BS83, Dés86, For89, AFI76]. just [CCL88].

Kent [DS83]. key [Hoo92]. Keyboard [FPR85]. Kézikönyv [WMS04]. Kingdom [Kin05]. Know [Sch88a, Loz13]. knowing [Hoo92]. Knowledge [ABD03]. Knuth [Gho83, Hof82, KD13]. komendach [Dyl94]. Komödie [DAN ]. Kompaktführer [Hei96]. Komputerowy [Dyl94]. Korea [HY14]. Kratkoe [Z^{+}94]. kxp [Kwo91]. kxpl [Kwo91].

labeling [Bei00]. labours [Bux88]. Labyrinth [Dru95]. Ładnego [Tat88]. Lamport [Mal19]. LAMS-TEX [Spi89]. Language [Ado85a, Ado85b, Ano88, Ben88, BK90, Bie88, Ker81b, Ker82, PLA04, Rei80, Rei88, SW92, POSxx, Van80, BK86, Fox87, Gou86, HBP04, IO89, Mor89, Van81, WBS82, You97, Bry88, Tun80]. Languages [CO04, ICC88, RS85]. Langues [Har92a]. large [Smi89]. Laser [Cer85, FN87]. LaserJet [Pap90b]. late [Bux88]. LaTeX [Kot23, MFW23, MFB^{+}23]. L^{A}TEX-Archiv [LS96]. L^{A}TEX-Begleiter [GMS94b]. L^{A}TEX-Buch [Abd92, Abd93, Abd96]. L^{A}TEX-Vademecum [Hei96]. L^{A}TEX2e [Bei00, Grä96, KD95]. LaTeX2HTML [Yua01]. L^{A}TEX3 [MR92a]. Latin [Kar94c, RS85]. laurea [Bec91b]. Lausanne [VC92]. Layout [ACE^{+}08, PR92, Atk04, CM12, Eic90, Fei88, HM92, HNJ^{+}04, HW97, JLS03a, JLS^{+}03b, JLS^{+}04, JMPS96, JMPS97, Lin05, Lin06, LFN04]. Layoutanpassung [Klö02]. LCCD [Tun80]. Learned [Knu83a]. Learning [GH97, GH16, Hec03, GH20]. lecons [And05, Spi86b]. Lectures [BU89, Zap87]. legacy [Row01]. Lehrbuch [Gün96, Sch91b]. leicht [Haf97, WKS^{+}93a, WKS^{+}93b]. Lernen [WKS^{+}93a, WKS^{+}93b]. Lesetypographie [WF97]. Leslie [Mal19]. Lessons [Knu83a, And05]. Letter [Knu80a, Knu80c, Tsc95c, vdLL89]. Letterform [Rug83]. Lettering [Tsc95c]. Letters [Dru95, Fir93]. Lettershape [GB83]. librarians [Gar74a]. Library [IM86a, IM86b]. Life [Fox98, Bux88]. Lifetree [Mil86b]. ligatures [Lag92]. Like [Woh98, Fat87, GP89]. Lilac [Bro91]. Line [AH05, BR93, Dil93, Dil99, Fin00, Hol06, Kan91, KR85, PK82]. linear [BBS01]. Lines [KP81]. lingual [Gho84]. Linux [Lip99]. Lion [Sie92c]. List [Pak05, Zap87, Pol93, Van94]. Lists [Har02b]. LITA [Cra87b]. Literacy [BH86]. Literate [Ben86b, Gur94a, Knu83b, Knu84b, Knu92, Pap90a, Sew89, SS91]. Livre [Ser89, Féd89]. Local [DS83]. Logical [Eic90]. logiciel [Dés84b, FPR84, Koł88b]. logon [McC88]. Loslegen [DL94]. Love [Pla06]. Low [Cer85, Les88, Thi91, FS92, Gun85, Lav92b]. low-cost [FS92, Lav92b]. Low-tech [Thi91].

Lua [Hag25]. Lucida [BH86]. ludzi [Paw93].

Mac [Lau04]. Machine [Nad90]. machines [Lav92a, Lav92b]. Macintosh [How86, KS87, SHS93, Mol93]. macro [Spi86a, Spi90]. macroistruzioni [Bec91b]. Macros [vB92d, vB93c, San89]. MACSYMA [Fat87]. MACSYMA-like [Fat87]. magazines [Hur78, POM12]. Mai [And83a]. maitrise [LR95]. make [Lav92a, Lav92b, Par90]. MakeIndex [Sch91a]. Makeup [KV89]. Making [DD94, LV23, Wal94]. Makropakete [App88]. Malo [HAB98]. Man [Nad90]. Man-Machine [Nad90]. Management [ABD03]. Managing [Prz04]. Mandarin [Liu13]. Manifesto [Sie86]. manifold [JLS03a]. Manipulation [ACM81, Fur90, VC92, vV86, vV88, And83b, And83a]. Manual [Ado85a, Ano82, Ano84, Ano03, Ano10, Ano17, BK90, HC88, Hob04, The03, Lab84, Lam86, Lam94, McL80, Sam83, Spi85, Spi86b, Spi91, WKS^{+}93b, Bau02, Col62, Fin03, Fin04, Ros92b, EE93]. manuels [Hob01]. Manuscript [AMS90]. Manuscripts [Ano87]. many [Hun94]. mapping [POM12]. Maps [Prz04]. Marāth¯ ı [AA04]. Marāth¯ ı-English [AA04]. March [HAB98, IEE92]. March/April [HAB98]. Mark [AJ88]. Mark-up [AJ88]. Markup [Ano88, Bry88, Koh08, Pad04, Shi94, Wit86, You97]. Maryland [Fur90]. master [KW90, Har93a]. Mastering [Whi83]. Matching [GH82]. Matematisk [Mor91]. materials [Bri89]. Math [AGQ85, Grä93c, Grä96, Grä00, Grä07, Grä16, Por97, vB92c, vB93b, Kna92, Ver00]. Mathematica [Soi95]. Mathematical [ABD03, Bor93, BEL04, ELS04, Hig93, Hig98, Hig20, HY14, Knu78a, Knu79b, KLR89, Kra96, Kra01, Mar94, Mor91, Pal86, Sch11, Soi95, Trz95, Doo02, EE93, Mat99, Rev91, Shi94, YF04, You97, vdLCL89]. mathematicheskikh [EE93, Shi94]. Mathematics [BP16, KC75, Ker81c, KZ89, Swa86, Day83, HW95a, HW95b, Hay96, Voß11b]. mathématiques [FPR84, Roy84]. MathML [Doo02, Pad04]. MathType [Har93b]. May [Hal87, Luc85b, NP89]. measures [HNJ^{+}04]. Mecum [Lip99]. Media [GR82]. Meeting [NP89, SBH^{+}04]. Memoir [Bux88]. men [Liu13]. Meta [Hof82, Knu81, HG83, Knu82b]. Meta-Font [Hof82, Knu81, HG83, Knu82b]. Metafont [Jac93, Ker81a, Sch94, Har93a, Kop90, CCL88, Des99, Knu79c, Li85, Sou85b, Tob85, KW90]. book [Knu86f]. Metafonte [Knu85]. metalanguage [Wit86]. Metamarks [Bil89]. Metamathematics [Hof82]. Metaphysics [Hof82]. Method [HS04, Ada91, Shi94]. Methoden [Kar92c]. methodology [CM12]. Methods [Cla90, Kar94b, Sow94, Wil83]. MF [MF86a, MF86b, MF87a, MF87b]. MG [Mar94]. Micro [Hàn Thế Thành00]. Micro-typographic [Hàn Thế Thành00]. Migrating [BP04]. Millennium [LS04]. minutes [McC88]. Mir [KM92]. mise [Sie92b]. Mitteilungen [Tsc86]. MKM [ABD03]. MLA [Gib96]. MlBibTEX [Huf04]. Mode [AGQ85]. Model [Lei13, CM12, Di 07]. model-driven [CM12]. Models [Ric78]. Modern [Fil04, Hug88, Vul92, Tsc95b, Knu80b, Knu86c]. modest [Sie92a]. modifiers [Mag11]. modular [Hur78]. molecular [AKWC82]. momentum [SS90]. monolithic [Vip96]. Monotonic [Lik04]. Montreal [Lev95]. Morris [McL95]. mots [Dés84b]. Mouse [Sie92c]. Moving [PHSB04]. ms [GL88, Bec91b]. MS-DOS [Bec91b]. multi [Day83, Gho84, MMH07]. multi-access [Day83]. multi-column [MMH07]. multi-lingual [Gho84]. multimedia [Ves92]. Multiple [Har93a, Bei00, Hol06]. multipurpose [KKL89]. Music [Fil98, Fox87, Gou86, Tau92]. MusicTEX [Tau92]. Musik [SKS97]. Mutation [DM91]. mv [GL88].

Nabor [L’v94]. naboru [EE93]. nadstavby [Bal97]. names [Gar74b]. Narrative [Tuf97]. naruby [Olš97]. National [AFI76, NP89]. native [GD10]. naturvidenskabelig [Mor91]. NC [BS83]. Nederlandstalige [NTG ]. Networks [DS83]. Neue [Tsc91a, Tsc28, Tsc95a]. News [Adoxx]. newsletter [Cra87b]. Newspaper [Har92c, Lag95]. newspapers [Hur78]. Newswriting [GR82]. Next [Por97]. Nice [vV88]. Nicely [BKW89]. nieco [Ryć93]. niepraktyczne [Jac93]. No [Van94, Kar89]. Non [App85, Pie91, GD10, JHR88, LS85]. Non-English [App85, JHR88]. Non-impact [Pie91]. non-native [GD10]. Non-Parametric [LS85]. norme [Sie92b]. North [BS83, WH86]. notation [Rev91, Sch11]. note [Bec91b]. Notes [Fir93, Shu94, Zap87]. Nottingham [vV86]. November [Kin05]. Novices [Tal12]. NP [KD15]. Number [Mao94]. Numbers [Wis01]. Numerical [Gar68, Pol93, Van94]. numérique [And92, Har04]. Numerous [Spi80]. Nuts [Ame01]. Nuts-and-bolts [Ame01]. n’y [Loz13].

o [Ryć93]. Object

[PHSB04, HM90a, HM92].

Object-Oriented [PHSB04, HM90a, HM92]. Objectif [GR95]. objective [Hol06]. objects [LA05, Zab82]. Observatory [Hec92]. Occidental [WH76]. October [Hec92, Mil84, Mil85, Mil86a, Mil87, Pie91]. ODA [Ves92]. Office [GvdHPR14, YKT90]. Offizin [Sch94]. oformleniia [Shi94]. Old [Fox97, AJ92]. On-Line [BR93, KR85]. One [Gri90, MZH22]. Online [NP89]. Only [CSR10a]. Open [Gau07]. opracowywania [Lam04]. Optimal [FK82, FK85, POM12, Pla81]. order [Gar68]. ordinary [Rev91]. ordinateur [Dés84b]. Oregon [ACM81]. Oriental [ICC88]. Orientales [Har92a]. Oriented [PHSB04, HM90a, HM92]. Original [Eng87]. Origination [EPO ]. oser [Loz13]. other [Sim87]. Ottmann [KSW03]. our [FvGGM90]. Outline [Dea92, PB03]. Output [IM86a, KV89, Lei13, Utt85, vB93a, Fat87, Fin03, Fin04, Pap90b]. Overhead [Les88]. Overview [Kwo91]. Oxford [Har91].

P [Man87b, KD15]. Package [Cas91, L’v94, Pap90a, Spi86a, Spi90, Voß12]. Packages [Gon04, Vos12]. pad [Ver00]. Page [Ale92, KV89, Wis01, Atk04, CYW^{+}03, Fin00, IO89]. Pages [Whi90, JMPS96, JMPS97]. Pagination [BKKW95, BKKW96, BKKW98, Cla87, Woh98, BKKW03, CDFV12, JMPS96, JMPS97, Lag95, MH11, POM12, Pla81]. Painters [Tsc95c]. pakete [L’v94]. Pandora [Bil89]. Paper [Arn72, AMS90]. Papers [BU89, Fuj93, Gib96, Trz95]. Paragon [Rug87]. Paragraphs [KP81, vB92c, vB93b]. Parametric [LS85]. Parametrization [Har93a]. parochial [JHR88]. Parsing [Smi90]. Part [BBW18, BBW19, AMS90, Kwo91, Mor94a, Mor94b, MFW23, MFB^{+}23]. Partenkirchen [BU89]. pasos [Gal92]. Past [Tsc95c]. Pattern [Di 07]. Pattern-based [Di 07]. Patterns [Fil04]. PC [otB86, Dem89, FM85, Pap89, Vin92, Wal87a, Ros92b]. PCTEX [Spi85, Spi86b, Shi94, Shi94]. PDF [HS04, PB03]. pdfTEX-Generated [HS04]. Pearl [Jac17]. Pearls [Ben86a]. People [BW09]. Perceptually [HBBG95]. perfect [Sie92a]. perfection [Bor90]. performance [SvR92]. Permuted [Che90]. Perq [Lei85]. Personal [Fuj93, Bia92, DE89].

Perspective [KSW03]. petit [Ser89]. Petites [And05]. Phase [DM89]. phen [Lia83, Lia84]. Philological [Won86b]. Philosophy [Zap87]. photo [Atk04]. Photocomposition [AK73]. Phototypesetting [BJK87]. Physics [Nad90, RSV16, vH89]. PIC [Ker81b, Ker82, Ros92a]. placement [MMH07]. Plain [Cla92]. Platform [PHSB04]. plots [Kot23]. podrecznik [Lam04, Wyr92]. Poles [JHR88]. policy [Gar74a]. Polish [Bie88]. Polishing [JR92]. polonais [Koł88b]. polskich [Koł87]. polskie [Wie92]. polyphonic [Tau92]. Polytonic [Lik04]. Popular [Wal87a]. Porting [Lub90]. Portland [ACM81, Pie91]. posobie [Shi94]. possibilities [Dem89]. Postprocessing [KV89]. PostScript [GRM97, Har93a, Rei88, Voß11a, YF04]. pour [Hob01]. power [Ude89, WBS82]. Practical [Grä14, Lab84, Pop92, Shu94, Trz95, van90, Dat17]. Practice [Wil83, vB92a, vB92b, vB92c, vB92d, vB93a, vB93b, vB93c, Knu91, Mor91, RKM88]. practitioner [Kot23]. Prague [Zla92]. praksis [Mor91]. praktyczne [Jac93]. pratique [Des99, Rol93, Rol95, Rol99]. praxisorientierte [Abd92, Abd93, Abd96]. Pre [Sie86]. Precision [LNN95, LNN00]. predicting [Ada91]. preface [Gho83]. Preliminary [Bil89]. Premiers [Urb91]. prepaging [FK85]. Preparation [Ala93, Ano15, CH87, CH88, Her70, KLO78a, Ker84, Ker85, Ker89, Ker90, KN14, KD93, KD95, KD99, Lam86, Lam94, NCNS82, GB82, KLO78b, KL82, Ohn91, PC09, Sey82, WBS82, AKWC82]. Preparing [Ano87, Fuj93, Shu94]. preprocessor [GP89]. presence [FS92]. Present [Tsc95c, McL95]. Presentation [Nai91, Sou92, II88]. Presentations [Voß12]. Presented [BU89]. Press [Har91, Ude89]. pretty [Ros92a]. Prime [Bro94]. Primer [Cla92, Kra96, SK95, Dea92, FN87]. primeros [Gal92]. Principles [GR82, Col62]. Print [Bro89]. Printer [Cer85, FN87, Pap90b]. Printers [Tsc95c]. Printing [Ben91, BK88, Dem89, Hug88, Pie91, Gou86, Lin06]. pro [Knu93, Ryb95]. Problem [sT81, CM12, Sou85a]. Problems [Vip96, Ueh90]. Proceedings [ACM81, Ano91b, BMJ88, Cla90, DS83, Dés86, Fur90, Hal87, IEE92, Kin05, Lev95, Luc85b, Mil84, Mil85, Mil86a, Mil87, AFI76, VC92, WH86, Zla92, vV86, vV88, Gut04, HAB98, NP89, HY14, Nad90, SBH^{+}04]. Process [Ano91b]. Processing [Abi88, CH88, Dil93, Dil99, Fil98, ICC88, Mil84, Mil85, Mil86a, Mil87, Mil86b, Pal86, Sri93, vV86, AJ92, BR87, Bec87, BBv88, Chr87, HM89, Rah87b, Won90]. Processor [Lab84]. Processors [Mil90, Mil92]. Produce [Win88]. Production [DM91, Lee79, AP83, Hur78, Ros84]. Profesjonalny [Bor92b, Bor97, Bor01]. Professional [DE89, KS87, HB95]. Professioneller [Sch89, Sch91b]. professionels [Féd89]. Program [BJK87, Knu86a, Knu86b, Knu90b, Rei88, Sew89, SvR92, Thi91]. Programmer [Hol89, Wal87a]. Programmiertechniken [App88]. Programmierung [Klö02]. Programming [Ben86b, Ben86a, Doo02, Gur94a, Her69, Hol87, Knu83b, Knu84b, Knu92, Sew89, SS91, Pap90a]. Programs [BM89, Mil86b, Kon01, Lub90, San89]. Project [Sie85]. Proof [Knu90b]. Proofreading [Swa86]. properties [AH05]. Prostoi [Shi94]. PROTEXT [Mil84, Mil85, Mil86a, Mil87]. proto [Doo00]. proving [L’v94]. Przewodnik [Dyl94, Wyr92, Lam04]. przygotowywania [Wyr92]. przykładami [Dyl94]. PSTricks [Voß04, Voß05, Voß11a]. Psychological [Ano84]. Psychonomic [vdVM88]. pu [Lia83, Lia84]. Publication [Ano84, Ear87, Jac94, Nel78, Nel91, Adoxx]. Publiceren [dBvLV88]. publikacji [Bor92b, Bor97, Bor01]. Publish [Yua01]. Published [AMS90, Bri89, Kra96]. Publisher [Gil89, Ise90, Jan86]. Publishers [Ano87, KM92]. Publishing [Bai70, Bai90, Bar86, BBW18, BBW19, EPO , Fur90, Hec92, HC88, Lun92, Man87b, Myh90, Sch88a, Ude89, VC92, vV88, Bla90, CDFV12, Cra87a, FN87, HAB98, SČS92, Vác92, YKT90]. purpose [IO89]. Pursuit [MR92c]. Putting [Wis01]. Pwn [CSR10a].

Q&AIX [Fox97, Fox98]. Quality

[Cla87, MR92c, Ada91, CDFV12, MR92b]. quand [Loz13]. Quantitative [Tuf83]. Quantities [Tuf97]. que [Loz13]. queue [Pap90a]. quick [PC09, Vos12].

R [Lei13]. Rôle [Utt85]. Raleigh [BS83]. rapporter [Mor91]. Raster [AH89, HAB98]. rasterization [Dea92]. rather [SP92]. Rationalized [FPR85]. read [Kra96]. readability [MBHL91]. Readable [BM89]. Readers [Har91, Wis01, Kir05]. ready [DL94, JR92]. Real [Rot88]. recipes [LU92b]. Recognition [Mat99]. Recommendation [Bri89]. Recommendations [Bri76]. Reconsidered [BKKW95, BKKW96]. Reconstruction [PR92]. recorded [Kra96]. Recovering [PR92]. Recursive [SS90]. redaktirovaniiu [EE93]. REDUCE [ASW89]. ReDuX [Mor94a, Mor94b]. referees [Gar74a]. Reference [Ado85a, Ado88, Eij92, Hah91, Hah93, Har02b, Hol89, The03, Lam86, Lam94, LNN00, Wal87a, Bau02, Ebe95, Fin03, Fin04, LNN95, Pol93, Van94, Vos12]. References [Ame78, Grä09c, ISO87, Wis01, Bri50, Bri76, Bri89, Bri10, Gar68, Hun94, O’C92]. Referenz [Gün96]. Referenzhandbuch [SP97]. regions [dSdOM^{+}05]. règles [Féd89]. Relevance [MZH22]. Remark [HM90b]. remarks [JHR88]. Rennes [And83a]. Repertory [WH76]. Report [Sch88a]. Reports [Her70, Mor91]. Representation [Kar92a]. Requirements [ACE^{+}08, RKM88]. Research [Ano15, BU89, Gib96, Gin94, KN14, Mil86b, GD10]. Resources [Prz04, Bri10, GC88]. Respond [Gau07]. Responsibilities [Sto92]. Results [Píš04]. Retrospective [Ker84, Ker89]. reusability [Pap90a]. Review [Gon04, Ker81a, Mil86b]. reviews [PH92, Pop92]. Revision [Bee87]. revival [Har92b]. Revolution [Grä09b, Wis01]. Riad [BK85]. RIDT [HAB98]. Ring [DS83, UCBB83]. Risø [Mor91]. Risø-rapporter [Mor91]. road [Sie92a]. Rochester [For89]. Roman [MF86b]. Routine [Lev90]. Routines [vB93a]. Ru [Liu13]. Rukovodstvo [EE93, Z^{+}94]. Rules [Col91, Har91, Sou92]. run [Dés84a]. Russian [Vin92].

S [Knu80a, Knu80c]. S.A.PR.I.S.T.I [Ros84]. Safe [CSR10a]. SäferTEX [Sch04]. Salute [FvGGM90]. sand [Sim87]. sans [Loz13]. Santander [Gut04]. Satzkunst [SP92]. Satzsystem [Haf97]. savoir [Loz13]. sazeciho [Bal97]. scalable [BTM^{+}01, Mat05]. Schaffende [Tsc28, Tsc95a]. Schnell [Kna97]. School [Nad90]. Schools [Tsc95c]. Schriften [Kar92a, Kar86]. Schriftenreihe [Sch94]. Schriftstatistik [Kar92b]. Schrifttechnologie [Kar92c]. Science [Bai70, Bai90, Eij12, GD10, KSW03, Yua01, Knu91]. Sciences [Hec92, Hig93, Hig98, Hig20, Kra01]. Scientific [CDMR85, Dés86, GvdHPR14, Her72, Luc85b, Mil86b, Mil92, Nad90, Dat17, Mor91, Ano97, HW95a, HW95b, HB95, HB96, Mil86b]. scientifica [Bec91b]. scientifically [ST05]. Scientists [Bue90, SK95]. scores [Fox87]. screen [MH11]. Scribe [Rei80]. script [HBP04, Mac83]. script-based [HBP04]. Scripts [Gho83]. Scriptwriting [Hay77].

Scrolling [MH11]. Sculptors [Tsc95c]. SDDS [CDMR85]. searching [MH11]. Second [ABD03, Dés86, Mil85]. segmentation [CYW^{+}03, Di 07]. Selected [Zap87, DE89]. Semantic [Koh08, Mat05]. sense [Cra87a]. Seoul [HY14]. September [BMJ88, CS99, DS83, Fur90, ICC88, SBH^{+}04, Zla92, Nad90]. sequence [Bei00]. server [UCBB83]. service [KR85]. set [San89]. Settings [Har02b]. Seventh [Pie91]. SGML [Bar89a, Bry88, Hic89, van90, vdLCL89, Pop92]. shade [Bei00]. shading [Bei00]. Shapes [Bil89]. Sharing [KLO78a]. Sharp [Man87b]. Shotai [Kar89]. Shotai-no-dijitaru-fomatto [Kar89]. Sign [Tsc95c]. SIGOA [ACM81]. SIGPLAN [ACM81]. Simple [Cla87, Knu90b, Shi94]. Simplement [Ros84]. Simplified [Grä93c, Gre05]. simulated [Lag95]. Singapore [Ano94]. Single [Doo90]. sistema [Bec91a]. Sixth [BS83]. size [MH11]. sizes [MBHL91]. Skalský [Nad90]. skills [FN87]. skład [Bor92b, Bor97, Bor01, Dyl94]. Slavic [MF87b]. Small [And05, MBHL91]. Society [vdVM88]. Soforthelfer [DL94]. Software [Cor86, DM91, HY14, Mil86b, Ude89, Win88, Ohn91, Wal87a]. solution [FS92, Lav92a]. Solved [MZH22]. solving [BBS01]. Some [Fir93, Li85]. Sonderheft [Tsc86]. sophisticated [Lub90]. Source [Gau07, Sch04, Tsc95c]. sourcebook [Bry96]. Sources [Píš04, KW90]. sous [Hob01]. South [HY14]. Space [Hec92]. Spain [Gut04]. Spanish [Mañ87a]. speakers [GD10]. Speaking [Hay96]. SPEC [Whi87]. Specification [Rei80, TM88]. Spectra [DM89]. Speicherung [Kar86]. Splines [LS85]. sposob [Shi94]. Springer [Bin93]. Springer-Verlag [Bin93]. Square [Doo93]. środowiskach [Dyl94]. St [HAB98]. Stammtisch [Sch93]. Standard [AH05, Krö90, vH85, Ame78, Bry88].

Standards [AJ88, HC88, MR92c, Krö90, vH89]. Stanford [McC20, Sie85]. Stanford’s [KW90]. start [DL94, PC09]. started [HB95]. Starting [Doo93]. State [BS83]. Statistical [Lei13]. statt [SP92]. Steps [Gin94, Grä99]. STET [Bos86]. stile [Bec91b]. sto [Doo00]. Storage [Les88]. Stories [Zap07a]. Story [Mao94]. STPL [vH85]. Strangers [CSR10b]. Strasbourg [Dés86, Hec92]. STRATEC [FPR84, FPR85]. Strategies [Hoe98]. String [GH82]. strings [Smi90]. Structural [Fuj94]. Structure [BJK87, ISO87, PR92, MR92b, MR92a]. Structured [Knu82e, KL93, MH92, Noo83, Won90, Knu83c]. structures [AKWC82, Eic90]. student [HB95]. Studies [WH86, Bil89]. Study [Gau07, PLA04]. Style [Ano82, Ano03, Ano10, Ano17, Bri92, Bri96, Bri04, Bri12, HC88, OC90, SvdL89, Tob85, MH11, Vip96]. submitting [vH89]. Substituting [PB03]. successful [Hoo92]. Suite [GvdHPR14]. Summary [BB85]. Summer [Nad90]. Support [PLA04, dSdOM^{+}05, Hoo92]. Supposed [Sch88a]. Survey [NCNS82, Rah89, Sow94, FSS82a, FSS82b]. Swiss [VC92]. Switzerland [VC92]. Symbol [Pak05]. Symbolic [Fil98, Gut04, Lev95]. Symbols [WH76, Vos12]. Symposium [ACM81, BMJ88, Gut04, Lev95, Kin05]. Syntax [Pad04, Voß12]. Synthesis [Spi89, Lin05, SvR92]. System [Ado88, BK91, Fil98, Geh87, Her70, Her72, Hob97, KC75, KLO78a, Ker89, Knu78b, Knu79c, Knu82e, KL93, Lam86, Lam94, Mar94, PR92, Rub88, Wyr92, AP83, Chr87, Geh86, GL88, Hàn Thế Thành00, Hur78, KKL89, Lam04, Smi89, Tut83, Z^{+}94, Olš95, CDMR85, Knu83c]. systema [Z^{+}94]. Système [Ros84, Har94]. systemie [Dyl94, Koł87]. Systems [Ano15, Cla87, Ear87, GKK14, Ker90, KN14, Mil84, Mil85, Mil86a, Mil87, NCNS82, Rug83, Day83, EP87, Fat87, FSS82a, FSS82b, HM90a, HM92, Pla81, Sey82]. systemu [Bal97]. Sztuka [Lao86].

Table [vdL92b]. Tables

[Ker81c, Lei13, WH76, SKS97, Voß11c]. Tabular [Wan96]. Tafel [SKS97]. Take [CSR10b]. Takes [Jac94]. Tangled [Mor93]. TAU [Knu78b]. TCI [Mil86b]. Teach [Chu92]. Teaching [And92, AS85]. tech [Thi91]. Technical [AK73, Ame01, otB86, Her93, Jan86, Knu78b, LK76, Shu94, Spi80, Ste81, Zap07a, Ude89]. Techniques [DFM97, Dil93, Dil99, KD04, Nad90, Píš04, WH76, Pla81, Rah87a]. technischer [Zap07b]. Technologies [Pie91]. Technology [Kar94b, VC92, DS83, Dea92, Rah87a, Ueh90]. Teil [WKS^{+}93a, WKS^{+}93b]. tekstov [Shi94]. tekstu [Dyl94]. Templates [vdLL89]. ten [McC88]. Tennessee [Eij12]. ter [Lia83, Lia84]. Terms [EAE^{+}94]. tesi [Bec91b]. Test [Knu86d, Knu82c, Knu84c]. Testing [MZH22]. tex [Har94, Ker81a, Spi80, KW90]. TEX- [LS96]. TEX-/L^{A}TEX-Archiv [LS96]. TEX-based [Wit86]. TEX-Euro [TEX a]. TEX-Installation [App90]. TEX-REDUCE [ASW89]. TEX/L^{A}TEX [Sow94, Fuj94, Sow94]. TEX/ [KW90]. TEX/WEB [Roy84]. TEX82 [Knu82a]. TEX88 [Cla90]. Texas [Hal87]. TEXbook [Knu84a, Knu89a, Olš97, Sal95]. TEXcetera [Bro94]. TEXHelp [BR93]. TEXikon [SP97]. TeXinfo [CS99]. TEXiraamat [Ibr92]. TEXline [TEX c]. TEXMACS [GvdHPR14]. TEXmag [TEX d]. TEXnical [Myh90]. TEXnician [Eij92]. TEXnicolor [Ano90a]. TEXnika [Tat88]. TEXnische [DAN ]. TEXnology [Pap89]. TEXowania [Lao86]. texreg [Lei13]. Text [ACM81, Abi88, AK73, ACE^{+}08, BR87, Ber06, Bro89, CSR10a, Cla87, Fuj94, GB82, Kan91, KV89, Knu78b, Mil84, Mil85, Mil86a, Mil87, Pal86, Ric78, Sou92, Spi80, Ste81, vV86, And83b, BB89, Chr87, CO91, EtPR94, Fis91, Noo83, vH89, Gho83]. Text-Only [CSR10a]. textbook [Shi94]. texte [Bor90]. Textes [Ros84, FPR84, Koł88b, Roy84]. textos [Val97]. textov [EE93]. texts [EE93, Shi94]. TEXtures [Ale92]. Textverarbeitung [Mau97]. TEXware [Knu82d, Knu86e]. Thames [McL80]. theme [Hun94]. theological [Won86a]. Theory [Knu91]. Third [Cla90, Mil86a, Won86a]. Thomas [KSW03]. Three [Mil92]. TikZ [Kol24, Kot23]. til [Mad08]. Time [Bro94, Gri90, KLO78a]. Time-Sharing [KLO78a]. tion [Lia83, Lia84]. Tipps [Klö02]. Tips [Dil93, Dil99, Shu94]. titles [Gar74a, Sha01]. together [Hoe92]. Tokens [vB92d, vB93c]. Too [Hun94]. Tools [BK88, Kar94b, Ker84, Ker85, Ker89, KD04, Píš04, Rah87a, RKM88]. Topic [Eij92, Prz04]. Topography [AH89]. Toronto [ICC88]. Torture [Knu86d, Knu84c]. toujours [Loz13]. Tour [Mol93]. Tout [Loz13]. Trade [Bos86]. Tradeoffs [Ker90]. traditional [Har92b]. traitement [Bor90, FPR84, Koł88b, Roy84]. Trajectories [Hob86]. transfer [Wu08]. Transition [Hal87, Wal82b]. translate [Gel95]. Translating [Ves92, Rev91]. translation [MBO93]. translator [Mar98]. Transporting [Hay87]. traps [WBS82]. Treasury [Tsc95c]. Trees [BKW89, SS90]. tri [BB89]. tri-directional [BB89]. Tricks [Bos86, Klö02]. triroff [BB89]. troff [EP87, GL88, Dic87, BB89]. TrueType [LL94]. TUG [SBH^{+}04]. Tuned [HBBG95]. TURES [KS87]. Tutorial [Ado85b, Bee13, BK90, Hah91, Hah93]. Two [Bro91, FM85, Mil86b]. Two-View [Bro91]. Type [Ado90, Bro89, Cra71, Dai67, DM89, Her93, Lab84, LNN00, Lun92, McL95, Por97, Rub88, SG74, Swa86, Whi87, WH76, Zap87, Gho84, LNN95, Sou85a, Gho83, GB83, Píš04]. Typeface [KZ89, Law90, MBHL91, Rug87]. Typefaces [Bla90, Ika83, Kar87, Kar94a, Knu80b, Knu86c, Sou85b, SB90, Zap87]. Types [BH86]. Typeset [AA04, Lab84, Kra96, McC88, Ros92a]. Typesetters [Lab84]. Typesetting [AJ88, BE92, Bee04, Ben88, BK90, Cas91, CO04, Day83, Dic87, Doo89, EAE^{+}94, EP87, Fuj94, Geh87, GSMN93, Gil89, Hah91, Hah93, Hoh88, How86, Kel85, KS87, KC75, Ker81a, Ker81b, Ker81c, Ker82, Knu79a, KD04, Lag92, LK76, MR92c, Mol93, Ros92b, Sch04, SW92, Soi95, Spi80, Spi86a, Spi90, Sri93, Tut83, Voß11b, Voß11c, Wal82a, WH76, Won86b, Woo87, Van80, vdL92c, AKWC82, BR87, BK86, Cra87a, Cra87b, FN87, Fox87, Geh86, GL88, Gho84, GB82, Hàn Thế Thành00, Jan86, KR85, MR92b, Myh90, Neg88, Pap91, Pla81, Sey82, Thi91, UCBB83, Vác92, Van81, Won86a, Sie86]. typing [EE93, L’v94]. typografi [Mor91]. Typografický [Olš95]. Typographers [McL95, Tsc95c]. Typographic [Bri92, Bri96, Bri04, Bri12, Har02b, Her70, OC90, Wil04, GC88, Hàn Thế Thành00, MBHL91]. Typographie [And85, Rud88, Tsc28, Tsc95a, And92, And05, Tsc86, Tsc91a]. Typographique [Féd89]. Typographische [Tsc86]. Typographisches [Sie89]. Typography [BM89, BD83, Col91, Fur90, Her69, Her72, Her81, Knu78a, Knu79b, Knu99, Kra01, LS04, McC20, McL80, MR92c, Rub88, SP92, SBH^{+}04, VC92, Wal82b, vV88, And92, And05, Bry96, Har92b, HAB98, Mor91, STS03, Tsc95b, Wit85, Dés85].

użytkownika [Wyr92]. Überblick [Sow94]. uchebnoe [Shi94]. uczytkownika [Lam04]. UK [DS83]. UKTEX [UKT ]. ultimate [Ebe95]. Unbound [Hoe98]. Understanding [Hol87, Lun92]. Unicode [AH05]. Uniformity [Gar74a]. United [Kin05]. universelles [Haf97]. University [BS83, DS83, For89, Gut04, vV86, Eij12, Har91]. UNIX [BR87, Chr87, EP87, Geh86, Geh87, GL88, KLO78a, KL82, Ker89, LK76, Lub90, Mau97, Sri93, Tut83, Ker84, Ker85]. Unlocking [Kol24]. URW [Kar94c]. USA [Pie91, Mil87]. usage [Féd89]. Use [Bar89a, CSR10a, DM91, Nai91, Bia92, JR92, Lav92a, Lav92b, Loz13, RS85, TM88, vH89, Dés85, Sie86]. Used [Ano15, KN14, Ros92a, Win88]. User [AMSV85, BK90, BR93, GKK14, Hob04, KS87, Lam86, Lam94, Bla90, Fin03, Fin04, LVC89, Ros92b, RKM88]. User-Friendly [AMSV85]. Users [KD93, KD95, KD99, SBH^{+}04, Hoo92]. Uses [Cla90, II88]. Using [AA04, Chu92, Eng87, Grä93c, GH82, Hah91, Hah93, Hic89, IM86a, IM86b, Koh08, LV23, Man87b, Pad04, Tri88, Bei00, Dic87, Gun85, How86, JHR88, Kot23, Lag95, STS03, Tau92]. uso [Bia92]. utilisateur [Hob01]. utiliser [Loz13]. Uzivatelske [Bal97].

v [L’v94, Grä13]. Vade [Lip99]. Vademecum [Hei96]. Väike [Ibr92]. variable [Lin06]. variation [MBHL91]. variations [Hun94]. VDM [BMJ88, TM88]. VDM-Europe [BMJ88]. vector [BTM^{+}01, HBP04, Mat05]. Vein [Ale92]. ver.0.93 [FPR84]. ver.2.3 [Knu83c]. Verfahren [Sow94]. Verheissung [Won86a]. Verlag [Bin93]. Version [Grä94b, Knu82e, KL93, LNN00, Sch91a, Spi80, CS99, Knu82d, LNN95, Mil86b]. Versprechen [Won86a]. verstka [L’v94]. versus [Gar68]. verze [LU92a]. via [JLS03a, vdL92c]. View [AJ88, Bro91, Doo89, Gil89, Ise90, Jan86]. viewed [YKT90]. Vieweg [Abd92, Abd93, Abd96]. VIIth [BU89]. vima [Doo00]. Virtual [Knu90a, Knu90c]. vision [CYW^{+}03]. vision-based [CYW^{+}03]. Visual [Her93, Tuf83, Tuf97, Den92, LFN04]. Visualization [GKK14, HM92]. visualizing [HM90a]. VM [AMSV85]. VM/CMS [AMSV85]. Vol [GL88]. Volkswriter [Mil86b]. volumes [AKWC82]. VOS3 [Hay87]. Voskhititel’nii [Spi93]. voulu [Loz13]. vous [Loz13]. vowels [Lag92]. Vse [Knu93]. VTEX [Cas91].

W [FvGGM90, Dyl94, Koł87]. want [Pol93, Van94]. wanted [Loz13]. Wat [EP90]. Way [BMJ88, Lav92b]. weapon [Ebe95]. Weaving [Sew89]. WEB [Knu82e, Knu83c, CM12, GR99, Mat05, Yua01, Roy84]. Wegweiser [Det97]. Well [De 88]. Well-established [De 88]. were [BU89]. Werkzeuge [Kar92c]. WG6.4 [DS83]. WG6.4/University [DS83]. which [Voß12]. Whose [Knu90b]. Wider [Sie86]. will [Gar74a]. William [McL95]. Windows [DF99, HB95]. Wissenschaftlich [ST05]. within [MH11]. without [Loz13]. Wizard [Spi91]. Wizards [Knu89b, Knu90a, Knu90c]. Word [Lab84, Lia83, Lia84, Mañ87a, Mil86b, Mil90, Mil92, Bec87, HB95, HB96, Mat19]. Word-Processing [Mil86b]. WordPerfect [McK91]. Wordprocessors [otB86]. Words [App85, SG74, Rah87b]. Wordworking [Cra87a]. Work [DD94, Ker85, Tal15, Wal94, Hoe92]. Working [ST05]. WorkPlace [Ano97, HB95, HW95a, HW95b, HB96]. Works [Mal19]. Workshop [DS83]. Workshops [Chu92]. workstation [Gho84]. Workstations [Ear87]. World [BW09, Rot88]. wraz [Dyl94]. write [Tau92]. Writers [Ame01, Bos86, Gib96, Kir05]. Writing [Ano91b, Bar86, GH20, Gur94b, Hig93, Hig98, Hig20, KLR89, Kra96, Sch88a, Trz95, Dat17, GD10]. Wuppertal [Sch93]. wyrazów [Koł87]. WYSIWYG [Grä11].

X [Lau04]. Xanthi [SBH^{+}04]. Xerox [HC88]. XML [SBH^{+}04, BP04, GR99, Prz04]. XSL [LA05, dSdOM^{+}05]. XSL-FO [dSdOM^{+}05]. XXXII [WH86]. XY [Ros92b]. XY-pc [Ros92b]. Xy-pic [Ros92a]. XYZapf [DE89].

Years [DE89]. Yellow [JMPS96, JMPS97]. Yellow-Pages [JMPS96, JMPS97]. York [AFI76, NP89].

z [Dyl94]. zacátecníky [Ryb95]. Zapf [DE89, Kar92a, Zap87]. zastosowania [Jac93]. Zeichensätze [Klö02]. zeitgemäss [Tsc28, Tsc95a]. Ziel [Kna97]. znaki [Wie92]. zum [DL94, Kat97].

References

Athale:2004:ULT

[AA04] Manasi Athale and Rahul Athale. Using L^{A}TEX to typeset a Marāt.h¯ ı-English dictionary. In Syropoulos et al. [SBH^{+}04], pages 55–58. CO-DEN LNCSD9. ISBN 3-540-22801-2 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN Z253.3 I58 2004. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t3130.htm;
http://www.springerlink.
com/openurl.asp?genre=
issue&issn=0302-9743&
volume=3130;   http:
//www.springerlink.

com/openurl.asp?genre=
volume&id=doi:10.1007/
b99374.

Abdelhamid:1992:VLB

[Abd92] Rames Abdelhamid. Das Vieweg L^{A}TEX-Buch: Eine praxisorientierte Einführung. Friedrich Vieweg und Sohn, Braunschweig, Germany, 1992. ISBN 3-528-05145-0. xiii + 169 pp.

Abdelhamid:1993:VLB

[Abd93] Rames Abdelhamid. Das Vieweg L^{A}TEX-Buch: Eine praxisorientierte Einführung. Friedrich Vieweg und Sohn, Braunschweig, Germany, second edition, 1993. ISBN 3-528-15145-

5. xvi + 169 pp.

Abdelhamid:1996:VLB

[Abd96] Rames Abdelhamid. Das Vieweg L^{A}TEX-Buch: Eine praxisorientierte Einführung. Friedrich Vieweg und Sohn, Braunschweig, Germany, third edition, 1996. ISBN 3-528-15145-

5. xviii + 226 pp. 44,00 DM. The ISBN did not change between second and third editions. Asperti:2003:MKM

[ABD03] Andrea Asperti, Bruno Buchberger, and James Harold Davenport, editors. Mathematical Knowledge Management: Second International Conference, MKM 2003 Bertinoro, Italy, February 16–18, 2003, volume 2594 of Lecture Notes in Computer Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2003. ISBN 3-540-00568-4. 65.50 SFR; 40.66 EUR.

Abrahams:1990:TI

[ABH90] Paul W. Abrahams, Karl Berry, and Kathryn A. Hargreaves. TEX for the Impatient. Addison-Wes-ley, Reading, MA, USA, 1990. ISBN 0-201-51375-7. xvii + 357 pp. LCCN Z253.4.T47 A27 1990.

Abrahams:1991:TI

[ABH91] Paul W. Abrahams, Karl Berry, and Kathryn A. Hargreaves. TEX per l’impaziente. Addison-Wesley, Reading, MA, USA, 1991. ISBN 88-7192-022-8. 396 pp. This is a translation to Italian of [ABH90], with additional material on Italian hyphenation.

Abikoff:1986:T

[Abi86] William Abikoff. TEX.

The Mathematical Intelligencer, 8(3):64–76, 1986.

Abikoff:1988:TEA

[Abi88] William Abikoff. TEX:

The ease and art of text processing. Abacus, 5 (4):10–29, summer 1988. CODEN ABACEJ. ISSN 0724-6722.

Anan:2008:RJT

[ACE^{+}08] Yasuhiro Anan, Hiroyuki Chiba, Junsaburo Edamoto, Richard Ishida, Keiichiro Ishino, Tatsuo Kobayashi, Toshi Kobayashi, Kenzou Onozawa, and Felix Sasaki. Requirements of Japanese text layout. Working draft, W3C, October 15, 2008. URL http:/

                                        /www.w3.org/TR/2008/
                                        WD-jlreq-20081015/.

ACM:1981:ASS

[ACM81] Proceedings of the ACM SIGPLAN SIGOA Symposium on Text Manipulation, Portland, Oregon, June 8–10, 1981, volume 16(6) of ACM SIGPLAN Notices. ACM Press, New York, NY, USA, June 1981. ISBN 0-89791-043-5. LCCN QA76.7 .S54 v.16:6.

Adams:1991:EFQ

[Ada91] Debra A. Adams. Evaluating font quality: a method of predicting character image defects. In Pietrowski [Pie91], pages 353–364.

Adobe:1985:PLR

[Ado85a] Adobe Systems Incorporated. PostScript Language Reference Manual. Addison-Wesley, Reading, MA, USA, 1985. ISBN 0-201-10174-2. ix + 321 pp. LCCN QA76.73.P67 A33 1985.

Adobe:1985:PLT

[Ado85b] Adobe Systems Incorporated. PostScript Language Tutorial and Cookbook. Addison-Wes-ley, Reading, MA, USA, 1985. ISBN 0-201-10179-3. x + 243 pp. LCCN QA76.73.P67 A34 1985.

Adobe:1988:DPS

[Ado88] Adobe Systems Incorporated, 1585 Charleston Road, P. O. Box 7900, Mountain View, CA 94039-7900, USA, Tel:

(415) 961-4400. The Display PostScript System Reference, October 10 1988.

Adobe:1990:ATF

[Ado90] Adobe Systems Incorporated, 1585 Charleston Road, P. O. Box 7900, Mountain View, CA 94039-7900, USA, Tel:

(415) 961-4400. Adobe Type 1 Font Format, March 1990.

Adobe:19xx:CAS

[Adoxx] Adobe Systems Incorporated. Colophon—Adobe Systems News Publication, 19xx.

NCC:1976:ACP

[AFI76] AFIPS Conference Proceedings of the 1976 National Computer Conference, June 7–10, 1976, New York City, volume 45. AFIPS Press, Montvale, NJ, USA, 1976. ISBN ?? LCCN ??

Andre:1985:TIM

[AGQ85] Jacques André, Yann Grundt, and Vincent Quint. Towards an interactive math mode in TEX. In Lucarella [Luc85b], pages 79– 92. ISBN 0-201-13399-7. LCCN Q224.E97 1985.

Andre:1989:RID

[AH89] Jacques André and Roger Hersch, editors. Raster Imaging and Digital Topography. Cambridge University Press, New York, 1989. ISBN 0-521-37490-1. LCCN Z253.3 .R37 1989.

Asmus:2005:LBP

[AH05] F. Asmus and A. Heninger. Line breaking properties: Unicode Standard Annex 14, line breaking properties. Technical report, Unicode Consortium, 2005. URL

                                        http://www.unicode.
                                        org/reports/tr14/.

Arrabito:1988:CBT

[AJ88] R. Arrabito and H. Jürgensen.

Computerized Braille typesetting: Another view of mark-up standards. Electronic Publishing—Origination, Dissemination, and Design, 1(2):117–131, September 1988. CODEN EPODEU. ISSN 0894-3982.

Allport:1992:OND

[AJ92] G. Allport and P. Jarratt. The old and the new in document processing. Electronic Library, 10(1): 41–45, February 1992. CODEN ELLIDZ. ISSN 0264-0473.

Alt:1973:CPT

[AK73] F. L. Alt and J. Kirk.

Computer photocomposition of technical text. Communications of the Association for Computing Machinery, 16(6):??, June 1973. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Allen:1982:CCD

[AKWC82] F. H. Allen, O. Kennard, D. G. Watson, and K. M. Crennell. Cambridge Crystallographic Data Centre. 6. Preparation and computer typesetting of “molecular structures and dimensions” bibliographic volumes. J. Chem. Inf. Comput. Sci, 22(3):129–138, August 1982. CODEN JCISD8. ISSN 0095-2338 (print), 1520-5142 (electronic).

Alamnehe:1993:EDP

[Ala93] Abass Alamnehe. eL^{A}TEX: Document Preparation Guide. EthiO Systems Co., P.O. Box 36921, Houston, TX 77236, USA, 1993. xvi + 230 pp. This book is typeset in eL^{A}TEX, an Ethiopicbased version of L^{A}TEX, and is entirely in the Amharic language, including many of the control sequences.

Alexander:1992:TTP

[Ale92] George Alexander. TEX and TEXtures: Page composition in a completely different vein. The Seybold Report on Desktop Publishing, 7(1): 3–7, September 4 1992. CODEN SRDFED. ISSN 0889-9762.

ANSI:1978:ANSe

[Ame78] American National Standards Institute. American National Standard for Bibliographic References. American National Standards Institute, 1430 Broadway, New York, NY 10018, USA, 1978. ???? pp.

Ament:2001:INB

[Ame01] Kurt Ament. Indexing:

a Nuts-and-bolts Guide for Technical Writers. William Andrew Publishing, Norwich, NY, USA, 2001. ISBN 0-8155-1481-6. x + 97 pp. LCCN T10.8 .A44 2001. URL

ftp://uiarchive.cso.
uiuc.edu/pub/etext/
gutenberg/;   http:/
/www.loc.gov/catdir/
description/wap041/
2001033646.html;   http:
//www.loc.gov/catdir/
toc/wap041/2001033646.
html;   http://www.
williamandrew.com/books.
asp?id=1481.

Department:1990:MPP

[AMS90] AMS Composition Services Department. From manuscript to published paper: How does it happen? part 2. Notices of the American Mathematical Society, 37(8):1027– 1028, October 1990. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Agostini:1985:IUF

[AMSV85] M. Agostini, V. Matano, M. Schaerf, and M. Vascotto. An Interactive User-Friendly TEX in VM/CMS Environment. In Lucarella [Luc85b], pages 117–132. ISBN 0-201-13399-7. LCCN Q224.E97 1985.

Andre:1983:AJM

[And83a] Jacques André, editor. Actes des Journ´ ees sur la Manipulation de Documents, Rennes 4–6 Mai 1983. INRIA/IRISA, Rennes, France, 1983.

Andre:1983:ABT

[And83b] Jacques André. Analytical bibliography on text manipulation. Technology and Science of Informatics, 1(5), 1983. ISSN 0264-7419.

Andre:1985:TIJ

[And85] Jacques André, editor. Typographie et Informatique, 21–25 janvier 1985. INRIA/IRISA, Rennes, France, 1985.

Andre:1992:ETN

[And92]  Jacques André.  En-

seigner la typographie numérique (English: Teaching digital typography), volume 1618 of Rapports de recherche. Institut National de Recherche en Informatique et en Automatique, Le Chesnay, France, 1992. 23 pp.

Andre:2005:PLT

[And05] Jacques André. Petites leçons de typographie. (French) [Small lessons of typography]. IRISA, Rennes, France, May 16, 2005. 50 pp. URL

http://www.irisa.fr/
faqtypo/lessons.pdf.

Revision of earlier versions of 1990 and 2003.

Anonymous:1982:CMS

[Ano82] Anonymous. The Chicago Manual of Style. University of Chicago Press, Chicago, IL, USA, 13th edition, 1982. ISBN 0-226-10390-0. ix + 737 pp. LCCN Z253 .U69 1982. UK£30.

Anonymous:1984:PMA

[Ano84] Anonymous. Publication Manual of the American Psychological Association. American Psychological Association, Washington, DC, USA, third edition, 1984. ISBN 0-912704-57-8. 208 pp. LCCN BF76.7 .P83 1983.

Anonymous:1987:CGP

[Ano87]  Anonymous.  Chicago

Guide to Preparing Electronic Manuscripts (For Authors and Publishers). University of Chicago Press, Chicago, IL, USA, 1987. ISBN 0-226-10392-7 (hardcover), 0-226-10393-5 (paperback). xi + 143 pp. LCCN Z286.E43 U54 1987. Short bibliography, examples, appendices, index.

Anonymous:1988:CDG

[Ano88] Anonymous. Composing documents with the generalized markup language. Technical Report S544-3421, IBM Corporation, San Jose, CA, USA, March 1988.

Anonymous:1990:JTB

[Ano90a] Anonymous. The Japanese TEXnicolor Book. ASCII Corporation, Tokyo, Japan, 1990. ISBN 4-7561-0405-3. Japanese; contents include chapters on the history of TEX and of the Japanese TEX extension, file formats (fonts and .dvi), and device drivers.

Anonymous:1990:NTF

[Ano90b] Anonymous. New TEX features from AMS. Notices of the American Mathematical Society, 37 (2):143–144, February 1990. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Anonymous:1991:B

[Ano91a] Anonymous. The Bluebook. Harvard Law Review Association, Cambridge, MA, USA, 15th edition, 1991. xvii + 343 pp. LCCN KF245 .U5 1991.

Anonymous:1991:PFA

[Ano91b] Anonymous, editor. Proceedings of the Fourth Annual Conference on Computers and the Writing Process. Univ. Sussex, Brighton, UK, March 1991.

Anonymous:1994:CCI

[Ano94] Anonymous, editor. Chinese computing: International conference — June 1994, Singapore, volume 4(1) of COMMUNICATIONS-COLIPS. 1994, Chinese and Oriental Languages Information Processing Society, 1994. ISSN 0218-7019.

Anonymous:1997:SW

[Ano97] Anonymous. Scientific WorkPlace 2.5. International Thomson Publishing GmbH, Königswintere Str. 418, 53227 Bonn, Germany, 1997. ISBN 3-87439-375-5. ???? pp. LCCN ???? DM 199,00. URL http:/

/www.itp.de/student/
index.html. Includes

CD-ROM.

Anonymous:2003:CMS

[Ano03] Anonymous. The Chicago Manual of Style. University of Chicago Press, Chicago, IL, USA, 15th edition, 2003. ISBN 0-226-10403-6. xvii + 956 pp. LCCN Z253 .U69 2003. US$55.00.

Anonymous:2010:CMS

[Ano10] Anonymous. The Chicago Manual of Style. University of Chicago Press, Chicago, IL, USA, 16th edition, 2010. ISBN 0-226-10420-6 (cloth). xvi + 1026 pp. LCCN Z253 .U69 2010.

Anonymous:2015:CEC

[Ano15] Anonymous. Correction: An Efficiency Comparison of Document Preparation Systems Used in Academic Research and Development. PLoS One, 10(4):e0125830, April 2015. CODEN POL-NCL. ISSN 1932-6203. See [KN14].

Anonymous:2017:CMS

[Ano17] Anonymous. The Chicago Manual of Style. University of Chicago Press, Chicago, IL, USA, 17th edition, 2017. ISBN 0-226-28705-X (hardcover). xvi + 1144 pp. LCCN Z253 .U69 2017.

Aiello:1983:SPD

[AP83] L. Aiello and S. Pavan. A system for the production of documents based on TEX (in italian). Rivista di Informatica (Milano), 13(3):261– 270, July 1983. ISSN 0390-668X.

Appelt:1985:HNE

[App85] Wolfgang Appelt. The hyphenation of non-English words with TEX. In Lucarella [Luc85b], pages 61–65. ISBN 0-201-13399-7. LCCN Q224.E97 1985.

Appelt:1988:TFP

[App88] Wolfgang Appelt. TEX für Fortgeschrittene: Programmiertechniken und Makropakete. Addison-Wesley, Reading, MA, USA, 1988. ISBN 3-89319-115-1. x + 179 pp. LCCN ????

Appelt:1990:TIG

[App90] Wolfgang Appelt. Die TEX-Installation in der GMD. Arbeitspapiere der gmd, Ges. f. Mathematik u. Datenverarbeitung, St. Augustin, Germany, 1990.

Appelt:1994:TF

[App94] Wolfgang Appelt. TEX für Fortgeschrittene. Addison-Wesley, Reading, MA, USA, second edition, 1994. ISBN 3-89319-695-1. xii + 183 pp.

Arnold:1972:IPH

[Arn72] Edmund C. Arnold. Ink on Paper 2: a Handbook of the Graphic Arts. Harper & Row, Publishers, New York, NY, USA, 1972. ISBN 0-06-010131-8. 374 pp. LCCN Z244 .A83 1972. Chapter bibliographies, sidebars, glossary, index, pp. 375.

Andre:1985:ETM

[AS85] Jacques André and Richard Southall. Experiments in teaching . In Lucarella [Luc85b], pages 141–153. ISBN 0-201-13399-7. LCCN Q224.E97 1985.

Antweiler:1989:TRI

[ASW89] Werner Antweiler, Andreas Strotmann, and Volker Winkelmann. A TEX-REDUCE interface. SIGSAM Bulletin, 23(2): 26–33, April 1989. CO-DEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Atkins:2004:APC

[Atk04] B. C. Atkins. Adaptive photo collection page layout. In ICIP ’04. 2004 International Conference on Image Processing, 24– 27 October 2004, Singapore, volume 5, pages 2897–2900. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 2004.

Bailey:1970:ASB

[Bai70] Herbert S. Bailey, Jr. The Art and Science of Book Publishing. Harper & Row, Publishers, New York, NY, USA, 1970. xii + 216 pp. LCCN Z278 .B15. Bibliography, tables, figures, appendix, index, pp. 217.

Bailey:1990:ASB

[Bai90] Herbert S. Bailey, Jr.

The Art and Science of Book Publishing. Ohio University Press, Athens, OH, USA, third edition, 1990. ISBN 0-8214-0970-0. xviii + 216 pp. LCCN Z278 .B15 1990.

Balda:1997:UNS

[Bal97] Miroslav Balda. Uzivatelske nadstavby sazeciho systemu L^{A}TEX 2.09. Zapadoceska univerzita, Plzen, Czech Republic, 1997. ISBN 80-7082-307-0. 119 pp. LCCN ????

Barzun:1986:WEP

[Bar86] Jacques Barzun. On Writing, Editing, and Publishing: Essays, Explicative and Hortatory. University of Chicago Press, Chicago, IL, USA, 1986. ISBN 0-226-03857-2 (hardcover), 0-226-03858-0 (paperback). xi + 148 pp. LCCN PN149 .B295 1986.

Barron:1989:WUS

[Bar89a]  D. Barron.  Why

use SGML? Electronic Publishing—Origination, Dissemination, and Design, 2(1):3–24, April 1989. CODEN EPODEU. ISSN 0894-3982.

Barwise:1989:TLA

[Bar89b]  Jon Barwise.  TEX,

L^{A}TEX, and AMS-TEX. Notices of the American Mathematical Society, 36(3):241–243, March 1989. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Bausum:2002:TRM

[Bau02] David Bausum. TEX reference manual. Kluwer Academic Publishers Group, Norwell, MA, USA, 2002. ISBN 0-7923-7673-0. xii + 384 pp. LCCN Z253.4.T47 B38 2002. US$99.00.

Botway:1985:LCS

[BB85] Lloyd Botway and Chris Biemesderfer. L^{A}TEX command summary (TEX 1.3, L^{A}TEX 2.08). Prepared for the Space Telescope Institute, Baltimore, MD; reprinted with permission and distributed by TUG, 1985.

Becker:1989:TAD

[BB89] Zeev Becker and Daniel Berry. triroff, an adaptation of the deviceindependent troff for formatting tri-directional text. Electronic Publishing—Origination, Dissemination, and Design, 2(3):119–142, October 1989. CODEN EPODEU. ISSN 0894-3982.

Badros:2001:CLA

[BBS01] Greg J. Badros, Alan Borning, and Peter J. Stuckey. The Cassowary linear arithmetic constraint solving algorithm. ACM Transactions on Computer-Human Interaction, 8(4):267–306, December 2001. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

Bouma:1988:DP

[BBv88] Louis G. Bouma, Jeroen Bruijning, and Johannes C. van Vliet. Document processing. In van der Veer and Mulder [vdVM88], pages 144– 169. ISBN 0-387-18901-7.

Beeton:2018:TBD

[BBW18] Barbara Beeton, Karl Berry, and David Walden. TEX: A branch of desktop publishing, part 1. IEEE Annals of the History of Computing, 40(3):78–93, July/ September 2018. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic). URL

                                        https://ieeexplore.
                                        ieee.org/document/8509554/

.

Beeton:2019:TBD

[BBW19] Barbara Beeton, Karl Berry, and David Walden. TEX: A branch in desktop publishing evolution, part 2. IEEE Annals of the History of Computing, 41(2):29–41, April 2019. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Bigelow:1983:DT

[BD83] Charles Bigelow and Donald Day. Digital typography. Scientific American, 249(2):106– 119, August 1983.

Barlow:1992:TC

[BE92] Geoff Barlow and Simon Eccles. Typesetting and Composition. Blueprint, London, UK, second edition, 1992. ISBN 0-948905-72-7. 274 pp.

Beckmann:1977:HP

[Bec77]  Petr Beckmann.  A

History of π. Golem Press, Boulder, CO, USA, fourth edition, 1977. ISBN 0-911762-18-3. 202 pp. LCCN QA484 .B4 1977. This book chronicles the story of the ultimate version number of TEX.

Becker:1987:AWP

[Bec87] Joseph D. Becker. Arabic word processing. Communications of the Association for Computing Machinery, 30(7):600– 610, July 1987. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Beccari:1991:LGS

[Bec91a] Claudio Beccari. L^{A}TEX —Guida a un sistema di editoria elettronica. Editore Ulrico Hoepli, Ulrico Hoepli Editore S.p.A., via Hoepli 5, 20121 Milano, Italy, 1991. ISBN 88-203-1931-4. ix + 399 pp. [From the author]: The book is the first and only one in Italian at this moment [18-Nov-1991]. Although it contains an introductory chapter for the beginner, its main purpose is to give the reader advanced information on L^{A}TEX (and partially on TEX) in order to prepare sophisticated macros, style files, and to make a clever use of fonts. Here is a reduced table of contents (translated to English): Chapter 1 - Introduction

- Introductory

L^{A}TEX

- General com-

mands for text mode - Mathematical commands

- Environ-

ments

- Figures and

tables - Macros

- Typical doc-

uments and style files

- Fonts

Appendix A - TEX Users Group

- Keyboards - Font tables - Hyphen-

ation patterns for Italian Bibliography Index The publisher received camera-ready pages that were prepared by the author with a 300-dpi laser printer. The page style dimensions were suitably enlarged while the photo plates were reduced so as to minimize the granulosity of the signs. In practice the book was prepared with 12pt option and reduced by the photografic ratio of 0.8 so that the final real font size is 9.6pt. The whole book was prepared with the standard Computer Roman fonts on purpose in order to show the potential user what he can get with a standard TEX/L^{A}TEX kit. The simple pictures were prepared with the “picture” environment. The cover and the title page were designed by a graphic expert of the publishing company.

Beccari:1991:TLS

[Bec91b]  Claudio Beccari.  La

tesi di laurea scientifica —Con note per la composizione in L^{A}TEX—disco MS-DOS da 5.25in contenente macroistruzioni di stile. Editore Ulrico Hoepli, Ulrico Hoepli Editore S.p.A., via Hoepli 5, 20121 Milano, Italy, 1991. ISBN 88-203-1932-2. vii + 116 pp. The book was typeset by the author and printed with a laser printer having a resolution of 300 dpi. The pages were then photographically reduced by a factor of 0.8 to the final size.. The fonts used are those of the standard TEX/L^{A}TEX set, that is Computer Modern, with a default normal size of 12pt that the photographic reduction made equivalent to approximately 10pt. The book is a set of instructions for the preparation of the thesis (Master and PhD levels) according to the Italian State Universities regulations; it deals mostly with literary style but it also includes extensive tables of scientific symbols and nomenclature and many references to Standards regulations. The final chapter deal with the typesetting of the thesis with a L^{A}TEX style, TOPTESI, whose source is recorded in a floppy disk that comes with the book.

Beebe:1987:TDD

[Bee87] Nelson H. F. Beebe. A TEXDVI Driver Family

— (Revision 2.07). University of Utah, Salt Lake City, UT 84112, USA, April 1987.

Beesley:2004:TDA

[Bee04] Kenneth R. Beesley. Typesetting the Deseret alphabet with L^{A}TEX and . In Syropoulos et al. [SBH^{+}04], pages 68–111. CO-DEN LNCSD9. ISBN 3-540-22801-2 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN Z253.3 I58 2004. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t3130.htm;
                                         http://www.springerlink.
                                        com/openurl.asp?genre=
                                        issue&issn=0302-9743&
                                        volume=3130;   http:
                                        //www.springerlink.
                                        com/openurl.asp?genre=
                                        volume&id=doi:10.1007/
                                        b99374.

Beebe:2013:BIT

[Bee13] Nelson H. F. Beebe.

BibTEX information and tutorial. Report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, February 16, 2013. 50 pp. URL https:

//www.math.utah.edu/
pub/bibnet/bibtex-info.
html.

Beitz:2000:TSS

[Bei00] Eric Beitz. TEXshade:

shading and labeling of multiple sequence alignments using L^{A}TEX2e. Bioinformatics, 16(2): 135–139, February 1, 2000. ISSN 1367-4803 (print), 1367-4811 (electronic).

Eddahibi:2004:AMD

[BEL04] Mostafa Banouni, Mohamed Elyaakoubi, and Azzeddine Lazrek. Arabic mathematical edocuments. In Syropoulos et al. [SBH^{+}04], pages 158–168. CO-DEN LNCSD9. ISBN 3-540-22801-2 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN Z253.3 I58 2004. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t3130.htm;
 http://www.springerlink.
com/openurl.asp?genre=

                                        issue&issn=0302-9743&
                                        volume=3130;   http:
                                        //www.springerlink.
                                        com/openurl.asp?genre=
                                        volume&id=doi:10.1007/
                                        b99374.

Bentley:1986:PPD

[Ben86a]  J. L. Bentley.  Pro-

gramming pearls: Document design. Communications of the Association for Computing Machinery, 29(9):832–839, September 1986. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Bentley:1986:LP

[Ben86b] Jon Bentley. Literate programming. Communications of the Association for Computing Machinery, 29(5):364– 369, 1986. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Bentley:1988:DLT

[Ben88] J. L. Bentley. dformat – A language for typesetting data formats. Comp. Sci. Tech. Rep. No. 142, AT&T Bell Laboratories, Murray Hill, NJ, USA, April 1988. 142.

Bennett:1991:BP

[Ben91] Paul A. Bennett, editor. Books and Printing. Biel, Savannah, Ga, 1991. ISBN 0-913720-72-0. 417 pp. LCCN Z116.A3 B58 1991. first published in 1951.

Berry:2006:TF

[Ber06]  D. M. Berry.  Text

formatting. In Keith Brown, Anne Anderson, et al., editors, Encyclopedia of Language & Linguistics, pages 607–621. Elsevier, Amsterdam, The Netherlands, second edition, 2006. ISBN 0-08-044299-4, 0-08-044854-2 (e-book). LCCN P29 .E48 2006. Fourteen volumes.

Bigelow:1986:DLI

[BH86] Charles Bigelow and Kris Holmes. The design of Lucida: an integrated family of types for electronic literacy. In van Vliet [vV86], pages 1– 17. ISBN 0-521-32592-7. LCCN QA76.9.T48 T481 1986.

Bianco:1992:TGU

[Bia92] Fiorella Bianco. TEX —Guida all’uso su personal computer (Guide for the use on personal computers). Pitagora Editrice, Via del legatore, 3, Bologna, Italy, 1992. ISBN 88-371-0583-5. ix + 157 pp. Prepared with Plain TEX and book macros written by the author. Final copy produced on a 300-dpi laser printer at magnification 1.2, then reduced for printing.

Bien:1988:PLT

[Bie88] Janusz S. Bień. Polish language and TEX. TEXline, 8:2, 1988.

Bien:1990:CTJ

[Bie90] Janusz S. Bień. Co to jest TEX? (what is TEX?). Wiadomości Matematyczne, 29(1): 132–156, 1990. In Polish.

Billawala:1989:MPS

[Bil89] Neenie Billawala. Metamarks: Preliminary studies for a Pandora’s box of shapes. Technical Report STAN-CS-89-1256, Stanford University, Department of Computer Science, Stanford, CA, USA, May 1989.

Binding:1993:ETS

[Bin93] Angelika Binding. Einsatz von TEX beim Springer-Verlag. Mitteilungen der Deutschen Mathematiker-Vereinigung, 19:19–22, 1993. ISSN 0942-5977, 0947-4471.

Bentley:1987:CPP

[BJK87] J. L. Bentley, L. W. Jelinski, and B. W. Kernighan. CHEM — A program for phototypesetting chemical structure diagrams. Computers and Chemistry, 11 (4):281–297, 1987. ISSN 0097-8485.

Bien:1985:TRC

[BK85] Janusz S. Bień and Hanna Kołodziejska. TEX for Riad Computers. In Lucarella [Luc85b], pages 133–140. ISBN 0-201-13399-7. LCCN Q224.E97 1985.

Bentley:1986:GLT

[BK86] Jon L. Bentley and Brian W. Kernighan. GRAP — a language for typesetting graphs. Communications of the Association for Computing Machinery, 29(8): 782–792, August 1986. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Bentley:1988:TPI

[BK88] J. L. Bentley and B. W.

Kernighan. Tools for printing indexes. Electronic Publishing—Origination, Dissemination, and Design, 1(1):3–18, April 1988. CODEN EPODEU. ISSN 0894-3982.

Bentley:1990:GLT

[BK90] J. L. Bentley and B. W.

Kernighan. GRAP — A language for typesetting graphs. tutorial and user manual. Unix Research System Papers, II: 109–146, 1990.

Bentley:1991:SAA

[BK91] Jon L. Bentley and Brian W. Kernighan. A system for algorithm animation. Computing Systems, 4(1):5–30, Winter 1991. CODEN CMSYE2. ISSN 0895-6340.

Bruggemann-Klein:1995:PR

[BKKW95] Anne Brüggemann-Klein, Rolf Klein, and Stefan Wohlfeil. Pagination reconsidered. Electronic Publishing—Origination, Dissemination, and Design, 8(2/3):139–152, June/September 1995. CODEN EPODEU. ISSN 0894-3982.

Bruggemann-Klein:1996:PR

[BKKW96] Anne Brüggemann-Klein, Rolf Klein, and Stefan Wohlfeil. Pagination reconsidered. Technical Report 205, Fernuniversität Hagen, Germany, 1996.

Bruggemann-Klein:1998:PCD

[BKKW98] Anne Brüggemann-Klein, Rolf Klein, and Stefan Wohlfeil. On the pagination of complex documents. Technical Report 234, Fernuniversität Hagen, Germany, July 27, 1998.

Bruggemann-Klein:2003:PCD

[BKKW03] Anne Brüggemann-Klein, Rolf Klein, and Stefan Wohlfeil. On the pagination of complex documents. In Klein et al. [KSW03], pages 49–68. CODEN LNCSD9. ISBN 3-540-00579-X (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76 .C5476 2003. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t2598.htm;
 http://www.springerlink.
com/content/978-3-540-
00579-7;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=
0302-9743&volume=2598.

Brueggeman-Klein:1989:DTN

[BKW89] A. Brüggeman-Klein and D. Wood. Drawing trees nicely with TEX. Electronic Publishing—Origination, Dissemination, and Design, 2(2):101– 115, July 1989. CODEN EPODEU. ISSN 0894-3982.

Black:1990:TDP

[Bla90] Alison Black. Typefaces for desktop publishing; a user guide. Architecture Design and Technology Press, Ltd., 140 Kensington Church St., London W84 BN, UK, 1990.

ISBN 1-85454-841-7. 106 pp. LCCN Z253.3.B53 1990.

Baecker:1989:HFT

[BM89] Ronald M. Baecker and Aaron Marcus. Human Factors and Typography for More Readable Programs. Addison-Wes-ley, Reading, MA, USA, 1989. ISBN 0-201-10745-7. xx + 348 pp. LCCN QA76.75 .B34 1990.

Bloomfield:1988:VVW

[BMJ88] R. Bloomfield, L. Marshall, and R. Jones, editors. VDM ’88: VDM, the Way Ahead: 2nd VDM-Europe Symposium, Dublin, Ireland, September, 1988: Proceedings. Springer-Ver-lag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1988. ISBN 0-387-50214-9 (New York), 3-540-50214-9 (Berlin). LCCN QA76.76.D47 V8541 1988.

Borceux:1990:LPT

[Bor90] Francis Borceux. L^{A}TEX: la perfection dans le traitement du texte. Éditions Ciaoco, Artel, Bruxelles, Belgium, 1990. ISBN 2-87085-194-4. LCCN ????

Borde:1992:TE

[Bor92a]  Arvind Borde.  TEX

by Example. Academic Press, New York, NY, USA, 1992. ISBN 0-12-117650-9. xiv + 169 pp. LCCN Z253.4.T47 B67 1992. This book was designed by the author, with some assistance from the production department of Academic Press, Boston. The covers were done by the publisher. It was typeset by the author using Plain TEX plus specially written macros. The code for all macros used for the book are reproduced in it so that the book itself contains all the information—page layout specifications, fonts used, and so on—about its own typesetting. Over half the book consists of TEX input and output shown side-by-side, providing an explicit example of how a part of the book was typeset. The Academic Press logo on page iii, the catalogingin-publication data on page iv, and the diagrams on page 33 were pasted in place by the production department of Academic Press. The diagrams were drawn using L^{A}TEX. Sample printouts of pages were made on various laser printers. When everything seemed satisfactory, the dvi files were copied onto IBM diskettes and were mailed to Academic Press, and from them to the American Mathematical Society. The final copy was produced by the A.M.S. using an APS micro-5 phototypesetter. All fonts used in the book are from the Computer Modern family. Parts of the covers of the book, inside and outside, were done in TEX by the author (the code for this is not included in the book). The outside cover was designed by Camille Pecoul of Academic Press and produced at Allison Associates in Boston. The font used for the title (on the covers) is Meridien. The background pattern was obtained by shooting a handmade marble paper as a black-and-white halftone and then overprinting with the Pantone Matching System (PMS). The placement of material on the cover was done by hand by Ms. Pecoul using a mechanical board. The drop shadow effect was done at Allison Associates.

Borkowski:1992:LPP

[Bor92b] Kazimierz M. Borkowski. L^{A}TEX. Profesjonalny skład publikacji. Adam Marszałek, Toruń, Poland, 1992. 88 pp.

Borde:1993:MTE

[Bor93] Arvind Borde. Mathematical TEX by Example. Academic Press, New York, NY, USA, 1993. ISBN 0-12-117645-2, 0-12-155940-8. xii + 352 pp. LCCN Z253.4.T47 B67 1993.

Borkowski:1997:LPS

[Bor97] Kazimierz Maria Borkowski.

L^{A}TEX — profesjonalny skład publikacji. Adam Marszałek, Toruń, Poland, 1997. ISBN 83-7174-023-9, 83-85263-64-0. 88 pp.

Borkowski:2001:LPS

[Bor01] Kazimierz Maria Borkowski.

L^{A}TEX: profesjonalny skład publikacji. Adam Marszałek, Toruń, Poland, 2001. ISBN 83-7322-004-6. 88 pp.

Boston:1986:STT

[Bos86] Bruce O. Boston. STET!

Tricks of the Trade for Writers and Editors. Editorial Experts, Inc., Alexandria, VA, USA, 1986. ISBN 0-935012-07-9. 310 pp. LCCN PN162 .S741 1986. US$15.95.

Bzyl:2004:MXC

[BP04] Włodzimierz Bzyl and Tomasz Przechlewski.

Migrating to XML: The case of the GUST Bulletin archive. In Syropoulos et al. [SBH^{+}04], pages 169–178. CO-DEN LNCSD9. ISBN 3-540-22801-2 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN Z253.3 I58 2004. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t3130.htm;
 http://www.springerlink.
com/openurl.asp?genre=
issue&issn=0302-9743&
volume=3130;   http:
//www.springerlink.
com/openurl.asp?genre=
volume&id=doi:10.1007/
b99374.

Beeton:2016:CMT

[BP16] Barbara Beeton and Richard Palais. Communication of mathematics with TEX. Visible Language, 50(2):40–51, August 2016. CODEN VSL-GAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL http://

visiblelanguage.herokuapp.
com/issue/202/article/
1362;   https://s3-
us-west-2.amazonaws.
com/visiblelanguage/
pdf/V50N2_2016_E.pdf.

Barron:1987:TPT

[BR87] David Barron and Mike Rees. Text processing and typesetting with UNIX.

International computer science series. Addison-Wesley, Reading, MA, USA, 1987. ISBN 0-201-14219-8. xiv + 447 pp. LCCN Z286.D47 B37 1987. US$24.95.

Borde:1993:UGL

[BR93] Arvind Borde and Tomas Rokicki. A User’s Guide for TEXHelp; The On-Line TEX Handbook. Academic Press, New York, NY, USA, 1993. ISBN 0-12-117640-1. v + 28 pp. This small booklet documents the TEXHelp program for IBM PC and compatible microcomputers.

Braswell:1989:IP

[Bra89] Frank Merritt Braswell.

Inside PostScript. Peachpit Press, Inc., 1085 Keith Avenue, Berkeley, CA 94708, USA, 1989. ISBN 0-938151-10-X. LCCN QA76.73.P67 B73 1989. US$37.50.

BSI:1950:BBR

[Bri50] British Standards Institution. BS 1629:1950: Bibliographical references. British Standards Institution, London, UK, February 24, 1950. 20

pp.  URL https:/
/shop.bsigroup.com/
ProductDetail?pid=000000000030307228.

BSI:1976:BRB

[Bri76] British Standards Institution. BS 1629:1976: Recommendations. Bibliographical references. British Standards Institution, London, UK, 1976. ISBN 0-580-09260-

7. ???? pp. URL https:

                                        //shop.bsigroup.com/
                                        ProductDetail?pid=000000000010033098.

BSI:1989:BRR

[Bri89] British Standards Institution. BS 1629:1989: Recommendation for references to published materials. British Standards Institution, London, UK, May 1989. 26 pp. URL https:

                                        //shop.bsigroup.com/
                                        ProductDetail?pid=000000000001551118.

Bringhurst:1992:ETS

[Bri92] Robert Bringhurst. The Elements of Typographic Style. Hartley & Marks Publishers, Point Roberts, WA, USA and Vancouver, BC, Canada, 1992. ISBN 0-88179-110-5 (hardback), 0-88179-033-8 (paperback). 254 pp. LCCN Z246 .B74 1992.

Bringhurst:1996:ETS

[Bri96] Robert Bringhurst. The Elements of Typographic Style. Hartley & Marks Publishers, Point Roberts, WA, USA and Vancouver, BC, Canada, second edition, 1996. ISBN 0-88179-133-4 (hardback), 0-88179-132-6 (paperback). 350 pp. LCCN Z246 .B74 1996.

Bringhurst:2004:ETS

[Bri04] Robert Bringhurst. The Elements of Typographic Style. Hartley & Marks Publishers, Point Roberts, WA, USA and Vancouver, BC, Canada, third edition, 2004. ISBN 0-88179-205-5, 0-88179-206-3 (paperback). 382 pp. LCCN Z246 .B74 2004. URL http:/

/www.loc.gov/catdir/
enhancements/fy1311/
2004053913-d.html.

BSI:2010:BII

[Bri10] British Standards Institution. BS ISO 690:2010: Information and documentation. Guidelines for bibliographic references and citations to information resources. British Standards Institution, London, UK, 2010. ISBN 0-580-56643-9. 50 pp. URL https:

//shop.bsigroup.com/
ProductDetail/?pid=
000000000030147089.

Bringhurst:2012:ETS

[Bri12] Robert Bringhurst. The Elements of Typographic Style 4.0. Hartley and Marks, Point Roberts, WA, USA, fourth edition, 2012. ISBN 0-88179-211-X (hardcover), 0-88179-212-8 (paperback). 398 pp. LCCN Z246 .B74 2012.

Brown:1989:IPT

[Bro89] Alex Brown. In Print: Text and Type. Watson-Guptill Publications, 1515 Broadway, New York, NY 10036, USA, 1989. ISBN 0-8230-2544-6. 191 pp. LCCN Z250.B868 1989.

Brooks:1991:LTV

[Bro91] Kenneth P. Brooks. Lilac: a two-view document editor. Computer, 24(6):7–19, June 1991. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Brown:1994:PTT

[Bro94] Vicki Brown, editor. Prime Time TEXcetera. Prime Time Freeware, 370 Altair Way, Suite 150, Sunnyvale, CA, USA, 1994. ISBN 1-881957-10-1. 96 pp. US$60.00. Includes CD-ROM.

Bryan:1988:SAG

[Bry88] Martin Bryan. SGML: An Author’s Guide to the Standard Generalized Markup Language. Addison-Wesley, Reading, MA, USA, 1988. ISBN 0-201-17535-5. xvii + 364 pp. LCCN QA76.73.S44 B79 1988. UK£16.95.

Bryan:1996:DTS

[Bry96]  Marvin Bryan.  Dig-

ital typography sourcebook. John Wiley and Sons, Inc., New York, NY, USA, 1996. ISBN 0-471-14811-3. xxiv + 384 pp. LCCN Z250.7 .B79 1996. Includes CD-ROM.

Burton:1983:SIC

[BS83] Sarah K. Burton and Douglas D. Short, editors. Sixth International Conference on Computers and the Humanities (North Carolina State University, Raleigh, NC, June 6–8, 1983). Computer Science Press, 11 Taft Court, Rockville, MD 20850, USA, 1983. ISBN 0-914894-96-X. LCCN AZ105 .I56 1983.

Badros:2001:CES

[BTM^{+}01] Greg J. Badros, Jojada J.

Tirtowidjojo, Kim Marriott, Bernd Meyer, Will Portnoy, and Alan Borning. A constraint extension to scalable vector graphics. In Vincent Y. Shen, editor, Proceedings of the 10th International Conference on World Wide Web: Hong Kong, Hong Kong, May 01– 05, 2001, pages 489–498. ACM Press, New York, NY, USA, 2001. ISBN 1-58113-348-0. LCCN TK5105.888 .I573 200.

Brandt:1989:CAC

[BU89] Josef Brandt and Ivar K. Ugi, editors. Computer Applications in Chemical Research and Education: a Collection of Papers Based on Lectures that were Presented at the VI-Ith International Conference on Computers in Chemical Research and Education at Garmisch-Partenkirchen, DE. Computeranwendung in der Chemie. Dr. Alfred Hüthing Verlag, Heidelberg, Germany, 1 edition, 1989. ISBN 3-7785-1804-6. xiii + 449 pp. LCCN QD39.3.E46 I56 1985. DM128.00.

Buerger:1990:LES

[Bue90] David J. Buerger. L^{A}TEX for Engineers and Scientists. McGraw-Hill, New York, NY, USA, 1990. ISBN 0-07-008845-4. xvi + 198 pp. LCCN Z253.4.L38 B84 1990.

Butcher:1981:CEC

[But81] Judith Butcher. Copy editing: the Cambridge handbook. Cambridge University Press, New York, second edition, 1981. ISBN 0-521-23868-4, 0-521-25638-0. LCCN PN162 .B86 1981. UK£22.

Buxton:1988:MLL

[Bux88] H. W. Buxton. Memoir of the life and labours of the late Charles Babbage Esq. F.R.S. Reprint series for the history of computing, vol. 13. MIT Press, Cambridge, MA, USA, 1988. ISBN 0-262-02269-9. xix + 401 pp. LCCN QA29.B2 B891 1988. US$50. From Computing Reviews: “Charles Babbage was a nineteenthcentury polymath who initially studied mathematics at Cambridge and subsequently pursued interests in fields as diverse as philosophy, political economy, and mechanical computing. While attempting to realize a machine for calculating and printing errorfree mathematical tables, he visited factories to learn metalworking techniques, invented a system of mechanical notation to represent the states of the machine, and devised a method for correct typesetting. As the editor of this series notes, the scope of Babbage’s work makes him a difficult subject to understand in our modern age of specialized scholarship.”.

Berry:2009:TPI

[BW09] Karl Berry and David Walden, editors. TEX People: Interviews from the World of TEX. TEX Users Group, Portland, OR, USA, 2009. ISBN 0-9824626-0-3. 312 (est.) pp. LCCN ???? US$36.00.

Bzyl:1993:TT

[Bzy93]  Włodek Bzyl.  TEX.

TEX?? broszura z konferencji członków GUSTu, Bachotek, 1993. In Polish.

GUTenberg:1988:CG

[Cah ] Cahiers GUTenberg, 1988– . Journal of Groupe des Utilisateurs de TEX Francophones (group of French-speaking TEX Users).

Casti:1991:VTP

[Cas91] John L. Casti. VTEX typesetting package. Notices of the American Mathematical Society, 38 (2):105–109, February 1991. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Casti:1993:BFT

[Cas93] John L. Casti. Bitmapfree TEX. Notices of the American Mathematical Society, 40(5):464–469, May/June 1993. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic). A review of Y&Y’s DVI-Windo and DVIPSONE, and comments on outline fonts for TEX.

Chan:1988:JAC

[CCL88] Kai Ming Chan, Kam Pui Chow, and Wai Man Lo. JACM—just another Chinese Metafont. In ICCPCOL ’88 [ICC88], pages 311–315. LCCN QA76.9.T48 I48 1988.

Ciancarini:2012:HQP

[CDFV12] Paolo Ciancarini, Angelo Di Iorio, Luca Furini, and Fabio Vitali. High-quality pagination for publishing. Software—Practice and Experience, 42(6):733–751, June 2012. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Canzii:1985:SSD

[CDMR85] G. Canzii, G. Degli Antoni, S. Mussi, and G. Rossi. SDDS: Scientific Document Delivery System. In Lucarella [Luc85b], pages 15– 25. ISBN 0-201-13399-7. LCCN Q224.E97 1985.

Cerofolini:1985:LCL

[Cer85] Luigi Cerofolini. A low cost laser beam printer controller. In Lucarella [Luc85b], pages 191– 195. ISBN 0-201-13399-7. LCCN Q224.E97 1985.

Chen:1987:AIP

[CH87] Pehong Chen and Michael A. Harrison. Automating index preparation. Technical Report 87/347, Computer Science Division, University of California, Berkeley, CA, USA, March 1987. This is an expanded version of [CH88].

Chen:1988:IPP

[CH88] Pehong Chen and Michael A. Harrison. Index preparation and processing. Software—Practice and Experience, 19(9):897–915, September 1988. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). The L^{A}TEX text of this paper is included in the makeindex software distribution.

Copeland:2001:CB

[CH01] Jeffrey Copeland and Jeffrey Haemer. Commonplace book. Server/ Workstation Expert, 12 (5):34–39, May 2001. Discusses book indexing with TEX and perl.

Cheswick:1990:PI

[Che90] Bill Cheswick. A permuted index for TEX and L^{A}TEX. Computing Science Technical Report 145, AT&T Bell Laboratories, Murray Hill, NJ, USA, February 1990. Also published by TUG as TEXniques No. 14.

Christian:1987:UTP

[Chr87] Kaare Christian. The UNIX text processing system. John Wiley and Sons, Inc., New York, NY, USA, 1987. ISBN 0-471-85581-2. xii + 250 pp. LCCN QA76.76.O63 C47 1987. US$24.95.

Church:1992:UWT

[Chu92] P. T. Church. Using workshops to teach TEX. Notices of the American Mathematical Society, 39(8):842, October 1992. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic). AMS-TEX input for two workshop handouts may be obtained by anonymous ftp from syr.edu in the directory

/math/papers.

Clarke:1987:BBS

[Cla87]  Mike Clarke.  Back

to basics: Simple but high-quality text pagination systems. In Earnshaw [Ear87], pages 203–211. ISBN 0-387-96527-0, 3-540-96527-0. LCCN Z286.E43 W67 1987. US$29.50.

Clark:1990:TAU

[Cla90] Malcolm Clark, editor. TEX: Applications, Uses, Methods. Proceedings, Third European TEX Conference, TEX88, Exeter, August 1988. Ellis Horwood, New York, NY, USA, 1990. ISBN 0-13-912296-6. LCCN Z253.4.T47 C45 1990. US$69.95.

Clark:1992:PTP

[Cla92] Malcolm Clark. A Plain TEX Primer. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 1992. ISBN 0-19-853784-0 (hardcover), 0-19-853724-7 (paperback). 481 pp. LCCN Z253.4.T47 C46 1992.

Comai:2012:MDM

[CM12] Sara Comai and Davide Mazza. A modeldriven methodology to the content layout problem in Web applications. ACM Transactions on the Web (TWEB), 6 (3):10:1–10:38, September 2012. CODEN ???? ISSN 1559-1131 (print), 1559-114X (electronic).

Cosar:1991:TIG

[CO91] A. Cosar and B. Ozguc.

A text, image, and graphics editor. Computers and Graphics, 15 (1):57–66, 1991. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Cho:2004:TCL

[CO04] Jin-Hwan Cho and Haruhiko Okumura. Typesetting CJK languages with Ω. In Syropoulos et al. [SBH^{+}04], pages 139–148. CO-DEN LNCSD9. ISBN 3-540-22801-2 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN Z253.3 I58 2004. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t3130.htm;
 http://www.springerlink.
com/openurl.asp?genre=
issue&issn=0302-9743&
volume=3130;   http:
//www.springerlink.
com/openurl.asp?genre=
volume&id=doi:10.1007/
b99374.

Collison:1962:IBM

[Col62] Robert Lewis Collison.

Indexing books: a manual of basic principles. De Graff, New York, NY, USA, 1962. 96 pp. LCCN Z695.9 .C64.

Collier:1991:CRD

[Col91] David Collier. Collier’s Rules for Desktop Design and Typography. Addison-Wesley, Reading, MA, USA, 1991. ISBN 0-201-54416-4. v + 135 pp. LCCN Z286.D47 C65 1991. US$14.95.

comp.text.tex:1989:CTT

[com ] comp.text.tex, 1989–. This is an unmoderated Usenet discussion list about TEX.

Cornell:1986:CHS

[Cor86]  Gary Cornell.  Chi-

writer: Horstman software group. The Mathematical Intelligencer, 8 (2):64–65, 1986.

Craig:1971:DT

[Cra71] James Craig. Designing with Type. Watson-Gup-till Publications, 1515 Broadway, New York, NY 10036, USA, 1971. ISBN 0-8230-1320-0. 175 pp. LCCN 70-159564.

Crawford:1987:CSW

[Cra87a] Walt Crawford. Common sense wordworking III: desktop publishing and desktop typesetting. Library Hi Tech, 5(1):43– 52, Spring 1987. CODEN LIHTD2. ISSN 0737-8831.

Crawford:1987:DTL

[Cra87b] Walt Crawford. Desktop typesetting and the LITA newsletter. Information Technology and Libraries, 6(2):89–96, June 1987. CODEN ITLBDC. ISSN 0730-9295.

Cross:1980:IB

[Cro80] Ruth Canedy Cross. Indexing books. A Word Guild guide. Word Guild, Cambridge, MA, USA, 1980. ISBN 0-9602524-1-X. xiii + 145 pp. LCCN Z695.9 .C76.

Chassell:1999:TGD

[CS99] Robert J. Chassell and Richard M. Stallman. TeXinfo: The GNU Documentation Format (for Texinfo version 4.0, 28 September 1999). Free Software Foundation, Inc., 51 Franklin Street, Fifth Floor, Boston, MA 02110-1301, USA, Tel:

(617) 876-3296, 1999. ISBN 1-882114-67-1. x + 244 pp. US$25.00.

Checkoway:2010:TOD

[CSR10a] Stephen Checkoway, Hovav Shacham, and Eric Rescorla. Are text-only data formats safe? Or, use this L^{A}TEX class file to pwn your computer. In LEET’10: 3rd Usenix Workshop on Large-Scale Exploits and Emergent Threats, San Jose, CA, April 27, 2010. USENIX, Berkeley, CA, USA, 2010. URL http:

//static.usenix.org/
events/leet10/tech/
full_papers/Checkoway.
pdf.

Checkoway:2010:DTL

[CSR10b] Stephen Checkoway, Hovav Shacham, and Eric Rescorla. Don’t take L^{A}TEX files from strangers. ;login: the USENIX Association newsletter, 35 (4):17–22, August 2010. CODEN LOGNEM. ISSN 1044-6397. URL https:/

/www.usenix.org/publications/
login/august-2010-volume-
35-number-4/dont-take-
latex-files-strangers;
 https://www.usenix.
org/system/files/login/
articles/73506-checkoway.
pdf.

Cai:2003:VBP

[CYW^{+}03] D. Cai, S. Yu, J. Rong Wen, W. Ying Ma, D. Cai, S. Yu, J. Rong Wen, and W. Ying Ma. 1VIPS: A vision-based page segmentation algorithm. Technical Report MSR-TR-203-79, Microsoft Corporation, Redmond, WA, USA, 2003.

Dair:1967:DT

[Dai67] Carl Dair. Design with Type. University of Toronto Press, Toronto, Ontario, Canada, 1967. ISBN 0-8020-1426-7 (hardcover), 0-8020-6519-8 (paperback). vi + 162 pp. LCCN Z246 D35 1967.

DANTE:1989:TK

[DAN ] Die TEXnische Komödie, 1989–. Journal of DANTE (Deutschsprachige Anwendervereinigung TEX e.V.); Group of German-speaking TEX Users.

Darling:1993:F

[Dar93] R. W. R. Darling.

FrameMaker 3.0. Notices of the American Mathematical Society, 40 (2):150–152, February 1993. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Datta:2017:LHP

[Dat17]  Dilip Datta.  L^{A}TEX

in 24 hours: a practical guide for scientific writing. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2017. ISBN 3-319-47830-3, 3-319-47831-1 (e-book). xxv + 296 pp. LCCN T11 .D388 2017; Z52.4. URL http:

//link.springer.com/
10.1007/978-3-319-47831-

Day:1983:TMM

[Day83] R. A. Day. Typesetting mathematics on multiaccess systems. Software—Practice and Experience, 13(2):131–138, February 1983. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

deBruin:1988:PML

[dBvLV88] R. de Bruin, C. G. van der Laan, J. R. Luyten, and H. F. Vogt. Publiceren met L^{A}TEX. Number 19. Centrum voor Wiskunde en Informatica, P. O. Box 4079, 1009 AB Amsterdam, The Netherlands, 1988. ISBN 90-6196-357-5. 196 pp. LCCN Z253.4.L38 .P97 1988.

DeRose:1994:MHW

[DD94] Steven J. DeRose and David G. Durand. Making Hypermedia Work. Kluwer Academic Publishers Group, Norwell, MA, USA, 1994. ISBN 0-7923-9432-1. xxii + 384 pp. LCCN QA76.76.I59 D47 1994. UK£51.00, US$68.00.

DeLaBeaujardiere:1988:WDI

[De 88] Jean-M. De La Beaujardiere. Well-established document interchange formats. In van Vliet [vV88], pages 83–94.

ISBN 0-521-36294-6. LCCN Z286.E43 I571 1988. US$47.50.

Dreyfus:1989:AXF

[DE89] John Dreyfus and Knut Erichson, editors. ABC-XYZapf: Fifty Years in Alphabet Design: Professional and personal contributions selected for Hermann Zapf. The Wynkyn de Worde Society and Bund Deutscher Buchkünstler, London, UK and Offenbach, Germany, 1989. ISBN 3-922715-35-4. 251 pp. LCCN Z250.A2 Z358 1989.

Deach:1992:OFH

[Dea92] Stephen Deach. Outline font hints and rasterization: a technology primer. The Seybold Report on Desktop Publishing, 6(7):21–31, March 9 1992. CODEN SRDFED. ISSN 0889-9762. This paper is followed by a dissent from Adobe (Terry O’Donnell), and a response from the author (p. 32). A sidebar (historical note, p. 25) contains a short description of and cites as a forerunner of the hinting mechanism used in TrueType.

Demmig:1989:PBP

[Dem89] S. Demmig. Printing books on the PC or the fantastic possibilities of TEX. Mikrocomputer Zeitschrift, 7:60–66, July 1989. CODEN MD-MZDL. ISSN 0720-4442.

Denton:1992:GVC

[Den92] Craig Denton. Graphics for visual communication. William C. Brown Company Publishers, Dubuque, IA, USA, 1992. ISBN 0-697-08540-6. xv + 383 pp. LCCN P93.5.D46 1992.

Desarmenien:1984:HRT

[Dés84a] Jacques Désarménien. How to run TEX in French. Technical Report STAN-CS-84-1013, Stanford University, Stanford, CA, USA, August 1984.

Desarmenien:1984:DOM

[Dés84b] Jacques Désarménien. La division par ordinateur des mots français avec le logiciel TEX. In How to run TEX in French, pages 19–42. Stanford University, Stanford, CA, USA, 1984.

Desarmenien:1985:UTF

[Dés85] Jacques Désarménien. The Use of TEX in French: Hyphenation and Typography. In Lucarella [Luc85b], pages 41–59. ISBN 0-201-13399-7. LCCN Q224.E97 1985.

Desarmenien:1986:TSD

[Dés86] Jacques Désarménien, editor. TEX for Scientific Documentation. Second European Conference. Strasbourg, France, June 19–21, 1986. Proceedings, number 236 in Lecture Notes in Computer Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1986. ISBN 3-540-16807-9, 0-387-16807-9 (paperback). LCCN Z253.4.T47 T49 1986.

Desgraupes:1999:MGP

[Des99]  Bernard  Desgraupes.

Metafont: Guide pratique. Vuibert, Paris, France, 1999. ISBN 2-7117-8642-0. 267 pp. LCCN ????

Detig:1997:LW

[Det97] Christine Detig. Der L^{A}TEX 2ε-Wegweiser. International Thomson Publishing GmbH, Königswintere Str. 418, 53227 Bonn, Germany, 1997. ISBN 3-8266-0256-0. xiv + 236 pp. DM 29,80.

Dol:1999:W

[DF99] Wietse Dol and Erik Frambach. 4TEX for Windows. Dutch-language oriented TEX Users Group (NTG), Amsterdam, The Netherlands, June 1999. ISBN 90-76669-01-5. xx + 531 pp. LCCN ???? US$25.

Dol:1994:TG

[DFKvdV94] Wietse Dol, Erik Frambach, Siep Kroonenberg, and Maarten H. van der Vlerk. TEX and 4TEX Guide. Nederlandstalige TEX Gebruikersgroep (Dutch-speaking TEX Users Group), P.O. Box 394, 1740 AJ Schagen, The Netherlands, 1994. US$35.00. The CD-ROM provides a complete runnable installation of TEX and - for the IBM PC.

Delpech:1997:TDL

[DFM97] Jean-Pierre Delpech, Marc-Andre Figueres, and Nicole Mari. Techniques du L^{A}TEX. Editions Eyrolles, Paris, France, 1997. ISBN 2-212-02609-9. 128 pp. LCCN ????

DiIorio:2007:PBS

[Di 07] A. Di Iorio. Patternbased segmentation of digital documents: model and implementation. Ph.D. thesis, Department of Computer Science, University of Bologna, Bologna, Italy, 2007. Available as Technical Report UBLCS-2007-05. Abstract published in the ACM SIGWEB Newsletter, March 2008, 3:1– 3:2).

Dickey:1987:TAU

[Dic87] Leroy J. Dickey. Typesetting APL using TROFF. In Haliburton [Hal87], pages 494–497. ISBN 0-89791-226-8. LCCN ???? US$37.00. APL quote quad, v. 17, no. 4.

Diller:1993:LLL

[Dil93] Antoni Diller. L^{A}TEX Line by Line: Tips and Techniques for Document Processing. Wiley Professional Computing. John Wiley and Sons, Inc., New York, NY, USA, 1993. ISBN 0-471-93471-2 (softcover). xiii + 291 pp. LCCN Z253.4.L38 D55 1993. US$29.95. Optionally accompanied by disk with examples, ISBN 0-471-93797-5.

Diller:1999:LLL

[Dil99] Antoni Diller. L^{A}TEX Line by Line: Tips and Techniques for Document Processing. Wiley, Chichester, UK, second edition, 1999. ISBN 0-471-97918-X (softcover). xv + 311 pp. LCCN Z253.4.L38 D55 1999.

Dietsche:1994:LLS

[DL94] Luzia Dietsche and Joachim Lammarsch. L^{A}TEX zum Loslegen: ein Soforthelfer für den Alltag [English: L^{A}TEX, ready to start: Instant help for everyday]. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1994. ISBN 3-540-56545-0. x + 116 pp. DM39.00.

Duan:1989:IPS

[DM89] G. Duan and R. A. Morris. The importance of phase in the spectra of digital type. Electronic Publishing—Origination, Dissemination, and Design, 2(1):47–60, April 1989. CODEN EPODEU. ISSN 0894-3982.

DeMillo:1991:USA

[DM91] Richard A. DeMillo and Aditya P. Mathur. On the use of software artifacts to evaluate the effectiveness of mutation analysis for detecting errors in production software. Technical Report SERC-TR-92-P, Software Engineering Research Center, Purdue University, W. Lafayette, IN 47907, USA, 1991.

Doob:1989:TTA

[Doo89]  Michael Doob.  TEX

and typesetting—an author’s view. Notices of the American Mathematical Society, 36(9):

1203–1204, November 1989. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Doob:1990:TSC

[Doo90] Michael Doob. TEX and the single CPU. Notices of the American Mathematical Society, 37(3):270–273, March 1990. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Doob:1993:TSS

[Doo93] Michael Doob. TEX:

Starting from Square 1. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1993. ISBN 0-387-56441-1. 114 pp. LCCN Z253.4.T47 D66 1993. US$25.00.

Doob:2000:PVS

[Doo00] Michael Doob. To proto vima sto TEX. Paratiritis Editions, Thessaloniki, Greece, 2000. ISBN 960-374-081-0. viii + 144 pp. Translation into Greek by Dimitrios Filippou.

Dooley:2002:PIM

[Doo02] Samuel S. Dooley. Programming the IBM MathML expression editor for interactive mathematical applications. In Mathematical software (Beijing, 2002), pages 421–431. World Scientific, Singapore, 2002.

Drucker:1995:ALL

[Dru95] Johanna Drucker. The Alphabetic Labyrinth: The Letters in History and Imagination. W. W. Norton & Co., New York, NY, USA, 1995. ISBN 0-500-01608-9. 328 pp. LCCN P211 .D78 1995 AAL. US$45.00.

Dallas:1983:IWU

[DS83] Ian Neil Dallas and Edward Brian Spratt, editors. Ring technology Local Area Networks: Proceedings of the IFIP WG6.4/University of Kent Workshop on Ring Technology Based Local Area Networks, Kent, UK, 28–30 September 1983. North-Holland Publishing Company, Amsterdam, The Netherlands, 1983. ISBN 0-444-86852-6. LCCN TK5105.5 .I3427 1983.

daSilva:2005:SAR

[dSdOM^{+}05] Ana Cristina B. da Silva, Joao B. S. de Oliveira, Fernando T. M. Mano, Thiago B. Silva, Leonardo L. Meirelles, Felipe R. Meneguzzi, and Fabio Giannetti. Support for arbitrary regions in XSL-FO. In King [Kin05], pages 64–73. ISBN 1-59593-240-2. LCCN HF5736 .A27 2005. ACM order number 605054.

Dyl:1994:KTS

[Dyl94]  Dariusz Dyl.  Kom-

puterowy skład tekstu w systemie L^{A}TEX. Przewodnik po komendach i środowiskach wraz z przykładami. Adam Marszałek, Toruń, Poland,

1. edition, 1994. składano w systemie L^{A}TEX (styl book11pt) przy użyciu programu VectorTEX ver. 2.24.

Eckersley:1994:GTT

[EAE^{+}94] Richard Eckersley, Richard Angstadt, Charles M. Ellertson, Richard Hendel, Naomi B. Pascal, and Anita Walker Scott. Glossary of Typesetting Terms. Chicago Guides to Writing, Editing, and Publishing. University of Chicago Press, Chicago, IL, USA, 1994. ISBN 0-226-18371-8. xi + 169 pp. US$20.00.

Earnshaw:1987:WPS

[Ear87] Rae A. Earnshaw, editor.

Workstations and Publication Systems. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1987. ISBN 0-387-96527-0, 3-540-96527-0. viii + 229 pp. LCCN Z286.E43 W67 1987. US$29.50.

Ebel:1995:UWM

[Ebe95] H. F. Ebel. The ultimate weapon: More about reference formats. The EASE Journal: European Science Editing, (54):21–22, February 1995. ISSN 0950-1835.

Evgrafov:1993:TRN

[EE93] Marat Evgrafov and Leonid Evgrafov. TEX: Rukovodstvo po naboru i redaktirovaniiu mathematicheskikh textov [English: Manual on typing and editing of mathematical texts]. Phismathlit, Moscow, Russia, 1993. ISBN 5-02-015116-5. 80 pp. The first Russian book about TEX that was written by Russians. First press run: 7000 copies.

Eickel:1990:LLS

[Eic90] Jürgen Eickel. Logical and layout structures of documents. In Nadrchal [Nad90], pages 201–208. Invited paper.

Eijkhout:1992:TTT

[Eij92] Victor Eijkhout. TEX by Topic, a TEXnician’s Reference. Addison-Wes-ley, Reading, MA, USA, 1992. ISBN 0-201-56882-9. vii + 307 pp. LCCN Z253.4.T47 E38 1992.

URL http://eijkhout.
net/texbytopic/.

Eijkhout:2012:CST

[Eij12] Victor Eijkhout. The computer science of TEX and L^{A}TEX; based on CS 594, fall 2004, University of Tennessee. Report, Texas Advanced Computing Center, The University of Texas at Austin, Austin, TX, USA, 2012. ISBN 1-105-41591-0. 294

pp.  URL http:/
/pages.tacc.utexas.
edu/~eijkhout/Articles/
TeXLaTeXcourse.pdf.

Eddahibi:2004:DAM

[ELS04] Mustapha Eddahibi, Azzeddine Lazrek, and Khalid Sami. Dynamic Arabic mathematical fonts. In Syropoulos et al. [SBH^{+}04], pages 149–

157. CODEN LNCSD9. ISBN 3-540-22801-2 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN Z253.3 I58 2004. URL

http://link.springer-
ny.com/link/service/
series/0558/tocs/t3130.
htm;   http://www.springerlink.
com/openurl.asp?genre=
issue&issn=0302-9743&
volume=3130;   http:
//www.springerlink.
com/openurl.asp?genre=

                                        volume&id=doi:10.1007/
                                        b99374.

English:1987:UMO

[Eng87] Paul M. English. Using for original font design. August 1987.

Emerson:1987:TTU

[EP87] Sandra L. Emerson and Karen Paulsell. troff Typesetting for UNIX systems. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 1987. ISBN 0-13-930959-4. xx + 359 pp. LCCN Z253.4.U53 E45 1987. US$29.95.

Eijkhout:1990:WT

[EP90] Victor Eijkhout and Nico Poppelier. Wat is TEX. TWIOscoop, 8(2):44–48, 1990.

EPODD:1988:EPO

[EPO ] Electronic publishing— origination, dissemination, and design. Wiley, 1988–. ISSN 0894-3982. This journal accepts papers in TEX form.

Ehrenfeucht:1994:CFT

[EtPR94] A. Ehrenfeucht, P. ten Pas, and G. Rozenberg. Context-free text grammars. Acta Informatica, 31(2):161–206, ?? 1994. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic). From Joachim Schrod: This paper describes “the basic theory behind text grammars needed for document structure specification”. Fateman:1987:TOM

[Fat87] Richard J. Fateman. TEX output from MACSYMAlike systems. SIGSAM Bulletin, 21(4):1–5, November 1987. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Anonymous:1989:CT

[Féd89] Fédération C. G. C.

de la Communication. Code Typographique— Choix de règles à l’usage des auteurs et professionels du livre, seizième edition, 1989. This book is a compendium of rules for French typography, and for the typesetting of several foreign languages in French texts. It includes tables of diacritics for most European languages, and hyphenation rules for English, German, Greek, Italian, Latin, Portuguese, and Spanish.

Feiner:1988:GBA

[Fei88] S. K. Feiner. A gridbased approach to automating display layout. In Graphics Interface ’88, 6–10 June, Edmonton, Alberta: proceedings = Graphics Interface ’88, 6–10 juin, Edmonton, Alberta; comptes rendus, pages 192–197. Canadian Information Processing Society, Toronto, ON, Canada, 1988. ISBN ????

Filgueiras:1998:ISM

[Fil98] Miguel Filgueiras. Implementing a symbolic music processing system. Software—Practice and Experience, 28(5):493– 512, May 1998. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http:

//www3.interscience.
wiley.com/cgi-bin/abstract?
ID=1793;   http://www3.
interscience.wiley.
com/cgi-bin/fulltext?
ID=1793&PLACEBO=IE.
pdf.

Filippou:2004:HPA

[Fil04] Dimitrios Filippou. Hyphenation patterns for ancient and modern Greek. In Syropoulos et al. [SBH^{+}04], pages 59–67. CO-DEN LNCSD9. ISBN 3-540-22801-2 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN Z253.3 I58 2004. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t3130.htm;
 http://www.springerlink.

com/openurl.asp?genre=
issue&issn=0302-9743&
volume=3130;   http:
//www.springerlink.
com/openurl.asp?genre=
volume&id=doi:10.1007/
b99374.

Fine:2000:LBP

[Fin00] Jonathan Fine. Line breaking and page breaking. TUGboat, 21 (3):210–221, September 2000. ISSN 0896-3207.

Finston:2003:URM

[Fin03] Laurence D. Finston.

3DLDF user and reference manual: 3-dimensional drawing with - output. ????, ????, 2003. ix + 269

pp.  URL http://
dante.ctan.org/CTAN/
graphics/3DLDF/3DLDF.

pdf. Manual edition 1.0 for 3DLDF version 1.0 September 2003.

Finston:2004:URM

[Fin04] Laurence D. Finston.

3DLDF user and reference manual: 3-dimensional drawing with - output. ????, ????, 2004. x + 296

pp.  URL http://
dante.ctan.org/CTAN/
graphics/3DLDF/3DLDF.

pdf. Manual edition 1.1.5.1 for 3DLDF version 1.1.5.1 January

Firmage:1993:AA

[Fir93] Richard A. Firmage. The Alphabet Abecedarium: Some Notes on Letters. David R. Godine, Publisher, Boston, MA, USA, 1993. ISBN 0-87923-987-5. xi + 307 pp. LCCN P211 .F6 1993.

Fisher:1991:IAI

[Fis91] A. J. Fisher. Incremental algorithms for interactive text formatting. The Journal of Systems and Software, 16(1):3– 16, September 1991. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Fuchs:1982:OFC

[FK82] David R. Fuchs and Donald E. Knuth. Optimal font caching. Technical Report STAN-CS-82-901, Stanford University, Stanford, CA, USA, 1982.

Fuchs:1985:OPF

[FK85] D. R. Fuchs and D. E. Knuth. Optimal prepaging and font caching. ACM Transactions on Programming Languages and Systems, 7(1):62–79, January 1985. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Furuta:1985:RTT

[FM85] Richard Furuta and Pierre A. MacKay. Two TEX implementations for the IBM PC. Dr. Dobb’s Journal of Software Tools, 10(9):80–89, 91, September 1985. CO-DEN DDJOEB. ISSN 1044-789X.

Felici:1987:DPS

[FN87] James Felici and Ted Nace. Desktop publishing skills: a primer for typesetting with computer and laser printer. Addison-Wesley, Reading, MA, USA, August 1987. ISBN 0-201-11537-9. vii + 180 pp. LCCN Z286.D47 F44 1987. UK£20, US$19.95.

Forsley:1989:RFC

[For89] Lawrence P. Forsley, editor. 1989 Rochester Forth Conference. Industrial Automation: June 20–24, 1989, University of Rochester. Institute for Applied FORTH Research, Rochester, NY, USA, 1989. ISBN 0-914593-09-9. LCCN TJ212.2.A15 1989.

Foxley:1987:MLT

[Fox87] Eric Foxley. Music— a language for typesetting music scores. Software—Practice and Experience, 17(8):485–502, August 1987. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Fox:1997:QLO

[Fox97] Jim Fox. Q&AIX: L^{A}TEX, the old and the new. Sun-Expert Magazine, 8(11): 72, 74, November 1997. ISSN 1053-9239.

Fox:1998:QLL

[Fox98] Jim Fox. Q&AIX: Life with L^{A}TEX. SunExpert Magazine, 9(4):53– 54, 56, April 1998. ISSN 1053-9239.

Foata:1984:SLT

[FPR84] Dominique Foata, Jean-Jacques Pansiot, and Yves Roy. STRATEC— un logiciel traitement de textes mathématiques en amont de TEX (ver.0.93). Laboratoire de Typographie Informatique, Université Louis Pasteur, Strasbourg, France, 1984.

Foata:1985:SRK

[FPR85] Dominique Foata, Jean-Jacques Pansiot, and Yves Roy. STRATEC and a Rationalized Keyboard for Inputting TEX. In Lucarella [Luc85b], pages 105–116. ISBN 0-201-13399-7. LCCN Q224.E97 1985.

Flipo:1992:HPA

[FS92] Daniel Flipo and Laurent Siebenmann. Hyphenation in presence of accents and diacritics. an easy and low-cost solution. In Zlatuška [Zla92], pages 87–96. ISBN 80-210-0480-0.

Furuta:1982:DFSa

[FSS82a] Richard Furuta, Jeffrey Scofield, and Alan Shaw. Document formatting systems: survey, concepts, and issues. ACM Computing Surveys, 14(3):417–472, September 1982.

Furuta:1982:DFSb

[FSS82b] Richard Furuta, Jeffrey Scofield, and Alan Shaw. Document formatting systems: survey, concepts, and issues. In Nievergelt et al. [NCNS82], chapter ??, pages 133–220. ISBN 0-444-86493-8. LCCN Z244 .D63 1982. US$46.50.

Fujita:1993:CBG

[Fuj93] Shinsaku Fujita. L^{A}TEX for Chemists and Biochemists. A Guide for Preparing Papers with Personal Computers [in Japanese]. Tokyo Kagaku Dozin, Tokyo, Japan, 1993. ISBN 4-8079-0386-1. xiv + 305 pp. Includes diskette with macros and styles for about thirty scientific journals.

Fujita:1994:TSF

[Fuj94] Shinsaku Fujita. Typesetting structural formulae with the text formatter TEX/L^{A}TEX. Computers in Chemistry, 18(2): 109–116, 1994.

Furuta:1990:PIC

[Fur90] R. Furuta, editor. Proceedings of the International Conference on Electronic Publishing, Document Manipulation & Typography, Gaithersburg, Maryland, September 1990. Cambridge University Press, New York, 1990. ISBN 0-521-40246-8. LCCN Z 286 E43 I58 1990.

Feijen:1990:BOB

[FvGGM90] W. H. J. Feijen, A. J. M. van Gasteren, David Gries, and J. Misra, editors. Beauty is our Business: a Birthday Salute to Edsger W. Dijkstra. Springer-Ver-lag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1990. ISBN 0-387-97299-4, 3-540-97299-4, 1-4612-8792-8 (print), 1-4612-4476-5 (online). ISSN 0172-603X. xix + 453 pp. LCCN QA76 .B326 1990. URL http:

//www.zentralblatt-
math.org/zmath/en/search/
?an=0718.68004. Con-

tains important treatment of accurate binaryto-decimal conversion [Gri90, Knu90b].

Gallego:1992:LPP

[Gal92] Francisco Ortegón Gallego. L^{A}TEX, primeros pasos. Manuales de Informática Masson. Masson, S.A., Barcelona, Spain, 1992. ISBN 84-311-0649-2. xix + 168 pp.

Garfield:1968:NVA

[Gar68] E. Garfield. Numerical versus alphabetic order for cited references. New Scientist, 39(614): 565–566, September 12, 1968. CODEN NWS-CAL. ISSN 0262-4079 (print), 1364-8500 (electronic).

Garfield:1974:UEP

[Gar74a] E. Garfield. Uniformity of editorial policy on titles in citations will aid referees, librarians and authors. Essays of an Information Scientist, 2: 229–230, 1974.

Garfield:1974:WII

[Gar74b] E. Garfield. Why initials instead of first names in ISI’s indexes? Essays of an Information Scientist, 2:323–324, 1974.

Gaudeul:2007:DOS

[Gau07] Alex Gaudeul. Do open source developers respond to competition? The L^{A}TEX case study. Review of Network Economics, 6(2):239–263, June 2007. CODEN ???? ISSN 2194-5993 (print), 1446-9022 (electronic). URL https:

//econpapers.repec.
org/article/bpjrneart/
v_3a6_3ay_3a2007_3ai_
3a2_3an_3a9.htm;   https:
//www.degruyter.com/
view/j/rne.2007.6.issue-
2/rne.2007.6.2.1119/
rne.2007.6.2.1119.xml.

GilmourBryson:1982:TPC

[GB82] A. Gilmour-Bryson. Text preparation for computer typesetting. ALLC Bull, 10(2):40–47, 1982. CO-DEN ALLCB5. ISSN 0951-1474.

Ghosh:1983:FAL

[GB83] Pijush K. Ghosh and Charles A. Bigelow. A Formal Approach to Lettershape Description for Type Design. Technical Report STAN-CS-83-966, Stanford University, Stanford, CA, USA, 1983.

Griffee:1988:ITF

[GC88] Alan W. Griffee and Cherrelyn A. Casey. An introduction to typographic fonts and digital font resources. IBM Systems Journal, 27(2):206, 1988. CODEN IBMSA7. ISSN 0018-8670.

Glasman-Deal:2010:SRW

[GD10] Hilary Glasman-Deal.

Science research writing for non-native speakers of English. Imperial College Press, London, UK, 2010. ISBN 1-84816-309-6, 1-84816-310-X (paperback). xi + 257 pp. LCCN PE1475 .G57

2010. URL http://www.
worldscientific.com/
worldscibooks/10.1142/
p605.

Gehani:1986:DFT

[Geh86] Narain Gehani. Document formatting and typesetting on the UNIX system. Silicon Press, 25 Beverly Road, Summit, NJ 07901, USA, 1986. ISBN 0-9615336-0-9. xv + 364 pp. LCCN Z52.5.U54 G43 1986. US$33.95.

Gehani:1987:DFT

[Geh87] Narain Gehani. Document Formatting and Typesetting on the UNIX System. Silicon Press, 25 Beverly Road, Summit, NJ 07901, USA, second edition, 1987. ISBN 0-13-938325-5, 0-9615336-2-5. xv + 377 pp. LCCN Z52.5.U54G43 1987. US$40.90.

Gella:1995:CTL

[Gel95]  Kamal  VenkataSatya

Gella. A compiler to translate L^{A}TEX document to hypertext. Thesis (m.s.), University of South Carolina, Columbia, SC, USA, 1995. iv + 40 pp.

Guoan:1982:USM

[GH82] Gu Guoan and John Hobby. Using String Matching to Compress Chinese Characters. Report STAN-CS-82-914, Stanford University, Department of Computer Science, Stanford, CA, USA, 1982.

Griffiths:1997:LL

[GH97] David F. (Francis) Griffiths and Desmond J. Higham. Learning L^{A}TEX. Society for Industrial and Applied Mathematics, Philadelphia, PA, USA, 1997. ISBN 0-89871-383-8, 0-89871-956-9 (ebook). x + 84 pp. LCCN Z253.4.L38G75 1997.

Griffiths:2016:LL

[GH16] David Francis. Griffiths and Desmond J. Higham.

Learning L^{A}TEX. Society for Industrial and Applied Mathematics, Philadelphia, PA, USA, second edition, 2016. ISBN 1-61197-441-0 (paperback). x + 103 pp. LCCN Z253.4.L38 G75 2016.

Griffiths:2020:WLL

[GH20] David F. (Francis) Griffiths and Desmond J. Higham. Writing Learning L^{A}TEX. SIAM News, ??(??):??, October 1, 2020. ISSN 0036-1437.

URL https://sinews.
siam.org/Details-Page/
writing-learning-latex.

Ghosh:1983:ATD

[Gho83] Pijush K. Ghosh. An Approach to Type Design and Text Composition in Indian Scripts (with a preface by donald E. knuth). Technical Report STAN-CS-83-965, Stanford University, Stanford, CA, USA, 1983.

Ghosh:1984:BDI

[Gho84] P. K. Ghosh. Basic design issues in a multilingual type font design and typesetting workstation. J. Inst. Electron. Telecommun. Eng, 30(6):196–202, November 1984. CODEN JI-ETAU. ISSN 0377-2063 (print), 0974-780X (electronic).

Gibaldi:1996:MHW

[Gib96] Joseph Gibaldi. MLA Handbook for Writers of Research Papers. Modern Language Association of America, New York, NY, USA, fourth edition, 1996. ISBN 0-87352-565-5. xviii + 293 pp. LCCN LB2369.G53 1995. US$16.75.

Gilchrist:1989:TTP

[Gil89] Martin Gilchrist. TEX and typesetting—A publisher’s view. Notices of the American Mathematical Society, 36(9): 1199–1201, November 1989. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Ginsparg:1994:FST

[Gin94] Paul Ginsparg. First steps towards electronic research communication. Computers in Physics, 8(4):390–??, July 1994. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL

                                        https://aip.scitation.
                                        org/doi/10.1063/1.4823313.

Galvez:2014:IIG

[GKK14] Akemi Gálvez, Kiyoshi Kitahara, and Masataka Kaneko. IFSGen4L^{A}TEX:

Interactive graphical user interface for generation and visualization of iterated function systems in L^{A}TEX. In Hong and Yap [HY14], pages 554– 561. ISBN 3-662-44198-5 (paperback), 3-662-44199-3 (e-book). LCCN QA76.9.M35.

Gehani:1988:DFT

[GL88] Narain Gehani and Steven Lally. Document formatting and typesetting on the UNIX system. Vol. 2: grap, mv, ms, and troff. Silicon Press, 25 Beverly Road, Summit, NJ 07901, USA, 1988. ISBN 0-9615336-3-3. xiii + 304 pp. LCCN Z52.5.U54 G431 1988. US$30.95.

Goossens:2008:LGC

[GMR^{+}08] Michel Goossens, Frank Mittelbach, Sebastian Rahtz, Denis Roegel, and Herbert Voß, editors. The L^{A}TEX Graphics Companion. Addison-Wesley series on tools and techniques for computer typesetting. Addison-Wesley, Reading, MA, USA, second edition, 2008. ISBN 0-321-50892-0. LCCN Z253.4.L38G663 2008.

Goossens:1994:LC

[GMS94a] Michel Goossens, Frank Mittelbach, and Alexander Samarin. The L^{A}TEX Companion. Tools and Techniques for Computer Typesetting. Addison-Wesley, Reading, MA, USA, 1994. ISBN 0-201-54199-8 (paperback), 0-201-77591-3 (hardcover). xxi + 528 + 2 pp. LCCN Z253.4.L38 G66 1994.

Goossens:1994:LB

[GMS94b] Michel Goossens, Frank Mittelbach, and Alexander Samarin. Der L^{A}TEX-Begleiter. Addison-Wes-ley Verlag, Bonn, Germany, 1994. ISBN 3-89319-646-3, 3-8273-1689-8. xxxiii + 554 pp. LCCN ???? DM 79,90; ATS 583,00; CHF 73,00.

Goncalves:2004:FRM

[Gon04] Luis Nobre Gonçalves. and a review of 3D packages. In Syropoulos et al. [SBH^{+}04], pages 112–124. CO-DEN LNCSD9. ISBN 3-540-22801-2 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN Z253.3 I58 2004. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t3130.htm;
 http://www.springerlink.
com/openurl.asp?genre=
issue&issn=0302-9743&
volume=3130;   http:
//www.springerlink.
com/openurl.asp?genre=

volume&id=doi:10.1007/
b99374.

Gourlay:1986:LMP

[Gou86] John S. Gourlay. A language for music printing. Software—Practice and Experience, 29(5): 388–401, May 1986. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Gonczarowski:1989:ICP

[GP89] Jakob Gonczarowski and On G. Paradise. InScript —a C-like preprocessor for PostScript. Electronic Publishing—Origination, Dissemination, and Design, 2(3):157– 167, October 1989. CO-DEN EPODEU. ISSN 0894-3982.

Garvey:1982:NEM

[GR82] Daniel E. Garvey and William L. Rivers. Newswriting for the Electronic Media: Principles, Examples, Applications. Wadsworth Publishing Company, Belmont, CA, USA, 1982. ISBN 0-534-01069-5. xiii + 250 pp. LCCN PN4784.B75 .G3. Exercises, appendices, index, pp. 250.

Granet:1995:OL

[GR95] Vincent Granet and Jean-Pierre Regourd. Objectif L^{A}TEX. Objectif, 1240-1013. Masson Editeur, Paris, France, 1995. ISBN 2-225-84993-5. 303 pp. LCCN ????

Goossens:1999:LWC

[GR99] Michel Goossens and Sebastian Rahtz. The L^{A}TEX Web companion: integrating TEX, HTML, and XML. Tools and Techniques for Computer Typesetting. Addison-Wesley Longman, Reading, MA, USA, 1999. ISBN 0-201-43311-7. xxii + 522 pp. LCCN QA76.76.H94G66 1999. US$36.95. With Eitan M. Gurari and Ross Moore and Robert S. Sutor.

Graetzer:1993:A

[Grä93a] George Grätzer. AMS-L^{A}TEX. Notices of the American Mathematical Society, 40(2):148–150, February 1993. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Graetzer:1993:ATI

[Grä93b] George Grätzer. Advances in TEX implementations. I. PostScript fonts. Notices of the American Mathematical Society, 40(7):834–838, September 1993. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Gratzer:1993:MTS

[Grä93c] George Grätzer. Math into TEX: a Simplified Introduction Using AMS-L^{A}TEX. Springer-Verlag and Birkhäuser, Berlin, Germany / Heidelberg, Germany / London, UK / etc. and Cambridge, MA, USA; Berlin, Germany; Basel, Switzerland, 1993. ISBN 0-8176-3637-4, 3-7643-3637-4. xxix + 294 pp. LCCN Z253.4.A65G7 1993. US$34, SFr 52, DM 57.

Gratzer:1994:ATIa

[Grä94a] George Grätzer. Advances in TEX implementations. II. integrated environments. Notices of the American Mathematical Society, 41(2):106– 111, February 1994. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Gratzer:1994:ATIb

[Grä94b] George Grätzer. Advances in TEX implementations. III. A new version of L^{A}TEX, finally. Notices of the American Mathematical Society, 41 (6):611–615, July/ August 1994. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Gratzer:1996:MLI

[Grä96] George Grätzer. Math into L^{A}TEX: an introduction to L^{A}TEX2e and AMS-L^{A}TEX 1.2. Birkhäuser, Cambridge, MA, USA; Berlin, Germany; Basel, Switzerland, 1996. ISBN 0-8176-3805-9, 3-7643-3805-9. xxvii + 451 pp. LCCN Z253.4.A65G69 199. US$49.50.

Gratzer:1999:FSL

[Grä99] George Grätzer. First steps in L^{A}TEX. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1999. ISBN 0-8176-4132-7, 3-7643-4132-7. xx + 131 pp. LCCN Z253.4.L38 G74 1999.

Gratzer:2000:ML

[Grä00] George Grätzer. Math into L^{A}TEX. Birkhäuser and Springer-Verlag, Cambridge, MA, USA; Berlin, Germany; Basel, Switzerland and Berlin, Germany / Heidelberg, Germany / London, UK / etc., third edition, 2000. ISBN 0-8176-4131-9, 3-7643-4131-9. xxv + 584 pp. LCCN Z253.4.L38 G745 2000. US$49.95.

Gratzer:2007:ML

[Grä07] George Andrew Grätzer. Math into L^{A}TEX. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., fourth edition, 2007. ISBN 0-387-32289-2. xxxiv + 619 pp. LCCN Z253.4.L38 G745 2007. Forward by Rainer Schöpf.

Gratzer:2009:WNLa

[Grä09a] George Grätzer. What is new in L^{A}TEX? I. Breaking free. Notices of the American Mathematical Society, 56(1):52–54, January 2009. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic). URL http://

www.ams.org/notices/
200901/tx090100052p.
pdf.

Gratzer:2009:WNLb

[Grä09b] George Grätzer. What is new in L^{A}TEX? II. TEX implementations, evolution or revolution? Notices of the American Mathematical Society, 56(5):627–629, May 2009. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic). URL http://

www.ams.org/notices/
200905/rtx090500627p.
pdf.

Gratzer:2009:WNLc

[Grä09c] George Grätzer. What is new in L^{A}TEX? III. Formatting references. Notices of the American Mathematical Society, 56 (8):954–956, September 2009. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic). URL http://

www.ams.org/notices/
200908/rtx090800954p.
pdf.

Gratzer:2011:WNL

[Grä11] George Grätzer. What is new in L^{A}TEX? IV. WYSIWYG L^{A}TEX. Notices of the American Mathematical Society, 58 (6):828–830, June/July 2011. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic). URL http://

www.ams.org/notices/
201106/rtx110600828p.
pdf.

Gratzer:2013:WNL

[Grä13] George Grätzer. What is new in L^{A}TEX? V. L^{A}TEX on an iPad. foundation. Notices of the American Mathematical Society, 60(3):332–334, March 2013. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic). URL http://

www.ams.org/notices/
201303/rnoti-p332.pdf.

Gratzer:2014:PL

[Grä14] George Grätzer. Practical L^{A}TEX. Springer International Publishing, Cham, Switzerland, 2014. ISBN 3-319-06424-X (print), 3-319-06425-8 (e-book). xvi + 216 + 55 + 20 pp. LCCN QA76.9.T48 2014; TK5105.875.I57; Z253.4.L38 G746 2014.

URL http://link.springer.
com/10.1007/978-3-319-
06425-3;   http://www.
springerlink.com/content/
978-3-319-06425-3.

Gratzer:2016:MML

[Grä16] George Grätzer. More Math Into L^{A}TEX. Springer International Publishing, Cham, Switzerland, fifth edition, 2016. ISBN 3-319-23795-0, 3-319-23796-9 (e-book). xxx + 609 pp. LCCN Z253.4.L38 G745 2016.

URL http://www.springerlink.
com/content/978-3-319-
23796-1.

Greenberg:2005:SIL

[Gre05] Harvey J. Greenberg.

Simplified introduction to L^{A}TEX. World-Wide Web document, March 21, 2005. URL http://

carbon.cudenver.edu/
~hgreenbe/aboutme/simplified-
intro.html;   http:/
/www.cudenver.edu/~
hgreenbe/courses/texinfo/

book/README;   http:
//www.cudenver.edu/~
hgreenbe/courses/texinfo/
book/simplified-intro.
zip.

Gries:1990:BDO

[Gri90] David Gries. Binary to decimal, one more time. In Feijen et al. [FvGGM90], chapter 16, pages 141–148. ISBN 0-387-97299-4, 3-540-97299-4, 1-4612-8792-8 (print), 1-4612-4476-5 (online). ISSN 0172-603X. LCCN QA76 .B326 1990. URL https:

//link.springer.com/
chapter/10.1007/978-
1-4612-4476-9_17. This

paper presents an alternate proof of Knuth’s algorithm [Knu90b] for conversion between decimal and fixed-point binary numbers.

Goossens:1997:LGC

[GRM97] Michel Goossens, Sebastian Rahtz, and Frank Mittelbach. The L^{A}TEX Graphics Companion: Illustrating Documents with TEX and PostScript. Tools and Techniques for Computer Typesetting. Addison-Wesley, Reading, MA, USA, 1997. ISBN 0-201-85469-4. xxi + 554 pp. LCCN Z253.4.L38G663 1997. US$39.75.

GeyerSchulz:1993:TAL

[GSMN93] A. Geyer-Schulz, J. Matulka, and G. Neumann. Typesetting APL with L^{A}TEX–the APL front end APL2TEX. APL Quote Quad, 23(3):22– 27, March 1993. CODEN APLQD9. ISSN 0163-6006.

Guntermann:1985:GDL

[Gun85] Klaus Guntermann. GEN-DRIV — a driver for low cost devices using builtin fonts. In Lucarella [Luc85b], pages 197– 204. ISBN 0-201-13399-7. LCCN Q224.E97 1985.

Gunther:1996:EL

[Gün96] Karsten Günther. Einführung in L^{A}TEX 2ε: Lehrbuch und Referenz. dpunkt-Verlag, Heidelberg, Germany, 1996. ISBN 3-920993-36-5. xxxiv + 522 pp. DM 68,00; ATS 496,00; CHF 60,00.

Gurari:1994:TLD

[Gur94a] Eitan M. Gurari. TEX and L^{A}TEX: Drawing and Literate Programming. McGraw-Hill, New York, NY, USA, 1994. ISBN 0-07-025208-4 (book), 0-07-911616-7 (book + diskette). xiv + 310 pp. LCCN Z253.4.T47G87 1994. US$34.95.

Gurari:1994:WT

[Gur94b] Eitan M. Gurari. Writing with TEX. McGraw-Hill, New York, NY, USA, 1994. ISBN 0-07-025207-6. xiv + 249 pp. LCCN Z253.4.T47 G87 1994.

Gutierrez:2004:IJU

[Gut04] Jaime Gutierrez, editor. ISAAC 2004: July 4– 7, 2004, University of Cantabria, Santander, Spain: proceedings of the 2004 International Symposium on Symbolic and Algebraic Computation. ACM Press, New York, NY, USA, 2004. ISBN 1-58113-827-X. LCCN ????

Gubinelli:2014:GTM

[GvdHPR14] Massimiliano Gubinelli, Joris van der Hoeven, François Poulain, and Denis Raux. GNU TEXMACS: towards a scientific office suite. In Hong and Yap [HY14], pages 562–569. ISBN 3-662-44198-5 (paperback), 3-662-44199-3 (ebook). LCCN QA76.9.M35.

Hersch:1998:EPA

[HAB98] Roger Hersch, Jacques André, and Heather Brown, editors. Electronic publishing, artistic imaging, and digital typography: 7th International Conference on Electronic Publishing, EP ’98 held jointly with the 4th International Conference on Raster Imaging and Digital Typography, RIDT ’98, St. Malo, France, March/ April 1998: proceedings, volume 1375 of Lecture Notes in Computer Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1998. CODEN LNCSD9. ISBN 3-540-64298-6. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN Z250.7.I58 1998.

Hafner:1997:TLL

[Haf97] Thomas Hafner. TEX und L^{A}TEX leicht gemacht: universelles Satzsystem für jeden Computer. Pflaum, München, Germany, 1997. ISBN 3-7905-1505-1. 108 pp. LCCN ???? DM 22,80.

Hagen:2025:LT

[Hag25] Hans Hagen. Lua in TEX.

Journal of Computer Languages, 85(??):1–13, November 2025. CO-DEN ???? ISSN 2665-9182 (print), 2590-1184 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S2590118425000401.

Hahn:1991:LER

[Hah91]  Jane Hahn.  L^{A}TEX

for Everyone—A Reference Guide and Tutorial for Typesetting Documents Using a Computer. PersonalTEX Inc., 12 Madrona Street, Mill Valley, CA 94941, USA, 1991. ISBN 0-9631044-0-3. xi + 346 pp. LCCN Z253.4.L38 H34 1991. US$19.95.

Hahn:1993:LER

[Hah93]  Jane Hahn.  L^{A}TEX

for Everyone—A Reference Guide and Tutorial for Typesetting Documents Using a Computer. Prentice-Hall, Upper Saddle River, NJ 07458, USA, second edition, 1993. ISBN 0-13-605908-2. xi + 346 pp. LCCN Z253.4.L38 H35 1993. US$19.95.

Haliburton:1987:ATC

[Hal87] John Haliburton, editor. APL in Transition: Conference Proceedings, APL87, May 10–14, 1987, Dallas, Texas. ACM Press, New York, NY, USA, 1987. ISBN 0-89791-226-8. LCCN ???? US$37.00. APL quote quad, v. 17, no. 4.

Han:2000:MTE

[Hàn Thế Thành00] Hàn Thế Thành. Microtypographic extensions to the TEX typesetting system. Ph.D. dissertation, Faculty of Informatics, Masaryk University, Brno, Czech Republic, October 2000. 118 pp. URL http:/

/www.pragma-ade.com/
pdftex/thesis.pdf.

Hart:1991:HRC

[Har91] Horace Hart. Hart’s Rules; For Compositors and Readers at the University Press, Oxford. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 39th edition, 1991. ISBN 0-19-212983-X. 182 pp. LCCN Z254 .H37 1983. UK£8.

Haralambous:1992:TLO

[Har92a] Yannis Haralambous.

TEX et les langues orientales. ????, ??:??, 1992.

Haralambous:1992:TRT

[Har92b] Yannis Haralambous.

Towards the revival of traditional Arabic typography … through TEX. In Zlatuška [Zla92], pages 293–305. ISBN 80-210-0480-0. Invited talk.

Harrower:1992:NDH

[Har92c]  Tim Harrower.  The

Newspaper Designer’s Handbook. William C. Brown Company Publishers, Dubuque, IA, USA, 1992. ISBN 0-697-13382-6. 180 pp. LCCN Z253.5 .H27 1992.

Haralambous:1993:PPF

[Har93a] Y. Haralambous. Parametrization of PostScript fonts through METAFONT: an alternative to Adobe Multiple Master fonts. Electronic Publishing— Origination, Dissemination, and Design, 6 (3):145–158, September 1993. CODEN EPODEU. ISSN 0894-3982.

Hartz:1993:M

[Har93b] David Hartz. MathType 3.0. Notices of the American Mathematical Society, 40(5):469–470, May/June 1993. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic). Review of a Macintosh equation editor front-end to TEX and various word processors.

Haralambous:1994:STB

[Har94] Yannis Haralambous. Un système tex berbère. Etudes et documents berbères, 11:43–53, 1994.

Hartley:2002:HA

[Har02a] James Hartley. Here’s the address: Improving information in abstracting journals and databases. Social Studies of Science, 32(4): 637–638, August 1, 2002. CODEN SSS-CDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https:

//journals.sagepub.
com/doi/pdf/10.1177/
0306312702032004007.

Hartley:2002:CTS

[Har02b]  James Hartley.  On

choosing typographic settings for reference lists. Social Studies of Science, 32(5–6): 917–932, December 1, 2002. CODEN SSS-CDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https:

//journals.sagepub.
com/doi/pdf/10.1177/
030631270203200512;
 https://www.jstor.
org/stable/3183059.

Haralambous:2004:FCG

[Har04] Yannis Haralambous.

Fontes & codages: Glyphes et caractères à l’ère du numérique. O’Reilly, Paris, France, 2004. ISBN 2-84177-273-X. xx + 990 pp. LCCN

???? URL http://www.
oreilly.com/catalog/
9782841772735.

Hayward:1977:SA

[Hay77] Stan Hayward. Scriptwriting for Animation. Hastings House, Publishers, New York, NY, USA, 1977. ISBN 0-8038-6741-7. 160 pp. LCCN TR 897.5 H335s, PN1996 .H37 1977.

Hayashi:1987:TTH

[Hay87]  T. Hayashi.  Trans-

porting TEX into the HITAC VOS3 environment. Software—Practice and Experience, 17 (1):17–28, January 1987. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Hayes:1996:SM

[Hay96] Brian Hayes. Speaking of mathematics. American Scientist, 84(2):110–113, March/April 1996.

Hunter:1995:GSS

[HB95] Roger Hunter and Susan Anne Bagby. Getting started with Scientific Word and Scientific WorkPlace for Windows: professional and student editions. Brooks/Cole Publishing Co., Pacific Grove, CA, USA, 1995. ISBN 0-534-34097-0. vi + 121 pp. LCCN Q183.9 .H86 1995.

Hunter:1996:CDS

[HB96] Roger Hunter and Susan Anne Bagby. Creating documents with Scientific WorkPlace and Scientific Word. Brooks/ Cole Publishing Co., Pacific Grove, CA, USA, 1996. ISBN 0-534-34369-4. xxiv + 246 pp. LCCN Q183.9 .H84 1996.

Hefferon:2009:TF

[HB09] Jim Hefferon and Karl Berry. The TEX family in 2009. Notices of the American Mathematical Society, 56(3):348– 354, March 2009. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic). URL

http://www.ams.org/
journals/notices/200903/
tx090300348p.pdf.

Hersch:1995:PTG

[HBBG95] Roger D. Hersch, Claude Bétrisey, Justin Bur, and André Gürtler. Perceptually tuned generation of grayscale fonts. IEEE Computer Graphics and Applications, 15(6):78– 90, November 1995. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Hammerlindl:2004:ASB

[HBP04] Andy Hammerlindl, John Bowman, and Tom

Prince. Asymptote: a

script-based vector graphics language. Faculty of Science, University of Alberta, Edmonton, AB, Canada, 2004. URL http://

asymptote.sourceforge.
net/.  Asymptote is

a powerful script-based vector graphics language for technical drawing, inspired by - but with an improved C++-like syntax. Asymptote provides for figures the same highquality level of typesetting that L^{A}TEX does for scientific text. Heinemann:1988:XPS

[HC88] Sue Heinemann and Virginia Croft, editors. Xerox Publishing Standards: a Manual of Style and Design. Watson-Guptill Publications, 1515 Broadway, New York, NY 10036, USA, 1988. ISBN 0-8230-5964-2. LCCN Z286.L68 X47 1988. US$35.00.

Heck:1992:DPA

[Hec92] André Heck, editor. Desktop Publishing in Astronomy & Space Sciences. Strasbourg Astronomical Observatory, [1–3] October, 1991. World Scientific, Singapore, 1992. ISBN 981-02-0915-0. LCCN QB14.2 .D47 1992. In addition to the papers presented at the colloquium, each paper is followed by a summary of the discussion about that paper.

Heck:2003:LMD

[Hec03]  André Heck.  Learn-

ing by doing. Report, AMSTEL Institute, ????, 2003. 93 pp. URL http:

//math.nju.edu.cn/~
meijq/tex/mptutans.
pdf.

Hegron:1984:AGR

[Hég84]  Gérard Hégron.  Al-

gorithmes de génération de caractères. Publication interne 244, IRISA, Rennes, France, 1984.

Heilmann:1996:LVK

[Hei96] Axel Heilmann. L^{A}TEX-Vademecum: ein Kompaktführer für Einsteiger und Fortgeschrittene. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1996. ISBN 3-540-60522-3. xix + 281 pp. DM 38,00.

Hershey:1969:FIP

[Her69] Allen V. Hershey. FOR-TRAN IV programming for cartography and typography. Technical Report TR-2339, U. S. Naval Weapons Laboratory, Dahlgren, VA 22448, USA, September 1969.

Hershey:1970:PRF

[Her70] Allen V. Hershey. Preparation of reports with the FORTRAN typographic system. Technical Report TN-K/27-70, U.

S. Naval Weapons Laboratory, Dahlgren, VA 22448, USA, September 1970.

Hershey:1972:CSS

[Her72] Allen V. Hershey. A computer system for scientific typography. Computer Graphics and Image Processing, 1:373– 385, 1972. CODEN CGIPBG. ISSN 0734-189X.

Hershey:1981:ACT

[Her81] Allen V. Hershey. Advanced computer typography. Technical Report NPS012-81-005, U. S. Naval Postgraduate School, Monterey, CA 93940, USA, December 1981.

Hersch:1987:CGU

[Her87] Roger D. Hersch. Character generation under grid constraints. Computer Graphics, 21(4): 243–252, July 1987. CO-DEN CGRADI, CPGPBZ. ISSN 0097-8930. URL

                                        http://www.acm.org:
                                        80/pubs/citations/proceedings/
                                        graph/37401/p243-hersch/

.

Hersch:1993:VTA

[Her93] Roger D. Hersch. Visual and Technical Aspects of Type. Cambridge University Press, New York, 1993. ISBN 0-521-44026-

2. 204 pp. LCCN Z250.7 .V62 1993. US$59.95.

Hobby:1983:CMF

[HG83] John Hobby and Gu Guoan.

A Chinese Meta-Font. Report STAN-CS-83-974, Stanford University, Department of Computer Science, Stanford, CA, USA, 1983.

Haralambous:2007:FE

[HH07] Yannis Haralambous and P. Scott Horne. Fonts and encodings. O’Reilly Media, Inc., 1005 Gravenstein Highway North, Sebastopol, CA 95472, USA, 2007. ISBN 0-596-10242-9. xx + 1016 pp. LCCN Z250. URL

http://www.oreilly.
com/catalog/9780596102425.

Hickey:1989:UST

[Hic89]  T. B. Hickey.  Us-

ing SGML and TEX for an interactive chemical encyclopedia. In National Online Meeting proceedings–1989, New York, May 9–11, 1989, pages 187–195. Learned Information, Medford, NJ, USA, 1989. ISBN 0-938734-34-2. LCCN QA76.55 .N37 1989.

Higham:1993:HWM

[Hig93] Nicholas J. Higham.

Handbook of Writing for the Mathematical Sciences. Society for Industrial and Applied Mathematics, Philadelphia, PA, USA, June 1993. ISBN 0-89871-314-5. xii + 241 pp. LCCN QA42.H54 1993. US$21.50. [From the publisher]: Having trouble with your latest math paper? Giving a presentation that you just can’t pull together? Struggling with your thesis or trying to get your first article published in a technical journal? Handbook of Writing for the Mathematical Sciences is the book for you! This handy volume provides information on virtually every issue you will face when writing a technical paper or talk, from choosing the right journal to handling your references. You’ll also get an overview of the entire publication process– invaluable for anyone hoping to publish in a technical journal. To write a truly impressive paper, you’ll need to understand the anatomy of a research paper and the steps involved in revising a draft. This book offers discussions of these fundamental topics, along with illustrative and provocative examples. Also included are chapters on standard English usage, using computers for writing and research, and writing technical material when English is a foreign language. This handbook provides much-needed advice on handling the basic ingredients of a research paper, like definitions, theorems, examples, and equations. In addition, appendices provide essential reference material, including summaries of L^{A}TEX symbols and Emacs commands, addresses of mathematical societies, and a list of papers that have won expository writing prizes. This book is ideal for graduate students and teachers. Among its special features:

• Detail and extensive use of examples make it an excellent teaching tool

• Discusses TEX and other software tools for preparing publications

• Thorough treatment, with examples, of how to write slides (transparencies) for a mathematical talk.

• Comprehensive index and extensive bibliography make it an excellent reference

• Readable for everyone in the field– from undergraduates to seasoned professionals

Contents Preface; Chapter 1: General Principles; Chapter 2: Writer’s Tools and Recommended Reading; Chapter 3: Mathematical Writing; Chapter 4: English Usage; Chapter 5: When English is a Foreign Language; Chapter 6: Writing a Paper; Chapter 7: Revising a Draft; Chapter 8: Publishing a Paper; Chapter 9: Writing a Talk; Chapter 10: Computer Aids for Writing and Research; Appendix A: The Greek Alphabet; Appendix B: Summary of TEX and L^{A}TEX Symbols; Appendix C: GNU Emacs– The Sixty+ Most Useful Commands; Appendix D: Mathematical Organizations in the UK and USA; Appendix E: Winners of Prizes for Expository Writing; Appendix F: Glossary; Bibliography; Index. About the Author: Nicholas J. Higham is a Reader in Mathematics at the University of Manchester, UK. He is the author of more than 40 publications and is a member of the editorial board of the SIAM Journal on Matrix Analysis and Applications.

Higham:1998:HWM

[Hig98] Nicholas J. Higham.

Handbook of Writing for the Mathematical Sciences. Society for Industrial and Applied Mathematics, Philadelphia, PA, USA, second edition, 1998. ISBN 0-89871-420-6 (paperback), 0-89871-955-0 (e-book). xvi + 302 pp. LCCN QA42 .H54 1998. Higham:2020:HWM

[Hig20] Nicholas J. Higham.

Handbook of Writing for the Mathematical Sciences. Society for Industrial and Applied Mathematics, Philadelphia, PA, USA, 2020. ISBN 91-611-9760-9-X (paperback), 1-61197-610-3 (ebook). xxi + 353 pp. LCCN QA42 .H54 2020.

Hirschberg:1987:NAF

[HL87] D. S. Hirschberg and L. L. Larmore. New applications of failure functions. Journal of the ACM, 34(3):616–625, July 1987. CODEN JA-COAH. ISSN 0004-5411 (print), 1557-735X (electronic). URL http:

                                        //www.acm.org/pubs/
                                        toc/Abstracts/0004-
                                        5411/28875.html.

Harrison:1989:IBP

[HM89] M. A. Harrison and E. V. Munson. On integrated bibliography processing. Electronic Publishing—Origination, Dissemination, and Design, 2(4):193–209, December 1989. CODEN EPODEU. ISSN 0894-3982.

Haarslev:1990:FVO

[HM90a] V. Haarslev and R. Moller. A framework for visualizing object-oriented systems. ACM SIGPLAN Notices, 25(10):237–244, October 1990. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). OOPSLA ECOOP ’90 Conference on Object-Oriented Programming: Systems, Languages, and Applications.

Hopkins:1990:RA

[HM90b] T. Hopkins and D. Morse.

Remark on algorithm 620. ACM Transactions on Mathematical Software, 16(4):401–403, December 1990. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Haarslev:1992:VGL

[HM92] V. Haarslev and R. Moller.

Visualization and graphical layout in objectoriented systems. Journal of Visual Languages and Computing, 3(1):1– 23, March 1992. CODEN JVLCE7. ISSN 1045-926X.

Harrington:2004:AMA

[HNJ^{+}04] Steven J. Harrington, J. Fernando Naveda, Rhys Price Jones, Paul Roetling, and Nishant Thakkar. Aesthetic measures for automated document layout. In Ethan V. Munson and Jean-Yves Vion-Dury, editors, Proceedings of the 2004 ACM symposium on Document Engineering: October 28–30, Milwaukee, Wisconsin, USA, DocEng ’04, pages 109–111. ACM Press, New York, NY, USA, 2004. ISBN 1-58113-938-1. LCCN HF5736 .A27 2004. URL http:/

                                        /portal.acm.org/toc.
                                        cfm?id=1030397.

Hobby:1986:DBT

[Hob86] John Douglas Hobby. Digitized Brush Trajectories. Ph.D. dissertation, Department of Computer Science, Stanford University, Stanford, CA, USA, June 1986. 151 pp.

                                        URL http://wwwlib.
                                        umi.com/dissertations/
                                        fullcit/8602484. Also

published as report STAN-CS-1070 (1985).

Hobby:1992:IM

[Hob92] John D. Hobby. Introduction to - . In Zlatuška [Zla92], pages 21–36. ISBN 80-210-0480-0. Invited talk.

Hobby:1995:DGM

[Hob95] John D. Hobby. Drawing Graphs with . AT&T Bell Laboratories, Murray Hill, NJ, USA, 1995. 18 pp. URL

                                        http://ctan.tug.org/
                                        tex-archive/macros/
                                        latex/contrib/pdfslide/
                                        mpgraph.pdf.

Hobby:1997:MS

[Hob97] John D. Hobby. The System. ????, ????, December 30, 1997.

                                        URL file:///texlive-
                                        2004-11/texmf-dist/
                                        doc/metapost/base/mpintro.
                                        pdf.

Hobby:2001:MDS

[Hob01]  John D. Hobby.  -

le dessin sous TEX: manuels de l’utilisateur pour et graph. Cahiers GUTenberg, 41(??):175, 2001. ISSN 1140-9304. French translation by Pierre Fournier et Jean-Côme Charpentier.

Hobby:2004:UMM

[Hob04] John D. Hobby. A User’s Manual for . ????, ????, 2004. i + 85 pp. URL file:

///texlive-2004-11/
texmf-dist/doc/metapost/
base/mpman.pdf.

Hoenig:1992:WTM

[Hoe92]  Alan Hoenig.  When

TEX and work together. In Zlatuška [Zla92], pages 1–20. ISBN 80-210-0480-0. Invited talk.

Hoenig:1998:TUL

[Hoe98]  Alan Hoenig.  TEX

Unbound: L^{A}TEX and TEX Strategies for Fonts, Graphics, & More. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 1998. ISBN 0-19-509686-X (paperback), 0-19-509685-1 (hardcover). ix + 580 pp. LCCN Z253.4.L38H64 1997. US$60.00 (hardcover), US$35.00 (paperback).

                                        URL http://www.oup-
                                        usa.org/gcdocs/gc_0195096851.
                                        html.

Hofstadter:1982:MFM

[Hof82] Douglas R. Hofstadter. Meta-font, metamathematics, and metaphysics: Comments on Donald Knuth’s article (16:1). Visible Language, 16(4): ??, October 1982. CO-DEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL

                                        https://s3-us-west-
                                        2.amazonaws.com/visiblelanguage/
                                        pdf/V16N4_1982_E.pdf.

Hohti:1988:TAT

[Hoh88]  Aarno Hohti.  Type-

setting APL with TEX. ACM SIGAPL APL Quote Quad, 18(3):13– 16, March 1988.

Holzgang:1987:UPP

[Hol87] David A. Holzgang. Understanding Post-Script Programming. Sybex, 2021 Challenger Driver, Suite 100, Alameda, CA 94501, USA, 1987. ISBN 0-89588-396-1. xxxii + 459 pp. LCCN QA76.73.P67 H65 1987.

Holzgang:1989:PPR

[Hol89] David A. Holzgang. PostScript Programmer’s Reference Guide. Scott, Foresman and Company, Glenview, IL, USA, 1989. ISBN 0-673-38574-4. x + 486 pp. LCCN QA76.73.P67 H64 1989. Holkner:2006:GMO

[Hol06]  A. Holkner.  Global

multiple objective line breaking. Master’s thesis, School of Computer Science and Information Technology, RMIT University, Melbourne, Victoria, Australia, 2006.

Hoover:1992:KSS

[Hoo92] Anita Z. Hoover. The key to successful support: knowing your TEX and L^{A}TEX users. In Zlatuška [Zla92], pages 71–85. ISBN 80-210-0480-0. Invited talk. Howland:1986:TAU

[How86] John E. Howland. Typesetting APL using a macintosh. ACM SIGAPL APL Quote Quad, 16(4): 301–305, 1986.

Holecek:2004:APG

[HS04] Jan Holeček and Petr Sojka. Animations in pdfTEX-generated PDF: a new method for directly embedding animation into PDF. In Syropoulos et al. [SBH^{+}04], pages 179–191. CO-DEN LNCSD9. ISBN 3-540-22801-2 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN Z253.3 I58 2004. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t3130.htm;
 http://www.springerlink.
com/openurl.asp?genre=
issue&issn=0302-9743&
volume=3130;   http:
//www.springerlink.
com/openurl.asp?genre=
volume&id=doi:10.1007/
b99374.

Hufflen:2004:MBL

[Huf04] Jean-Michel Hufflen. MlBibTEX: Beyond L^{A}TEX. In Syropoulos et al. [SBH^{+}04], pages 203–215. CO-DEN LNCSD9. ISBN 3-540-22801-2 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN Z253.3 I58 2004. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t3130.htm;
 http://www.springerlink.
com/openurl.asp?genre=
issue&issn=0302-9743&
volume=3130;   http:
//www.springerlink.
com/openurl.asp?genre=
volume&id=doi:10.1007/
b99374.

Huggins:1988:EHM

[Hug88] K. Cleo R. Huggins. Egyptian hieroglyphs for modern printing devices. Technical Report STAN-CS-89-1251, Stanford University, Stanford, CA, USA, June 1988.

Hunt:1994:BRT

[Hun94] R. Hunt. Bibliographic references: Too many variations on a theme? The EASE Journal: European Science Editing, (52):7–8, July 1994. ISSN 0950-1835.

Hurlburt:1978:GMS

[Hur78] Allen Hurlburt. The grid:

a modular system for the design and production of newspapers, magazines, and books. Van Nostrand Reinhold, New York, NY, USA, 1978. ISBN 0-442-23598-4. 96 pp. LCCN Z246 .H82. US$15.00.

Hardy:1995:DMSa

[HW95a] Darel W. Hardy and Carol L. Walker. Doing mathematics with Scientific WorkPlace. Brooks/ Cole Publishing Co., Pacific Grove, CA, USA, 1995. ISBN 0-534-25284-2. xviii + 291 pp. LCCN QA76.95 .H37 1995.

Hardy:1995:DMSb

[HW95b] Darel W. Hardy and Carol L. Walker. Doing mathematics with Scientific WorkPlace. Brooks/ Cole Publishing Co., Pacific Grove, CA, USA, revised edition, 1995. ISBN 0-534-34049-0. xix + 340 pp. LCCN QA76.95 .H37 1995b. Heckmann:1997:FDT

[HW97] Reinhold Heckmann and Reinhard Wilhelm. A functional description of TEX’s formula layout. Journal of Functional Programming, 7 (5):451–485, September 1997. CODEN JF-PRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https:

//www.cambridge.org/
core/product/18A13702FF264BBA3894B55E01832976.

Hong:2014:MSI

[HY14] Hoon Hong and Chee Yap, editors. Mathematical Software — ICMS 2014: 4th International Conference, Seoul, South Korea, August 5– 9, 2014, Proceedings, volume 8592. Springer-Ver-lag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2014. ISBN 3-662-44198-5 (paperback), 3-662-44199-3 (e-book). LCCN QA76.9.M35.

Ibrus:1992:VET

[Ibr92] Hans Ibrus. Väike Eesti TEXiraamat. a/S “KompuT” kirjastus, Tartumaa, Estonia, 1992. 144 pp. In Estonian. Appendix A is a translation into Estonian of [Sam83].

ICCPCOL:1988:ICC

[ICC88] 1988 International Conference on Computer Processing of Chinese and Oriental Languages, August 29–September 1, 1988, Toronto, Canada. Concordia University, Montréal, Québec, Canada, 1988. LCCN QA76.9.T48 I48 1988.

IEEE:1992:PEC

[IEE92] Proceedings /The European Conference on Design Automation, Brussels, Belgium, March 16– 19, 1992. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1992. ISBN 0-8186-2645-3. LCCN TK7867 .E93 1992.

Ion:1988:TEU

[II88] P. D. F. Ion and B. D. R.

Ion. TEX and its uses for information presentation. Bulletin of the American Society for Information Science, 14(4): 21–22, April–May 1988. CODEN BASICR. ISSN 0095-4403.

Ikarus:1983:ITD

[Ika83] URW Unternehmensberatung Karow Rubow Weber GMBH, Hamburg, Germany. IKARUS for Typefaces in Digital Form, 1983.

Incerpi:1986:GFT

[IM86a] Janet Incerpi and Francis Montagnac. Getting faster TEX output. using the DVI library. Rapports Techniques 71, INRIA/IRISA, Rennes, France, July 1986.

Incerpi:1986:UDL

[IM86b] Janet Incerpi and Francis Montagnac. Using the DVI library. In Getting Faster TEX Output. INRIA/IRISA, Rennes, France, 1986.

Iseri:1989:GPP

[IO89] M. Iseri and T. Ohno. General purpose page composition language— TEX. Journal of the Society of Instrument and Control Engineers, 28(3): 213–217, March 1989. CODEN KESEA4. ISSN 0453-4662.

Iseri:1990:PVD

[Ise90] M. Iseri. A publisher’s view of DTP. Journal of Information Processing, 31(11):1501–1507, 1990. CODEN JOSHA4. ISSN 0447-8053.

ISO:1987:DBR

[ISO87]  ISO.  Documentation

— Bibliographic References — Content, Form and Structure (ISO 690). International Organization for Standardization, Geneva, Switzerland, 1987. ???? pp.

Jackowski:1993:MZP

[Jac93] Bogusław Jackowski. Metafont: zastosowania praktyczne i niepraktyczne. broszura z konferencji członków GUSTu, Bachotek, 1993. In Polish.

Jackson:1994:WHN

[Jac94] Allyn Jackson. What’s Happening: New publication takes the AMS in new direction. Notices of the American Mathematical Society, 41(6): 626–627, July/August 1994. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Jackowski:2017:PD

[Jac17] Paweł Jackowski. Pearls’ Decade. GUST, Bachotek, Poland, 2017. ISBN 83-939016-3-4. ???? pp. LCCN ????

Janson:1986:TTP

[Jan86] B. J. Janson. Technical typesetting: a publisher’s view. In Constance U. Greaser and Ethel G. Langlois, editors, Seeking scholarly perspectives: Proceedings of the Eighth Annual Meeting of the Society for Scholarly Publishing, pages 66–68. Society for Scholarly Publishing, Washington, DC, USA, 1986. US$15.00.

Jackowski:1988:TPF

[JHR88] Bogusław Jackowski, Tomasz Hołdys, and Marek Ry´ cko. With TEX to the Poles: a few parochial remarks on using TEX by the non-English. Anima Ltd, Gda´ nsk, 1988.

Jacobs:2003:ADL

[JLS03a] Charles Jacobs, Wilmot Li, and David H. Salesin. Adaptive document layout via manifold content. In ????, editor, Second International Workshop on Web Document Analysis (wda2003), Liverpool, UK, 2003, page ?? ????, ????, 2003. ISBN ???? LCCN ????

Jacobs:2003:AGB

[JLS^{+}03b] Charles Jacobs, Wilmot Li, Evan Schrier, David Bargeron, and David Salesin. Adaptive gridbased document layout. ACM Transactions on Graphics, 22(3):838–847, July 2003. CODEN AT-GRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Jacobs:2004:ADL

[JLS^{+}04] Charles Jacobs, Wil Li, Evan Schrier, David Bargeron, and David Salesin. Adaptive document layout. Communications of the Association for Computing Machinery, 47(8):60–66, August 2004. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Johari:1996:AYP

[JMPS96] Ramesh Johari, Joe Marks, Ali Partovi, and Stuart Shieber. Automatic Yellow-Pages pagination and layout. Technical Report TR-96-29, Mitsubishi Electric Research Laboratory, 201 Broadway, Cambridge, MA 02139, USA, December 1996. 29 pp. URL

http://www.merl.com/
publications/TR1996-
029/.

Johari:1997:AYP

[JMPS97] Ramesh Johari, Joe Marks, Ali Partovi, and Stuart Shieber. Automatic Yellow-Pages pagination and layout. Journal of Heuristics, 2(4):321–342, ???? 1997. CODEN JO-HEFD. ISSN 1381-1231 (print), 1572-9397 (electronic). URL http:

//link.springer.com/
article/10.1007/BF00132503.

Johnstone:1992:LC

[Joh92] Adrian Johnstone. L^{A}TEX, Concisely. Ellis Horwood, New York, NY, USA, 1992. ISBN 0-13-524539-7 (paperback). xviii + 170 pp. LCCN Z253.4.L38 J64 1992. US$30.00.

Jackowski:1992:PTR

[JR92] Boguslav Jackowski and Market Ryćko. Polishing TEX: from ready to use to handy in use. In Zlatuška [Zla92], pages 119–134. ISBN 80-210-0480-0.

Jurriens:1992:TE

[Jur92] Theo Jurriens. TEX for everybody? In Zlatuška [Zla92], page 86. ISBN 80-210-0480-0.

Kanai:1991:TLE

[Kan91] J. Kanai. Text line extraction and baseline detection. In A. Lichnerowicz, editor, Intelligent Text and Image Handling: Proceedings of a Conference on Intelligent Text and Image Handling ‘RIAO 91’, Barcelona, Spain, 2–5 April 1991, pages 194–210. Elsevier Science Publishers B.V., Amsterdam, The Netherlands, April 1991. ISBN 0-444-89361-X. LCCN QA76.9.T48 C66 199.

Karow:1986:DSS

[Kar86]  Peter Karow.  Dig-

itale Speicherung von Schriften. URW (Unternehmensberatung Rubow Weber) Verlag, Hamburg, Germany, 1986. ISBN 3-926515-00-7. 331 pp.

Karow:1987:DFT

[Kar87] Peter Karow. Digital Formats for Typefaces. URW (Unternehmensberatung Rubow Weber) Verlag, Hamburg, Germany, 1987. ISBN 3-926515-01-5. 400 pp. LCCN Z253.3 .K371 1987.

Karow:1989:SND

[Kar89] Peter Karow. Shotai-no-dijitaru-fomatto. URW (Unternehmensberatung Rubow Weber) Verlag, Hamburg, Germany, 1989. ISBN 3-926515-00-7 (Hamburg), 4-946454-01-2 (Japan). 436 pp. LCCN ???? In Japanese.

Karow:1992:DSD

[Kar92a]  Peter Karow.  Digi-

tale Schriften, Darstellung und Formate, Geleitwort von Hermann Zapf (German) [Digital Fonts, Representation and Formats, Forward by Hermann Zapf]. Springer-Verlag and URW (Unternehmensberatung Rubow Weber) Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc. and Hamburg, Germany, 1992. ISBN 3-540-54917-X. xiii + 457, with 230 illustrations pp. LCCN ????

Karow:1992:S

[Kar92b] Peter Karow. Schriftstatistik. URW (Unternehmensberatung Rubow Weber) Verlag, Hamburg, Germany, 1992. ISBN 3-926515-07-4. 347 pp. LCCN ????

Karow:1992:SMW

[Kar92c] Peter Karow. Schrifttechnologie: Methoden und Werkzeuge. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1992. ISBN 3-540-54918-8. xvi + 460 pp. LCCN ????

Karow:1994:DTD

[Kar94a] Peter Karow. Digital Typefaces: Description and Formats. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1994. ISBN 3-540-56509-4. xiii + 448 pp. LCCN Z250.7 .K37 1994. US$89.00.

Karow:1994:FTM

[Kar94b] Peter Karow. Font Technology: Methods and Tools. Springer-Ver-lag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1994. ISBN 0-387-57223-6 (New York), 3-540-57223-6 (Berlin). xvi + 457 pp. LCCN Z250.7.K3813 1994. US$79.00. Foreword by Gerard Unger.

Karow:1994:UEL

[Kar94c]  Peter Karow.  URW

EuroWorks: 500 Latin fonts for Europe. URW (Unternehmensberatung Rubow Weber) Verlag, Hamburg, Germany, 1994. ISBN 3-926515-11-

2. 216 pp. LCCN ????

Katzenbeisser:1993:ET

[Kat93] Stefan Katzenbeisser.

Einführung in TEX. R. Oldenbourg Verlag, Munich, Germany and Vienna, Austria, second edition, 1993. ISBN 3-486-22682-7. 108 pp. 29,80 DM.

Katzenbeisser:1997:IDE

[Kat97] Stefan Katzenbeisser.

Von der Idee zum Dokument: Einführung in TEX und L^{A}TEX. R. Oldenbourg Verlag, Munich, Germany and Vienna, Austria, second edition, 1997. ISBN 3-486-24182-

6. 222 pp. DM 68,00.

Kernighan:1975:STM

[KC75] Brian W. Kernighan and Lorinda L. Cherry. A system for typesetting mathematics. Communications of the Association for Computing Machinery, 18(3):151– 157, 1975. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Kopka:1993:GLD

[KD93] Helmut Kopka and Patrick W. Daly. A Guide to L^{A}TEX: Document Preparation for Beginners and Advanced Users. Addison-Wesley, Reading, MA, USA, 1993. ISBN 0-201-56889-6. xvi + 436 pp. LCCN Z253.4.L38 K66 1993. US$34.50.

Kopka:1995:GDP

[KD95] Helmut Kopka and Patrick W. Daly. A Guide to L^{A}TEX2e: Document Preparation for Beginners and Advanced Users. Addison-Wesley, Reading, MA, USA, second edition, 1995. ISBN 0-201-42777-X. x + 554 pp. LCCN Z253.4.L38 K66 1995. US$37.61.

Kopka:1999:GLD

[KD99] Helmut Kopka and Patrick W. Daly. A Guide to L^{A}TEX 2ε: Document Preparation for Beginners and Advanced Users. Addison-Wes-ley, Reading, MA, USA, third edition, 1999. ISBN 0-201-39825-7. xv + 600 pp. LCCN Z253.4.L38 K66 1999. US$44.95.

Kopka:2004:GLT

[KD04] Helmut Kopka and Patrick W.

Daly. Guide to L^{A}TEX: Tools and Techniques for Computer Typesetting. Addison-Wes-ley, Reading, MA, USA, fourth edition, 2004. ISBN 0-321-17385-6 (paperback). xii + 597 pp. LCCN Z253.4.L38 K66 2004. US$49.99, CAN$75.99. Includes CD-ROM.

Knuth:2013:EK

[KD13] Donald Ervin Knuth and Edgar G. Daylight. Essential Knuth. Lonely Scholar Scientific Books, Geel, Belgium, 2013. ISBN 94-91386-03-4. 94 pp. LCCN ????

Knuth:2015:ABF

[KD15] Donald E. Knuth and Edgar G. Daylight. Algorithmic Barriers Falling: P = NP? Conversations. Lonely Scholar Scientific Books, Geel, Belgium, 2015. ISBN 94-91386-04-2. xix + 116 pp. LCCN QA76.9.A43 K58 2014.

Keller:1985:TA

[Kel85] Arthur M. Keller. Typesetting by authors. In Lucarella [Luc85b], pages 1–14. ISBN 0-201-13399-7. LCCN Q224.E97 1985.

Kernighan:1981:BRM

[Ker81a] B. W. Kernighan. Book review: “TEX & metafont: New directions in typesetting”. Computing Reviews, 1981. CODEN CPGRA6. ISSN 0010-4884, 0149-1202.

Kernighan:1981:PLT

[Ker81b] B. W. Kernighan. PIC: a language for typesetting graphics. In ACM-SIGOA’81 [ACM81], pages 92–96. ISBN 0-89791-043-5. LCCN QA76.7 .S54 v.16:6.

Kernighan:1981:TMT

[Ker81c] B. W. Kernighan. Typesetting mathematics, tables, and diagrams. In IBM Conference on Text Processing. ????, Bad Neuenahr, Germany, September 1981. Reprinted in Oldenbourg Verlag, Munich, 1982.

Kernighan:1982:PLT

[Ker82] B. W. Kernighan. PIC: a language for typesetting graphics. Software —Practice and Experience, 12(1):1–21, January 1982. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Kernighan:1984:UDP

[Ker84] B. W. Kernighan. The Unix document preparation tools — A retrospective. In Miller [Mil84], pages vii + 59. ISBN 0-906783-41-0 (hardcover), 0-906783-42-9 (paper). LCCN QA76.9.T48 I591 1984.

Kernighan:1985:RWU

[Ker85] B. W. Kernighan. Recent work in Unix document preparation tools. In Proceedings of the European UNIX Users Group Conference, September, Copenhagen, Denmark, page ?? European UNIX Users Group, Buntingford, Herts, UK, September 1985.

Kernighan:1989:USD

[Ker89] B. W. Kernighan. The UNIX system document preparation tools: a retrospective. AT&T Technical Journal, 68(4):5– 20, August 1989. CO-DEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Kernighan:1990:ITD

[Ker90] B. W. Kernighan. Issues and tradeoffs in document preparation systems. In Furuta [Fur90], pages 1–16. ISBN 0-521-40246-8. LCCN Z 286 E43 I58 1990.

King:2005:DPA

[Kin05] Peter R. King, editor. Doceng ’05: Proceedings of the 2005 ACM symposium on Document Engineering: November 2– 4, 2005, Bristol, United Kingdom. ACM Press, New York, NY, USA, 2005. ISBN 1-59593-240-2. LCCN HF5736 .A27 2005. ACM order number 605054.

Kingston:2006:FDF

[Kin06] J. H. Kingston. The future of document formatting. Technical report, Basser Department of Computer Science, University of Sydney, Sydney, NSW, Australia, 2006.

Kirschenbaum:2005:GGH

[Kir05] Valerie Kirschenbaum. Goodbye Gutenberg: hello to a new generation of readers and writers. Designer Writers series. Global Renaissance Society, New York, NY, USA, 2005. ISBN 0-9745750-3-8. 416 pp. LCCN Z116.A3 K57

                                        2005. URL http://www.
                                        goodbyegutenberg.com/

.

Kim:1989:MFG

[KKL89] Cha Jong Kim, Han Jong Kang, and Soo Youn Lee. A multipurpose font generation system. Journal of the Korea Information Science Society = Chongbo Kwahakhoe nonmunji, 16 (5):422–433, September 1989. CODEN HJKHDC. ISSN 0258-9125.

Kernighan:1982:UDP

[KL82] B. W. Kernighan and M. E. Lesk. UNIX document preparation. In Nievergelt et al. [NCNS82], pages 1– 20. ISBN 0-444-86493-8. LCCN Z244 .D63 1982. US$46.50.

Knuth:1993:CSS

[KL93] Donald E. Knuth and Silvio Levy. The CWEB System of Structured Documentation, Version 3.0. Addison-Wesley, Reading, MA, USA, 1993. ISBN 0-201-57569-8. 226 pp. LCCN QA76.9.D3 K6 1993.

Kernighan:1978:UTS

[KLO78a] B. W. Kernighan, M. E.

Lesk, and J. F. Ossanna. UNIX time-sharing system: Document preparation. The Bell System Technical Journal, 57(6): 2115–2135, 1978. CO-DEN BSTJAN. ISSN 0005-8580.

Kernighan:1978:DP

[KLO78b] B. W. Kernighan, M. E. Lesk, and J. F. Ossanna, Jr. Document preparation. The Bell System Technical Journal, 57(2): 2115–2135, 1978. CO-DEN BSTJAN. ISSN 0005-8580.

Klockl:2002:LTT

[Klö02] Ingo Klöckl. L^{A}TEX — Tipps und Tricks: Layoutanpassung, Programmierung, Grafik, Hilfsprogramme, Zeichensätze. dpunkt/PRO, ????, second edition, 2002. ISBN 3-89864-145-7. x + 606 pp. LCCN ???? 80 SFR, 49 EUR.

Knuth:1989:MW

[KLR89] Donald E. Knuth, Tracy Larrabee, and Paul M. Roberts. Mathematical Writing. Notes Number 14. Mathematical Association of America, Washington, DC, USA, 1989. ISBN 0-88385-063-X. Khodulev:1992:TEM

[KM92] Andrei B. Khodulev and Irina A. Makhovaya. On TEX experience in Mir Publishers. In Zlatuška [Zla92], pages 37–42. ISBN 80-210-0480-0.

Knauff:2014:ECD

[KN14] Markus Knauff and Jelica Nejasmic. An efficiency comparison of document preparation systems used in academic research and development. PLoS One, 9(12): e115069, December 2014. CODEN POLNCL. ISSN 1932-6203. See correction [Ano15].

Knappen:1992:CAM

[Kna92] Jörg Knappen. Changing the appearance of math. In Zlatuška [Zla92], pages 212–216. ISBN 80-210-0480-0.

Knappen:1997:SAZ

[Kna97] Jörg Knappen. Schnell ans Ziel mit L^{A}TEX 2ε. R. Oldenbourg Verlag, Munich, Germany and Vienna, Austria, 1997. ISBN 3-486-24199-0. xx + 187 pp. DM 48,00.

Knuth:1978:MT

[Knu78a] Donald E. Knuth. Mathematical typography. Technical Report STAN-CS-78-648, Stanford University, Stanford, CA, USA, February 1978. Also published as [Knu79b].

Knuth:1978:TEC

[Knu78b] Donald E. Knuth. TAU EPSILON CHI. A system for technical text. Report STAN-CS-78-675, Stanford University, Department of Computer Science, Stanford, CA, USA, 1978.

Knuth:1979:TMN

[Knu79a] Donald E. Knuth. TEX and : New Directions in Typesetting. Digital Press, 12 Crosby Drive, Bedford, MA 01730, USA, 1979. ISBN 0-932376-02-9. xi + 201 + 105 pp. LCCN Z253.3 .K58 1979.

Knuth:1979:MT

[Knu79b] Donald E. Knuth. Mathematical typography. Bulletin of the American Mathematical Society (new series), 1:337– 372, March 1979. Also published as a Stanford CS Report, STAN-CS-78-648.

Knuth:1979:MSA

[Knu79c] Donald E. Knuth. Metafont — A system for alphabet design. Report STAN-CS-79-762, Stanford University, Department of Computer Science, Stanford, CA, USA, 1979.

Knuth:1979:LS

[Knu80a] Donald E. Knuth. The letter S. The Mathematical Intelligencer, 2 (3):114–122, 1979, 1980. Also published as [Knu80c].

Knuth:1980:CMF

[Knu80b] Donald E. Knuth. The Computer Modern family of typefaces. Report STAN-CS-80-780, Stanford University, Department of Computer Science, Stanford, CA, USA, 1980.

Knuth:1980:LS

[Knu80c] Donald E. Knuth. The letter S. Technical Report STAN-CS-80-795, Stanford University, Stanford, CA, USA, April 1980. Also published as [Knu80a].

Knuth:1981:CMF

[Knu81] Donald E. Knuth. The concept of a meta-font. Technical Report STAN-CS-81-886, Stanford University, Stanford, CA, USA, October 1981.

Knuth:1982:T

[Knu82a] Donald E. Knuth. TEX82.

Stanford University, Stanford, CA, USA, 1982.

Knuth:1982:CMF

[Knu82b] Donald E. Knuth. The concept of a Meta-Font. Visible Language, 16(1): 3–27, January 1982. CO-DEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL

https://s3-us-west-
2.amazonaws.com/visiblelanguage/
pdf/V16N1_1982_E.pdf.

See comments [Hof82]. French translation by M. R. Delorme in Communication et Langages no. 55, (1983), 40–53; reprinted in Typographie et Informatique, proceedings of INRIA conference held 21–25 January 1985, organized by Jacques André and Patrick Sallio (Rennes, France: INRIA/IRISA – CCETT, 1985), 119–132.

Knuth:1982:TFT

[Knu82c] Donald E. Knuth. A test file for TEX. Stanford University, Stanford, CA, USA, 1982.

Knuth:1982:TV

[Knu82d] Donald E. Knuth. TEXware (version 1). Stanford University, Stanford, CA, USA, 1982.

Knuth:1982:WSS

[Knu82e] Donald E. Knuth. The WEB System of Structured Documentation— Version 1. Stanford University, Stanford, CA, USA, 1982.

Knuth:1983:LLM

[Knu83a] Donald E. Knuth. Lessons learned from - . Report STAN-CS-83-978, Stanford University, Department of Computer Science, Stanford, CA, USA, 1983.

Knuth:1983:LP

[Knu83b] Donald E. Knuth. Literate programming. Report STAN-CS-83-981, Stanford University, Department of Computer Science, Stanford, CA, USA, 1983.

Knuth:1983:WSS

[Knu83c] Donald E. Knuth. The WEB System of Structured Documentation— ver.2.3. Report STAN-CS-83-980, Stanford University, Department of Computer Science, Stanford, CA, USA, 1983.

Knuth:1984:TB

[Knu84a] Donald E. Knuth. The TEXbook, volume A of Computers and Typesetting. Addison-Wes-ley, Reading, MA, USA, 1984. ISBN 0-201-13447-0 (hardcover), 0-201-13448-9 (paperback). ix + 483 pp. LCCN Z253.4.T47 K58 1986. US$15.95 (paperback), US$32.95 (hardcover). Second printing, revised, October 1984. Sixth printing, revised, January 1986; also published as Computers & Typesetting, Vol. A. Twentyfourth printing, summer 1993, contains final revisions. Twenty-first printing, June 1992.

Knuth:1984:LP

[Knu84b] Donald E. Knuth. Literate programming. The Computer Journal, 27 (2):97–111, May 1984. CODEN CMPJA6.

Knuth:1984:TTT

[Knu84c] Donald E. Knuth. A torture test for TEX. Report STAN-CS-84-1027, Stanford University, Department of Computer Science, Stanford, CA, USA, 1984.

Knuth:1985:CM

[Knu85] Donald E. Knuth. Le concept de Metafonte. In André [And85], pages 119–132. Przedruk z “Communication et langages” no. 55 (1983) pp. 40–53.

Knuth:1986:MP

[Knu86a]  Donald E. Knuth.  -

: The Program, volume D of Computers and Typesetting. Addison-Wesley, Reading, MA, USA, 1986. ISBN 0-201-13438-1. xv + 560 pp. LCCN Z250.8.M46 K578 1986.

Knuth:1986:TP

[Knu86b] Donald E. Knuth. TEX: The Program, volume B of Computers and Typesetting. Addison-Wes-ley, Reading, MA, USA, 1986. ISBN 0-201-13437-3. xv + 594 pp. LCCN Z253.4.T47 K578 1986.

Knuth:1986:CMT

[Knu86c] Donald E. Knuth. Computer Modern Typefaces, volume E of Computers and Typesetting. Addison-Wesley, Reading, MA, USA, 1986. ISBN 0-201-13446-2. xv + 588 pp. LCCN Z250.8.M46 K574 1986.

Knuth:1986:TTT

[Knu86d] Donald E. Knuth. A torture test for . Report STAN-CS-86-1095, Stanford University, Department of Computer Science, Stanford, CA, USA, 1986.

Knuth:1986:T

[Knu86e] Donald E. Knuth. TEXware.

Report STAN-CS-86-1097, Stanford University, Department of Computer Science, Stanford, CA, USA, April 1986.

Knuth:ct-c

[Knu86f] Donald E. Knuth. The book, volume C of Computers and Typesetting. Addison-Wesley, Reading, MA, USA, 1986. ISBN 0-201-13445-4 (hardcover), 0-201-13444-6 (paperback). xi + 361 pp. LCCN Z250.8.M46 K58

Knuth:1987:DHD

[Knu87] Donald E. Knuth. Digital halftones by dot diffusion. ACM Transactions on Graphics, 6(4):245– 273, October 1987. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Knuth:1988:ET

[Knu88] Donald E. Knuth. The errors of TEX. Technical Report STAN-CS-88-1223, Stanford University, Department of Computer Science, Stanford, CA, USA, September 1988.

Knuth:1989:TB

[Knu89a] Donald E. Knuth. The TEXbook. ASCII Corporation, Tokyo, Japan, 1989. ISBN 4-7561-0010-4. Translation of the TEXbook into Japanese, with addition of Appendix K, and new drawings by Duane Bibby.

Knuth:1989:CAG

[Knu89b] Donald E. Knuth. Calling all grand wizards. TEXhax, 89(98), November 1989.

Knuth:1989:ET

[Knu89c] Donald E. Knuth. The errors of TEX. Software—Practice and Experience, 19(7):607–681, July 1989. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). This is an updated version of [Knu88].

Knuth:1990:VFMa

[Knu90a]  Donald Knuth.  Vir-

tual fonts: More fun for grand wizards. TUGboat, 11(1):13–23, April 1990. ISSN 0896-3207.

Knuth:1990:SPW

[Knu90b] Donald E. Knuth. A simple program whose proof isn’t. In Feijen et al. [FvGGM90], chapter 27, pages 233– 242. ISBN 0-387-97299-4, 3-540-97299-4, 1-4612-8792-8 (print), 1-4612-4476-5 (online). ISSN 0172-603X. LCCN QA76 .B326 1990. URL https:

//link.springer.com/
chapter/10.1007/978-
1-4612-4476-9_28. This

paper discusses the algorithm used in TEX for converting between decimal and scaled fixedpoint binary values, and for guaranteeing a minimum number of digits in the decimal representation. See also [?] for decimal to binary conversion, [?] for binary to decimal conversion, and [Gri90] for an alternate proof of Knuth’s algorithm.

Knuth:1990:VFMb

[Knu90c] Donald E. Knuth. Virtual fonts: More fun for grand wizards. TEXhax, 90(11 and 12), January 1990.

Knuth:1991:TPC

[Knu91] D. E. Knuth. Theory and practice (computer science). Theoretical Computer Science, 90(1):1– 15, November 1991. CO-DEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Knuth:1992:LP

[Knu92] Donald E. Knuth. Literate Programming, volume 27 of CSLI Lecture Notes. Stanford University Center for the Study of Language and Information, Stanford, CA, USA, 1992. ISBN 0-937073-80-6 (paperback), 0-937073-81-4 (hardcover). xv + 368 pp. LCCN QA76.6.K644.

                                        URL http://csli-www.
                                        stanford.edu/publications/
                                        literate.html.

Knuth:1993:VT

[Knu93] Donald E. Knuth. Vse pro TEX [English: Everything about TEX]. Protvino, Moscow, Russia, 1993. ISBN 5-900614-01-8. xvi + 575 pp. Russian translation of [Knu84a], by M. V. Lisina, edited by S. V. Klimenko and S. N. Sokolov. AO RDTEX (AO means Inc; RDTEX stands for Russian Database TEXnology). First press run: 10000 copies.

Knuth:1999:DT

[Knu99] Donald E. Knuth. Digital Typography, volume 78 of CSLI Lecture Notes. CSLI Publications, Stanford, CA, USA, 1999. ISBN 1-57586-010-4 (paperback), 1-57586-011-2 (hardcover). xvi + 685 pp. LCCN Z249.3.K59 1998. US$90.00 (cloth), US$39.95 (paperback).

Kohlhase:2008:USM

[Koh08] Michael Kohlhase. Using L^{A}TEX as a semantic markup format. Mathematics in Computer Science, 2(2):279–304, December 2008. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic). URL http:

//www.springerlink.
com/openurl.asp?genre=
article&issn=1661-8270&
volume=2&issue=2&spage=
279.

Kolodziejska:1987:DWP

[Koł87] Hanna Kołodziejska. Dzielenie wyrazów polskich w systemie TEX. Technical Report 165, Sprawozdania Instytutu Informatyki Uniwersytetu Warszawskiego, Warszawa, Poland, 1987.

Kolodziejska:1988:T

[Koł88a] Hanna Kołodziejska. TEX. Mikrolan, 1:12, 1988.

Kolodziejska:1988:TTP

[Koł88b] Hanna Kołodziejska. Le traitement des textes polonais avec le logiciel TEX. Cahiers GUTenberg, 0:3–10, April 1988. ISSN 1140-9304.

Kolda:2024:ULG

[Kol24] Tamara G. Kolda. Unlocking L^{A}TEX Graphics: a Concise Guide to TikZ. MathSci.ai, Dublin, CA, USA, 2024. iv + 120 pp. LCCN

                                        ????  URL https:/
                                        /epubs.siam.org/doi/
                                        book/10.1137/1.9798991229500;
                                         https://tikz.dev/pgfplots/

.

Koning:2001:CDL

[Kon01] Ruud H. Koning. A comparison of different L^{A}TEX programs. Journal of Applied Econometrics, 16(1):81–92, January– February 2001. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Kopka:1988:LE

[Kop88] Helmut Kopka. L^{A}TEX: Eine Einführung. Addison-Wesley Verlag, Bonn, Germany, second edition, 1988. ISBN 3-89319-136-4, 3-89319-199-2. xvi + 340 pp.

Kopka:1990:LEE

[Kop90] Helmut Kopka. L^{A}TEX —Erweiterungsmöglichkeiten: mit einer Einführung in METAFONT. Addison-Wesley Verlag, Bonn, Germany, 1990. ISBN 3-89319-287-5. xv + 463 pp. LCCN ????

Kopka:1991:EWE

[Kop91a] Helmut Kopka. L^{A}TEX —Erweiterungsmöglichkeiten mit einer Einfürung in . Addison-Wesley Verlag, Bonn, Germany, second edition, 1991. ISBN 3-89319-356-1. 552 pp.

Kopka:1991:LE

[Kop91b] Helmut Kopka. L^{A}TEX— Eine Einführung. Addison-Wesley Verlag, Bonn, Germany, third edition, 1991. ISBN 3-89319-338-3. xvii + 375 pp.

Kopka:1992:LE

[Kop92] Helmut Kopka. L^{A}TEX— Eine Einführung. Addison-Wesley Verlag, Bonn, Germany, fourth edition, 1992. ISBN 3-89319-434-7. xvii + 445 pp.

Kopka:1994:LBE

[Kop94] Helmut Kopka. L^{A}TEX Band 1: Einführung. Addison-Wesley Verlag, Bonn, Germany, 1994. ISBN 3-89319-664-1. xix + 428 pp. Replaces the old series of two volumes. Kopka:1995:LBEa

[Kop95] Helmut Kopka. L^{A}TEX Band 2: Ergänzungen— mit einer Einfürung in . Addison-Wesley Verlag, Bonn, Germany, 1995. ISBN 3-89319-665-X. xiv + 428 pp. DM69.90.

Kopka:1997:LBE

[Kop97] Helmut Kopka. L^{A}TEX Band 3: Erweiterungen. Addison-Wesley Verlag, Bonn, Germany, 1997. ISBN 3-89319-666-8. xiii + 412 pp. DM69.90.

KTS:2007:AJT

[Kor07] Korean TEX Society. The Asian journal of TEX. Online journal in Chinese, English, Japanese, and Korean., 2007. ISSN 1976-1228. URL http:/

                                        /ajt.ktug.org/.

Kottwitz:2011:LBG

[Kot11] Stefan Kottwitz. L^{A}TEX beginner’s guide. Packt Publishing, Birmingham, UK, 2011. ISBN 1-84719-986-0, 1-84719-987-9 (ebook). xi + 314 pp.

LCCN Z253.4.L38 K67 2011.

Kottwitz:2015:LC

[Kot15] Stefan Kottwitz. L^{A}TEX Cookbook. Packt Publishing, Birmingham, UK, 2015. ISBN ????

????  pp.  LCCN

????  URL http://
latex-cookbook.net/.

Kottwitz:2023:LGT

[Kot23] Stefan Kottwitz. La-TeX graphics with TikZ: a practitioner’s guide to draw 2D and 3D images, diagrams, charts, and plots in L^{A}TEX using TikZ. Packt Publishing, Birmingham, UK, 2023. ISBN 1-80461-823-3 (paperback), 1-80461-704-0 (e-book). xviii + 279 pp. LCCN Z253.4.L38 K68 2023.

Knuth:1981:BPL

[KP81] Donald E. Knuth and Michael F. Plass. Breaking paragraphs into lines. Software—Practice and Experience, 11(11):1119– 1184, November 1981. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Kendall:1985:IOC

[KR85] M. Kendall and A. Reed.

Introduction of an online computer typesetting service. University Computing, 7(3):134– 138, Winter 1985. CO-DEN UNCOET. ISSN 0265-4385.

Krantz:1996:PMW

[Kra96] Steven G. Krantz. A Primer of Mathematical Writing: Being a disquisition on having your ideas recorded, typeset, published, read & appreciated. American Mathematical Society, Providence, RI, USA, 1996. ISBN 0-8218-0635-1. 216 pp. LCCN QA42.K73 1996. US$19.00. URL

http://www.ams.org/

bookstore. Includes discussions of computers and TEX.

Krantz:2001:HTM

[Kra01] Steven G. Krantz. Handbook of Typography for the Mathematical Sciences. Chapman and Hall/CRC, Boca Raton, FL, USA, 2001. ISBN 1-58488-149-6. xv + 173 pp. LCCN Z253.4.T47 K65 2001.

Kroenert:1990:IIS

[Krö90] Günther Krönert. Importance of the ISO standard 8613 for document interchange. Basic standard, functional standards, initial experience, benefits. In Nadrchal [Nad90], pages 209–218. Invited paper.

Kellerman:1987:TPT

[KS87] David Kellerman and Barry Smith. TEXTURES: Professional Typesetting for the Macintosh: User’s Guide. Addison-Wesley, Reading, MA, USA, 1987. ISBN 0-201-17221-6. 123 pp. LCCN Z253.4.T47 U85 1987.

Krieger:1990:IT

[KS90] Jost Krieger and Norbert Schwarz. Introduction to TEX. Addison-Wes-ley Europe, Amsterdam, The Netherlands, 1990. ISBN 0-201-51141-X. 278 pp. LCCN Z253.4.T47 K75 1990. English translation of [Sch88b]. See also the Dutch translation, [Sch90b].

Klein:2003:CSP

[KSW03] Rolf Klein, H.-W. (Hans-Werner) Six, and Lutz Wegner, editors. Computer Science in Perspective: Essays Dedicated to Thomas Ottmann, volume 2598 of Lecture Notes in Computer Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2003. CODEN LNCSD9. ISBN 3-540-00579-X (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). x + 355 pp. LCCN QA76 .C5476 2003. URL http:

                                        //link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t2598.htm;
                                         http://www.springerlink.
                                        com/content/978-3-540-
                                        00579-7;   http://www.
                                        springerlink.com/openurl.
                                        asp?genre=issue&issn=
                                        0302-9743&volume=2598.

Kernighan:1989:PMP

[KV89] Brian W. Kernighan and Christopher J. Van Wyk. Page makeup by postprocessing text formatter output. Computing Systems, 2(2):103– 132, Spring 1989. CO-DEN CMSYE2. ISSN 0895-6340.

Knuth:1990:NTM

[KW90] Donald E. Knuth and Joe Weening. New TEX/ sources available on Stanford’s master archive. TEXhax, 90(13), January 1990.

Kwong:1991:KKB

[Kwo91] Man Kam Kwong. kxp and kxpl: a busy man’s L^{A}TEX: Part I: An overview. Technical Report ANL-91/14, Argonne National Laboratory, 9700 South Cass Avenue, Argonne, IL 60439-4801, USA, March 1991.

Knuth:1989:AEN

[KZ89] Donald E. Knuth and Hermann Zapf. AMS euler—A new typeface for mathematics. Scholarly Publishing, 20(3):131– 157, April 1989. ISSN 0036-634X.

Loureiro:2005:CXF

[LA05] Gil Loureiro and Francisco Azevedo. Constrained XSL formatting objects for adaptive documents. In King [Kin05], pages 95–97. ISBN 1-59593-240-2. LCCN HF5736 .A27 2005. ACM order number 605054.

Labuz:1984:HTW

[Lab84] Ronald Labuz. How to Typeset from a Word Processor: a Practical Manual for Type Buyers and Typesetters. R. R. Bowker Company, New York, NY, USA, 1984. ISBN 0-8352-1899-6. xii + 218 pp. LCCN Z253.3.L33 1984. US$24.95.

Lagally:1992:ATA

[Lag92] Klaus Lagally. Arab-TEX — typesetting Arabic with vowels and ligatures. In Zlatuška [Zla92], pages 153–172. ISBN 80-210-0480-0.

Lagus:1995:APG

[Lag95]  Krista Lagus.  Au-

tomated pagination of the generalized newspaper using simulated annealing. Master’s thesis, Helsinki University of Technology, Helsinki, Finland, 1995.

Lamport:1986:LDP

[Lam86] Leslie Lamport. L^{A}TEX: A Document Preparation System: User’s Guide and Reference Manual. Addison-Wesley, Reading, MA, USA, 1986. ISBN 0-201-15790-X. xiv + 242 pp. LCCN Z253.4.L38 L35 1986.

Lamport:1991:L

[Lam91] Leslie Lamport. L^{A}TEX. Cyfronet, Kraków, 1991. x + 202 pp. Polish translation of “L^{A}TEX: A Document Preparation System”, 1986, by Piotr Wyrostek.

Lamport:1994:LDP

[Lam94] Leslie Lamport. L^{A}TEX: A Document Preparation System: User’s Guide and Reference Manual. Addison-Wes-ley, Reading, MA, USA, second edition, 1994. ISBN 0-201-52983-1. xvi + 272 pp. LCCN Z253.4.L38L35 1994.

Lamport:1995:DLH

[Lam95] Leslie Lamport. Das L^{A}TEX Handbuch. Addison-Wesley, Reading, MA, USA, 1995. ISBN 3-89319-826-1. xviii + 325 pp. DM69.90. German translation of [Lam94] by Rebecca Stiels.

Lamport:2004:LSO

[Lam04] Leslie Lamport. L^{A}TEX:

system opracowywania dokumentów: podrecznik i przewodnik uçzytkownika. Wydawnictwa Naukowo-Techniczne, Warszawa, 2004. ISBN 83-204-2878-5. xviii + 366 pp. Polish translation of “L^{A}TEX: a document preparation system”, 1994, by Marta Wolińska and Marcin Woliński.

Lao:1986:ST

[Lao86] Marek J. Lao. Sztuka TEXowania. Wiadomości Matematyczne, 27(1):81– 87, 1986. Roczniki Polskiego Towarzystwa Matematycznego, Seria II. In Polish.

Larsen:1989:LD

[Lar89] Steen Larsen. L^{A}TEXpå dansk. UNI•C, Danmarks EDB-Center for Forskning og Uddannelse, Copenhagen, Denmark, 1989. ISBN 87-7252-089-2.

Laurens:2004:IIT

[Lau04] Jérôme Laurens. iTEXmac:

An integrated TEX environment for Mac OS X. In Syropoulos et al.

[SBH^{+}04], pages 192–

202. CODEN LNCSD9. ISBN 3-540-22801-2 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN Z253.3 I58 2004. URL

http://link.springer-
ny.com/link/service/
series/0558/tocs/t3130.
htm;   http://www.springerlink.
com/openurl.asp?genre=
issue&issn=0302-9743&
volume=3130;   http:
//www.springerlink.
com/openurl.asp?genre=
volume&id=doi:10.1007/
b99374.

Lavaud:1992:SHE

[Lav92a] Michel Lavaud. A solution to help ensuring the future of TEX: make its use easier on cheap machines. In Zlatuška [Zla92], pages 66–68. ISBN 80-210-0480-0.

Lavaud:1992:WEF

[Lav92b] Michel Lavaud. A way to ensure the future of TEX: make its use easier on low-cost machines. In Zlatuška [Zla92], pages 307–330. ISBN 80-210-0480-0. Invited talk.

Lawson:1990:AT

[Law90] Alexander Lawson. Anatomy of a Typeface. Hamish Hamilton, London, 1990. ISBN 0-241-13267-3. 428 pp. LCCN Z250 .L34 1990.

Lee:1979:BIG

[Lee79] Marshall Lee. Bookmaking: the Illustrated Guide to Design/Production/ Editing. R. R. Bowker Company, New York, NY, USA, second edition, 1979. ISBN 0-8352-1097-9. 485 pp. LCCN Z116.A2 .L44 1979. US$25.00.

Leitch:1985:IMI

[Lei85] S. Leitch. Implementing on an ICL Perq. In Lucarella [Luc85b], pages 171– 182. ISBN 0-201-13399-7. LCCN Q224.E97 1985.

Leifeld:2013:TCS

[Lei13] Philip Leifeld. texreg:

Conversion of statistical model output in R to L^{A}TEX and HTML tables. Journal of Statistical Software, 55(8):??, November 2013. CO-DEN JSSOBK. ISSN 1548-7660. URL http:

//www.jstatsoft.org/
v55/i08.

Lesk:1988:GIL

[Les88] Michael Lesk. GRAB— inverted indexes with low storage overhead. Computing Systems, 1(3): 207–220, Summer 1988. CODEN CMSYE2. ISSN 0895-6340.

Levine:1990:LGR

[Lev90] Michael J. S. Levine. A L^{A}TEX graphics routine for drawing Feynman diagrams. Computer Physics Communications, 58(1 and 2):181– 198, February/March 1990. Available in the CPC library, and

from listral@ib.rl.

ac.uk with an e-mail

request GET CPC INTRO
CPCINDEX.

Levelt:1995:IPI

[Lev95] A. H. M. (Antonius Henricus Maria) Levelt, editor. ISSAC ’95: Proceedings of the 1995 International Symposium on Symbolic and Algebraic Computation: July 10–12, 1995, Montreal, Canada. ACM Press, New York, NY, USA, 1995. ISBN 0-89791-699-

9. LCCN QA 76.95 I59 1995. ACM order number: 505950.

Levin:2019:B

[Lev19] Roy Levin. Biography. In Malkhi [Mal19], pages 131–170. ISBN 1-4503-7271-6 (paperback), 1-4503-7270-8 (hardcover), 1-4503-7273-2. LCCN

????  URL https://
www.morganclaypoolpublishers.
com/catalog_Orig/product_
info.php?cPath=24&products_
id=1468.

Lok:2004:EVB

[LFN04] Simon Lok, Steven Feiner, and Gary Ngai. Evaluation of visual balance for automated layout. In Nuno Jardim Nunes and Charles Rich, editors, IUI 04: 2004 International Conference on Intelligent User Interfaces, Funchal, Madeira, Portugal, January 13– 16, 2004, pages 101–108. ACM Press, New York, NY, USA, 2004. ISBN 1-58113-815-6. LCCN QA76.9.H85 I573 200.

Li:1985:GSC

[Li85] Jiarong Li. Generation of Some Chinese Characters with Metafont. In Lucarella [Luc85b], pages 161–170. ISBN 0-201-13399-7. LCCN Q224.E97 1985.

Liang:1983:WHP

[Lia83] Franklin Mark Liang.

Word hy-phen-a-tion by com-pu-ter. Technical Report STAN-CS-83-977, Stanford University, Stanford, CA, USA, August 1983. URL http:

//www.tug.org/docs/
liang/.

Liang:1984:WHP

[Lia84] Franklin Mark Liang.

Word Hy-phen-a-tion by Com-pu-ter. Ph.D. dissertation, Computer Science Department, Stanford University, Stanford, CA, USA, March 1984. 92 pp. URL

http://www.tug.org/
docs/liang/;   http:
//wwwlib.umi.com/dissertations/
fullcit/8329742.

Likos:2004:JBC

[Lik04] Johannis Likos. μoνo2πoλυ: Java-based conversion of monotonic to polytonic Greek. In Syropoulos et al. [SBH^{+}04], pages 34–54. CO-DEN LNCSD9. ISBN 3-540-22801-2 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN Z253.3 I58 2004. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t3130.htm;
 http://www.springerlink.
com/openurl.asp?genre=
issue&issn=0302-9743&
volume=3130;   http:
//www.springerlink.
com/openurl.asp?genre=
volume&id=doi:10.1007/
b99374.

Lin:2005:ADL

[Lin05]  Xiaofan Lin.  Active

document layout synthesis. In IEEE, editor, Proceedings. Eighth International Conference on Document Analysis and Recognition, 2005, pages 86–90. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 2005. ISBN 0-7695-2420-6. ISSN 1520-5263. LCCN TA1640 .I57 2005.

Lin:2006:ALE

[Lin06]  Xiaofan Lin.  Active

layout engine: Algorithms and applications in variable data printing. Computer-Aided Design, 38(5):444–456, May 2006. CODEN CAIDA5. ISSN 0010-4485 (print), 1879-2685 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0010448505001958.

Lipkin:1999:LLV

[Lip99] Bernice Sacks Lipkin.

L^{A}TEX for Linux: a Vade Mecum. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1999. ISBN 0-387-98708-8 (paperback). xxxi + 568 pp. LCCN Z253.4 L38 L56 1999. US$49.95.

Liu:2013:LRM

[Liu13] Haiyang ‘Leo’ Liu. L^{A}TEX Ru men. (Mandarin Chinese) [Introduction to L^{A}TEX]. Dian zi gong ye chu ban she, Beijing, China, 2013. ISBN 7-121-20208-5. xiv + 566 pp. LCCN ????

Lesk:1976:CTT

[LK76] M. E. Lesk and B. W. Kernighan. Computer typesetting of technical journals on UNIX. In AFIPS NCC ’76 [AFI76], pages 879–888. ISBN ?? LCCN ??

Lin:1994:CMF

[LL94] Y.-J. Lin and T.-L. Lee. Conversion from - files to TrueType. In Anonymous [Ano94], pages 65–75. ISSN 0218-7019.

Level:1995:PTF

[LNN95] Jeff Level, Bruce Newman, and Brenda Newman. Precision type font reference guide, version 5.0. Precision Type, Commack, NY, USA, 1995. ISBN 0-9646252-0-2. xxv + 653 pp. LCCN Z250.7 .P74 1995.

Level:2000:PTF

[LNN00] Jeff Level, Bruce Newman, and Brenda Newman. Precision Type Font Reference Guide, Version 5.0. Hartley & Marks Publishers, Point Roberts, WA, USA and Vancouver, BC, Canada, 2000. ISBN 0-88179-182-2. xxv + 653 pp. LCCN Z250.7 .P74 2000. US$39.95, CDN$59.95. Republication of [LNN95].

Lozano:2013:TCQ

[Loz13] Vincent Lozano. Tout ce que vous avez toujours voulu savoir sur L^{A}TEX sans jamais oser le demander, ou Comment utiliser L^{A}TEX quand on n’y connaît goutte. (French) [Everything you always wanted to know about L^{A}TEX without ever daring to ask, or How use L^{A}TEX when you don’t know a bit], volume 5 of Framabook. Framasoft, Lyon, 2013. 304 pp. LCCN ???? URL http:

//catalogue.bnf.fr/
ark:/12148/cb437602764.

Lachand-Robert:1995:MTL

[LR95] Thomas Lachand-Robert.

La maitrise de TEX et L^{A}TEX. Masson Editeur, Paris, France, 1995. ISBN 2-225-84832-7. 644 pp. LCCN ????

Leitch:1985:GAD

[LS85] S. Leitch and F. J.

Smith. Generalised Algorithm for Drawing Non-Parametric Splines. In Lucarella [Luc85b], pages 155–161. ISBN 0-201-13399-7. LCCN Q224.E97 1985.

Lammarsch:1996:CTL

[LS96] Joachim Lammarsch and Harald Schoppmann. CTAN-3: das TEX- /L^{A}TEX-Archiv von DANTE e.V. Addison-Wesley, Reading, MA, USA, 1996. 123 pp. DM 49,90. Includes CD-ROM.

Loverdos:2004:DTN

[LS04] Christos K. K. Loverdos and Apostolos Syropoulos. Digital typography in the new millennium: Flexible documents by a flexible engine. In Syropoulos et al. [SBH^{+}04], pages 1–16. CODEN LNCSD9. ISBN 3-540-22801-2 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN Z253.3 I58 2004. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t3130.htm;
 http://www.springerlink.
com/openurl.asp?genre=
issue&issn=0302-9743&
volume=3130;   http:
//www.springerlink.
com/openurl.asp?genre=
volume&id=doi:10.1007/
b99374.

Licha:1992:AV

[LU92a] Michaela Lichá and Oldřich Ulrych. AMS-TEX verze 2.1. CzechoslovakTEX Users Group, CSTUG MÚ UK, Sokolovská 83, 186 00 Praha 8, The Czech Republic, 1992. 80 pp. This publication provides a short (but complete) tour through AMS-TEX 2.1 together

with amsppt.sty 2.1.

Almost all features are demonstrated with a lot of examples. The book is organized in such a way that it is useful for novices (basic textbook), and for advanced users (quick reference guide) as well. Contains a table of all symbols.

Luvisetto:1992:LER

[LU92b] M. L. Luvisetto and E. Ugolini. L^{A}TEX: easy recipes. Technical Report INFN/TC-92/05, Ist. Naz. Fis. Nucl., Bologna, Italy, March 1992. 17 pp.

Lubkin:1990:PSP

[Lub90] Saul Lubkin. Porting sophisticated programs to your UNIX environment for free. Computer Shopper, pages 642–644, March 1990.

Lucarella:1985:TFD

[Luc85a] Dario Lucarella. TEX formulae dictionary. In TEXSCIDOC85 [Luc85b], pages 93–103. ISBN 0-201-13399-7. LCCN Q224.E97 1985.

Lucarella:1985:PFE

[Luc85b] Dario Lucarella, editor.

Proceedings of the First European Conference on TEX for Scientific Documentation, 16–17 May 1985, Como, Italy. Addison-Wesley, Reading, MA, USA, 1985. ISBN 0-201-13399-7. LCCN Q224.E97 1985.

Luna:1992:UTD

[Lun92] Paul Felix John Luna. Understanding Type for Desktop Publishing. Blueprint, London, UK, 1992. ISBN 0-948905-76-X. 136 pp. LCCN Z250.7 .L86 1992. Highly regarded English typographer: designed the Oxford English Dictionary and the New English Bible.

Lvovskii:1994:NVP

[L’v94] Sergei L’vovskii. Nabor i verstka v pakete L^{A}TEX [English: Typing and proving in the L^{A}TEX package]. Kosmosinform, Moscow, Russia, 1994. ISBN 5-900242-11-0. 328 pp. The first complete description of L^{A}TEX 2.09 in Russian. First press run: 5000 copies.

Larson:2023:MAD

[LV23] Eric Larson and Isabel Vogt. Making accessible documents using L^{A}TEX. Notices of the American Mathematical Society, 70(1):68– 71, January 2023. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic). URL

https://www.ams.org/
journals/notices/202301/
rnoti-p68.pdf.

Linton:1989:CUI

[LVC89] Mark A. Linton, John M.

Vissides, and Paul R. Calder. Composing user interfaces with Inter-Views. IEEE Computer, 22(2):8–22, February 1989. From Computing Reviews: “Inter-Views objects are combined through a model similar to the boxes and glues in Knuth’s TEX, which allows interface objects to be combined with little concern for undesired interactions.”.

MacKay:1983:TAS

[Mac83] Pierre A. MacKay. TEX for Arabic script. In Burton and Short [BS83], pages 391–400. ISBN 0-914894-96-X. LCCN AZ105 .I56 1983.

Madsen:2008:ITL

[Mad08] Lars Madsen. Introduktion til L^{A}TEX. Corpus Daleif, ????, 2008. ISBN ???? xx + 420 pp. LCCN ????

URL http://www.imf.
au.dk/system/latex/
bog/version3beta.html.

Magat:2011:HHC

[Mag11] Joan Ames Magat. Hawking hyphens in compound modifiers. Legal Communication and Rhetoric: JALWD, 11 (??):153–175, Fall 2011. ISSN 2328-0255. URL

https://www.alwd.org/
index.php?option=com_
attachments&task=download&
id=73.

Malkhi:2019:CWL

[Mal19] Dahlia Malkhi, editor. Concurrency: the Works of Leslie Lamport. Morgan and Claypool Publishers, Williston, VT 05495 USA, 2019. ISBN 1-4503-7271-6 (paperback), 1-4503-7270-8 (hardcover), 1-4503-7273-2. xx + 345 pp. LCCN ???? URL https:

//www.morganclaypoolpublishers.
com/catalog_Orig/product_
info.php?cPath=24&products_
id=1468.

Manas:1987:WDS

[Mañ87a] Jose A. Mañas. Word division in Spanish. Communications of the Association for Computing Machinery, 30(7):612– 616, July 1987. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Manson:1987:UEP

[Man87b] David Manson. Using TEX for electronic publishing at I. P. Sharp Associates. ACM SIG-DOC Asterisk*, 13(1):3– 7, March 1987. ISSN 0731-1001.

Maor:1994:SN

[Mao94] Eli Maor. e: The Story of a Number. Princeton University Press, Princeton, NJ, USA, 1994. ISBN 0-691-03390-0. xiv + 223 pp. LCCN QA247.5.M33 1994. URL http:/

/www-gap.dcs.st-and.
ac.uk/~history/HistTopics/

e.html. This book chronicles the story of the ultimate version number of .

Margolis:1994:MMG

[Mar94] Marvin S. Margolis. MG mathematical graphics system. Notices of the American Mathematical Society, 41(3):200–201, March 1994. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Maranget:1998:HLH

[Mar98]  Luc Maranget.  He-

VeA: L^{A}TEX to HTML translator. World-Wide Web document., 1998. URL http://

para.inria.fr/~maranget/

hevea/. The translator is written in Objective Caml.

Matsakis:1999:RHM

[Mat99] Nicholas E. (Nicholas Elias)

Matsakis. Recognition of handwritten mathematical expressions. Master’s thesis, Department of Electrical Engineering and Computer Science, Massachusetts Institute of Technology, Cambridge, MA, USA, 1999. 59 pp. URL http://

hdl.handle.net/1721.
1/16727.

Mathis:2005:CSV

[Mat05] R. M. Mathis. Constraint scalable vector graphics, accessibility and the Semantic Web. In Yair Levy, editor, Proceedings of the IEEE Southeast-Con 2005: “Excellence in Engineering, Science and Technology”: April 8–10, 2005, Ft. Lauderdale, FL, pages 588– 593. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 2005. ISBN 0-7803-8865-8. LCCN TK7801 .I117 2005.

Matthews:2019:CBE

[Mat19] David Matthews. Craft beautiful equations in Word with L^{A}TEX. Nature, 570(7760):263–264, June 13, 2019. CO-DEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic). URL

https://www.nature.
com/articles/d41586-
019-01796-1.

Mauerer:1997:TLU

[Mau97] Wolfgang Mauerer. Textverarbeitung mit L^{A}TEX 2ε unter UNIX. UNIX easy. Carl Hanser, München, Germany, 1997. ISBN 3-446-18909-2. xiv + 361 pp. LCCN ???? DM 49,80.

Morris:1991:HTV

[MBHL91] Robert A. Morris, Karl Berry, Kathryn A. Hargreaves, and Dimitrios Liarokapis. How typeface variation and typographic variation affect readability at small sizes. In Pietrowski [Pie91], pages 344–352.

Mamrak:1993:BAD

[MBO93] S. Mamrak, J. Barnes, and C. S. O’Connell. Benefits of automating data translation. IEEE Software, 10(4):82–88, July 1993. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

McClelland:1988:LTD

[McC88] Bruce McClelland. From logon to typeset document in under ten minutes. Online, 12(6):66– 71, November 1988. CO-DEN ONLIDN. ISSN 0146-5422.

McCarthy:2020:DTS

[McC20] Steven McCarthy. Digital typography at Stanford. She Ji: The Journal of Design, Economics, and Innovation, 6(4): 546–560, Winter 2020. ISSN 2405-8726. McKibben:1991:EEW

[McK91] L. McKibben. Equation editing with WordPerfect. Laboratory microcomputer, 10(3):117–118, 1991. CODEN LMICDI. ISSN 0262-2955.

McLean:1980:THM

[McL80]  Ruari McLean.  The

Thames and Hudson Manual of Typography. Thames and Hudson, London, UK, 1980. ISBN 0-500-68022-1. 216 pp. LCCN Z244.3 .M45.

McLean:1995:TTI

[McL95] Ruari McLean, editor. Typographers on Type: An Illustrated Anthology from William Morris to the present day. W. W. Norton & Co., New York, NY, USA, 1995. ISBN 0-393-70201. xii + 188 pp. LCCN Z 250 A2 T96 1995. US$27.00.

MF:1986:MCF

[MF86a] The MF Chel Family of fonts. The Metafoundry, OCLC Online Computer Library Center, Inc., Dublin, OH, USA, 1986.

MF:1986:MRF

[MF86b] The MF Roman Family of fonts. The Metafoundry, OCLC Online Computer Library Center, Inc., Dublin, OH, USA, 1986.

MF:1987:MDF

[MF87a] The MF Decorative Family of fonts. The Metafoundry, OCLC Online Computer Library Center, Inc., Dublin, OH, USA, 1987.

MF:1987:MSF

[MF87b] The MF Slavic Families of fonts. The Metafoundry, OCLC Online Computer Library Center, Inc., Dublin, OH, USA, 1987.

Mittelbach:2023:LCPb

[MFB^{+}23] Frank Mittelbach, Ulrike Fischer, Javier Bezos, Johannes Braams, and Joseph Wright. The LaTeX Companion: Part II. Tools and techniques for computer typesetting. Addison-Wesley, Reading, MA, USA, third edition, 2023. ISBN 0-201-36300-3 (Part II) (hardcover), 0-13-816657-9 (Part II) (PDF), 0-13-816652-8 (Part II) (ePub). xxxviii + 970 pp. LCCN Z253.4.L38 M58

Mittelbach:2023:LCPa

[MFW23] Frank Mittelbach, Ulrike Fischer, and Joseph Wright. The LaTeX Companion: Part I. Tools and techniques for computer typesetting. Addison-Wesley, Reading, MA, USA, third edition, 2023. ISBN 0-13-465894-9 (Part I) (hardcover), 978-013-465-9-5-9-6 (Part I) (PDF), 978-013-465-9-6-0-2 (Part I) (e-pub), 0-13-816648-0-X (Part I + II). xxviii + 947 pp. LCCN Z253.4.L38 M58 2023.

Mittelbach:2004:LC

[MGB^{+}04] Frank Mittelbach, Michel Goossens, Johannes Braams, David Carlisle, Chris Rowley, Christine Detig, and Joachim Schrod. The L^{A}TEX Companion. Tools and Techniques for Computer Typesetting. Addison-Wesley, Reading, MA, USA, second edition, 2004. ISBN 0-201-36299-6 (paperback), 0-321-51443-2 (hardcover). xxvii + 1090 pp. LCCN Z253.4.L38 G66 2004. US$59.99, CAN$86.99.

Murata:1992:FHS

[MH92] Makoto Murata and Koichi Hayashi. Formatter hierarchy for structured documents. In Vanoirbeek and Coray [VC92], pages 77–94. ISBN 0-521-43277-4. LCCN Z 286 E43 I58 1992. UK£35. Prepared with L^{A}TEX, and Microsoft Word converted to L^{A}TEX.

Marcial:2011:SPW

[MH11] Laura Marcial and Brad Hemminger. Scrolling and pagination for within document searching: The impact of screen size and interaction style. Proceedings of the American Society for Information Science and Technology, 48(1):1–4, 2011. ISSN 1550-8390 (print), 1936-1734 (electronic). URL http://

onlinelibrary.wiley.
com/doi/10.1002/meet.
2011.14504801302/abstract.

Miller:1984:PPF

[Mil84] John J. H. Miller, editor.

PROTEXT I: Proceedings of the First International Conference on Text Processing Systems held in Dublin from 24th to 26th October, 1984. Boole Press, Dublin, Ireland, 1984. ISBN 0-906783-41-0 (hardcover), 0-906783-42-9 (paper). LCCN QA76.9.T48 I591 1984.

Miller:1985:PIP

[Mil85] John J. H. Miller, editor. PROTEXT II:

Proceedings of the Second International Conference on Text Processing Systems held in Dublin from 23rd to 25th October, 1985. Boole Press, Dublin, Ireland, 1985. ISBN 0-906783-50-X (hardcover), 0-906783-40-2 (paperback). LCCN QA76.9.T48 I561 1985.

Miller:1986:PIP

[Mil86a] John J. H. Miller, editor. PROTEXT III: Proceedings of the Third International Conference on Text Processing Systems held in Dublin from 22nd to 24th October, 1986. Boole Press, Dublin, Ireland, 1986. ISBN 0-906783-57-7 (hardcover), 0-906783-58-5 (paperback). LCCN QA76.9.T48 I561 1986.

Milne:1986:RTS

[Mil86b] J. S. Milne. Review of two scientific wordprocessing programs: T^{3} version 2.02 by TCI Software Research Inc., and Volkswriter Scientific version 1.0 by Lifetree Software Inc. The Mathematical Intelligencer, 8(1):66–71, 1986.

Miller:1987:PIP

[Mil87] John J. H. Miller, editor. PROTEXT IV: Proceedings of the Fourth International Conference on Text Processing Systems, 20–22 October 1987, Boston, USA. Boole Press, Dublin, Ireland, 1987. ISBN 0-906783-80-1 (hardcover), 0-906783-79-8 (paperback). LCCN QA76.9.T48 I571 1987.

Milne:1990:FWP

[Mil90] J. S. Milne. Four word processors with TEX capabilities. Notices of the American Mathematical Society, 37(8):1018– 1022, October 1990. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Milne:1992:TSW

[Mil92] J. S. Milne. Three scientific word processors. Notices of the American Mathematical Society, 39(8):838–840, October 1992. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic). The word processors ChiWriter 4.02, Leo 1.4, and Scientific Word 1.0 are discussed in terms of ease of use, effectiveness in producing TEX code, and other features of interest to users who require TEX output but do not wish to use TEX directly.

Mittelbach:1990:TGF

[Mit90] Frank Mittelbach. E-TEX: Guidelines for future TEX. TUGboat, 11 (3):337–345, September 1990. ISSN 0896-3207.

Marriott:2007:AFP

[MMH07] Kim Marriott, Peter Moulder, and Nathan Hurst. Automatic float placement in multicolumn documents. In Proceedings of the 2007 ACM symposium on Document Engineering, DocEng ’07, pages 125– 134. ACM Press, New York, NY, USA, 2007. ISBN 1-59593-776-5.

Molnar:1993:TGT

[Mol93] Suzanne M. Molnar. A tour of graphical typesetting on the Macintosh. The UMAP Journal, 14 (2):97–112, 1993. ISSN 0197-3622 (print), 1938-338X (electronic).

Morishita:1989:DLT

[Mor89]  J. Morishita.  DTP

language TEX. Systems, Control and Information, 33(3):117–125, 1989. CODEN SSEJE3. ISSN 0374-4507.

Mortensen:1991:RRT

[Mor91] Niels G. Mortensen. Risø-rapporter i TEX. Matematisk og naturvidenskabelig typografi i praksis (Risø reports in TEX. Mathematical and scientific typography in practice). Risø National Laboratory, Roskilde, Denmark, 1991. ISBN 87-550-1635-9. Prepared using L^{A}TEX, and typeset on a Chelgraph IBX photo-typesetter.

Morin:1993:TT

[Mor93] Richard Morin. Tangled up in TEX. SunExpert Magazine, 4(10):42– 46, October 1993. ISSN 1053-9239. A brief review of TEX support, including the Comprehensive TEX Archive Network (CTAN), the TEX Users Group, and several books on TEX.

Morin:1994:TRPa

[Mor94a] Richard Morin. TEX, Re-DuX (part I). SunExpert Magazine, 5(11):28, 30–31, November 1994. ISSN 1053-9239. Comments on why use TEX. Mentions the TEX Users Group, and the Comprehensive TEX Archive Network (CTAN).

Morin:1994:TRPb

[Mor94b] Richard Morin. TEX, Re-DuX (part II). SunExpert Magazine, 5(12):28– 31, December 1994. ISSN 1053-9239. Lists sources of TEX, and gives short reviews of several books about TEX, AMS-TEX, L^{A}TEX, and literate programming.

Mittelbach:1992:LSD

[MR92a] Frank Mittelbach and Chris Rowley. L^{A}TEX3: structure and design. In Zlatuška [Zla92], pages 69–70. ISBN 80-210-0480-0. Invited talk. Mittelbach:1992:FHQ

[MR92b] Frank Mittelbach and Chris Rowley. The future of high quality typesetting: structure and design. In Zlatuška [Zla92], pages 255–256. ISBN 80-210-0480-0. Invited talk. Mittelbach:1992:PQH

[MR92c] Frank Mittelbach and Chris Rowley. The pursuit of quality: How can automated typesetting achieve the highest standards of craft typography? In Vanoirbeek and Coray [VC92], pages 261–273. ISBN 0-521-43277-4. LCCN Z 286 E43 I58 1992. UK£35. Prepared with L^{A}TEX, and Microsoft Word converted to L^{A}TEX.

Mulvany:1994:IB

[Mul94] Nancy C. Mulvany. Indexing books. Chicago guides to writing, editing, and publishing. University of Chicago Press, Chicago, IL, USA, 1994. ISBN 0-226-55014-1. xiii + 320 pp. LCCN Z695.9 .M8 1994. URL http:/

/www.loc.gov/catdir/
description/uchi052/
93022934.html;   http:
//www.loc.gov/catdir/
toc/uchi051/93022934.
html.

Mulvany:2005:IB

[Mul05] Nancy C. Mulvany. Indexing books. University of Chicago Press, Chicago, IL, USA, second edition, 2005. ISBN 0-226-55276-4. xiv + 315 pp. LCCN Z695.9 .M8 2005. URL http:/

/www.loc.gov/catdir/
enhancements/fy0621/
2005004214-b.html;
http://www.loc.gov/
catdir/enhancements/
fy0621/2005004214-d.
html;   http://www.
loc.gov/catdir/toc/
ecip057/2005004214.
html.

Myhill:1990:PTI

[Myh90]  M. Myhill.  Publish-

ing TEXnical INFOR-MATION: typesetting by database. Program, 24(3):281–289, July 1990. CODEN PRGMBD. ISSN 0033-0337.

Miller:2022:RCF

[MZH22] Barton P. Miller, Mengxiao Zhang, and Elisa R.

Heymann. The relevance of classic fuzz testing: Have we solved this one? IEEE Transactions on Software Engineering, 48(6):2028–2039, June 2022. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL https://arxiv.
org/abs/2008.06537;
 https://ieeexplore.
ieee.org/document/9309406.

Nadrchal:1990:MMI

[Nad90] J. Nadrchal, editor. Man-Machine Interface in the Scientific Environment. Proceedings of the 8th European Summer School on Computing Techniques in Physics. Skalský Dvůr, Czechoslovakia, 19–28 September 1989, volume 61 of Computer Physics Communications. European Summer Schools on Computing Techniques in Physics, North-Holland Publishing Company; Elsevier Science Publishers B. V., Amsterdam, The Netherlands, 1990. In addition to the papers presented at the colloquium, each paper is followed by a summary of the discussion about that paper.

Naiman:1991:UGI

[Nai91] Avi Naiman. The Use of Grayscale for Improved Character Presentation. Ph.d. thesis, University of Toronto, Toronto, Ontario, Canada, 1991. Technical report CSRI-253. To order by mail, send a letter to ATTN: Technical Reports, Computer Systems Research Institute, University of Toronto, 6 King’s College Road, Toronto, Ontario M5S 1A1, CANADA, Tel: (416) 978-8751. The cost is CDN$16 if inside Canada, US$16 if outside. Avi’s e-mail address

is avi@cvs.rochester.
edu.

Nievergelt:1982:DPS

[NCNS82] J. Nievergelt, G. Coray, J.-D. Nicoud, and A. C. Shaw, editors. Document Preparation Systems: a Collection of Survey Articles. Elsevier North-Holland, Inc., New York, NY, USA, 1982. ISBN 0-444-86493-8. xiv + 274 pp. LCCN Z244 .D63 1982. US$46.50.

Negru:1988:CT

[Neg88] John Negru. Computer typesetting. Van Nostrand Reinhold, New York, NY, USA, 1988. ISBN 0-442-26696-0. xiv + 185 pp. LCCN Z253.3 .N4 1988. US$39.95.

Nelson:1978:PD

[Nel78] Roy Paul Nelson. Publication Design. William C. Brown Company Publishers, Dubuque, IA, USA, second edition, 1978. ISBN 0-697-04324-X. 295 pp. LCCN Z246 .N44 1978.

Nelson:1991:PD

[Nel91] Roy Paul Nelson. Publication Design. William C. Brown Company Publishers, Dubuque, IA, USA, fifth edition, 1991. ISBN 0-697-08620-8. xiv + 319 pp. LCCN Z246 .N44 1991.

Noot:1983:STF

[Noo83] Han Noot. Structured text formatting. Software —Practice and Experience, 13(1):79–94, January 1983. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Nixon:1989:NOM

[NP89] Carol Nixon and Lauree Padgett, editors. National Online Meeting proceedings–1989, New York, May 9–11, 1989. Learned Information, Medford, NJ, USA, May 1989. ISBN 0-938734-34-2. LCCN QA76.55 .N37 1989.

NTG:1989:NTG

[NTG ] Nederlandstalige TEX gebruikersgroep, 1989– . Journal of the NTG (Dutch-speaking TEX Users Group).

Oman:1990:TSI

[OC90] Paul W. Oman and Curtis R. Cook. Typographic style is more than cosmetic. Communications of the Association for Computing Machinery, 33(5):506–520, May 1990. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

OConnor:1992:RR

[O’C92]  M. O’Connor.  With

references to references. The EASE Journal: European Science Editing, (47):4–6, September 1992. ISSN 0950-1835.

Ohl:1995:DFD

[Ohl95] Thorsten Ohl. Drawing Feynman diagrams with L^{A}TEX and . Computer Physics Communications, 90(2–3): 340–354, October 1995. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/001046559590137S.

Ohno:1991:DPS

[Ohn91] Y. Ohno. Document preparation software. Journal of the Institute of Electronics, Information and Communication Engineers, 74(2): 175–177, February 1991. CODEN DJTGEB. ISSN 0913-5693.

Olsak:1995:TST

[Olš95] Petr Olšák. Typografický systém TEX. CSTUG, Praha, Czech Republic, 1995. ISBN 80-901950-0-8. 272 pp.

Olsak:1997:TNC

[Olš97] Petr Olšák. TEXbook naruby. (Czech) [TEXbook inside out]. Konvoj, Berkova 22, Brno, 612 00, Czech Republic, 1997. ISBN 80-85615-64-9. 467 pp. 250 Kc. URL http:

//math.feld.cvit.cz/
olsak/tbn.html.

BCS:1986:TWI

[otB86] Report of the Boston Computer Society. Technical Wordprocessors for the IBM PC and Compatibles. Notices of the American Mathematical Society, 33(1):8–37, January 1986. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Padovani:2004:IEM

[Pad04] Luca Padovani. Interactive editing of MathML markup using TEX syntax. In Syropoulos et al. [SBH^{+}04], pages 125–138. CO-DEN LNCSD9. ISBN 3-540-22801-2 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN Z253.3 I58 2004. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t3130.htm;
 http://www.springerlink.
com/openurl.asp?genre=
issue&issn=0302-9743&
volume=3130;   http:
//www.springerlink.
com/openurl.asp?genre=
volume&id=doi:10.1007/
b99374.

Pakin:2005:CLS

[Pak05] Scott Pakin. The comprehensive L^{A}TEX symbol list. Technical report, TeX User Group, Portland, OR, USA, September 22, 2005. URL http:

//www.ctan.org/tex-
archive/info/symbols/
comprehensive/.

Palais:1986:MTP

[Pal86] Richard S. Palais. Mathematical text processing. Notices of the American Mathematical Society, 33(1):3–7, January 1986. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Pappas:1989:TIP

[Pap89] T. L. Pappas. TEXnology on the IBM PC. Computer, 22(8):111–120, August 1989. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Pappas:1990:LPR

[Pap90a]  T. L. Pappas.  Lit-

erate programming for reusability: a queue package example. In Proceedings of Eighth Annual National Conference on Ada Technology, pages 500–514. U.S. Army Commun.- Electron. Command, Fort Monmouth, NJ, USA, 1990.

Pappas:1990:POL

[Pap90b] T. L. (Frank) Pappas. PostScript output on a LaserJet printer. Computer, 23(10):82–88, October 1990. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Pappas:1991:TTD

[Pap91] T. L. (Frank) Pappas. TEX typesetting with a difference. Computer, 24(6):92–93, June 1991. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Partl:1990:HMT

[Par90] Hubert Partl. How to make TEX and L^{A}TEX international. In Nadrchal [Nad90], pages 190–200. CODEN CPHCBZ. Invited paper.

Pawlowski:1993:LCT

[Paw93] Wiesław Pawłowski. L^{A}TEX, czyli TEX dla ludzi. broszura z konferencji członków GUSTu, Bachotek, 1993. In Polish.

Probets:2003:SOF

[PB03] S. G. Probets and D. F.

Brailsford. Substituting outline fonts for bitmap fonts in archived PDF files. Software —Practice and Experience, 33(9):885–899, July 25, 2003. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http:

//brain.mhri.edu.au/
text/texpdf/texpdf-
ffont.html;   http:/
/www.cs.nott.ac.uk/~
dfb/Publications/Download/
2003/Probets03.pdf;
http://www.eprg.org/
research/.

Price:2009:LQS

[PC09] Thomas E. Price and Lance Carnes. L^{A}TEX quick start: a first guide to document preparation. Personal TEX, Inc., San Francisco, CA, USA, 2009. ISBN 0-9631044-4-6. iv + 135 pp. LCCN ????

Poppelier:1992:BR

[PH92] Nico Poppelier and Amy Hendrickson. Book reviews. Nederlandstalige TEX Gebruikersgroep (Dutch-speaking TEX Users Group), 92 (1):135–138, May 1992.

Plaice:2004:MOO

[PHSB04] John Plaice, Yannis Haralambous, Paul Swoboda, and Gábor Bella. Moving Ω to an objectoriented platform. In Syropoulos et al. [SBH^{+}04], pages 17–26. CO-DEN LNCSD9. ISBN 3-540-22801-2 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN Z253.3 I58 2004. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t3130.htm;
 http://www.springerlink.
com/openurl.asp?genre=
issue&issn=0302-9743&
volume=3130;   http:
//www.springerlink.
com/openurl.asp?genre=
volume&id=doi:10.1007/
b99374.

Pietrowski:1991:ISI

[Pie91] Ken Pietrowski, editor. IS&T’s Seventh International Congress on Advances in Non-impact Printing Technologies, October 6–11, Portland, OR, USA., volume 2.

Imaging Science & Technology, ????, 1991.

Piska:2004:CTF

[Píš04] Karel Píška. Creating Type 1 fonts from - sources: Comparison of tools, techniques and results. In Syropoulos et al. [SBH^{+}04], pages 240–256. CO-DEN LNCSD9. ISBN 3-540-22801-2 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN Z253.3 I58 2004. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t3130.htm;
http://www.springerlink.
com/openurl.asp?genre=
issue&issn=0302-9743&
volume=3130;   http:
//www.springerlink.
com/openurl.asp?genre=
volume&id=doi:10.1007/
b99374.

Plass:1982:CBL

[PK82] Michael F. Plass and Donald E. Knuth. Choosing better line breaks. In Nievergelt et al. [NCNS82], chapter ??, pages 221–242. ISBN 0-444-86493-8. LCCN Z244 .D63 1982. US$46.50.

Plass:1981:OPT

[Pla81] Michael F. Plass. Optimal pagination techniques for automatic typesetting systems. Ph.D. dissertation, Computer Science Department, Stanford University, Stanford, CA, USA, 1981. vi + 72 pp. URL http://

wwwlib.umi.com/dissertations/
fullcit/8124134.

Perez:2004:BCS

[PLA04] Jagoba Arias Pérez, Jesús Lázaro, and Juan M. Aguirregabiria. Basque: a case study in generalizing language support. In Syropoulos et al. [SBH^{+}04], pages 27–33. CODEN LNCSD9. ISBN 3-540-22801-2 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN Z253.3 I58 2004. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t3130.htm;
 http://www.springerlink.
com/openurl.asp?genre=
issue&issn=0302-9743&
volume=3130;   http:
//www.springerlink.
com/openurl.asp?genre=
volume&id=doi:10.1007/
b99374.

Platoni:2006:LFB

[Pla06] Kara Platoni. Love at first byte. Stanford Magazine, 35(3):64–69, May/ June 2006. ISSN 1063-

2778. URL http://www.
stanfordalumni.org/
news/magazine/2006/
mayjun/features/knuth.

html. Photography by Timothy Archibald.

Polderman:1993:YWN

[Pol93] A. K. S. Polderman. So you want a numerical reference list? The EASE Journal: European Science Editing, (50):18–19, September 1993. ISSN 0950-1835.

Piccoli:2012:OPC

[POM12] Ricardo Piccoli, João Oliveira, and Isabel Manssour. Optimal pagination and content mapping for customized magazines. Journal of the Brazilian Computer Society, 18(4):331–349, November 2012. ISSN 0104-6500 (print), 1678-4804 (electronic). URL

http://link.springer.
com/article/10.1007/
s13173-012-0066-6.

Poppelier:1992:BRJ

[Pop92] Nico Poppelier. Book reviews: Jane Hahn, L^{A}TEX for Everyone; Eric van Herwijnen, Practical SGML. TUGboat, 13(2): 182–184, July 1992. ISSN 0896-3207.

Porrat:1992:DHT

[Por92] Rama Porrat. Developments in Hebrew TEX. In Zlatuška [Zla92], pages 135–148. ISBN 80-210-0480-0.

Porter:1997:MTN

[Por97] Richard D. Porter. Math into Type: The Next Generation. American Mathematical Society, Providence, RI, USA, 1997. ISBN 0-8218-0644-0. 120 pp. US$19.00.

Wood:19xx:PLJ

[POSxx] PostScriptLanguage Journal. Pipeline Associates, Inc., P. O. Box 5763, Parsippany, NJ 07054, 19xx.

Porter:1992:DRA

[PR92] Gilbert B. Porter III and Emil V. Rainero. Document reconstruction: a system for recovering document structure from layout. In Vanoirbeek and Coray [VC92], pages 127–141. ISBN 0-521-43277-4. LCCN Z 286 E43 I58 1992. UK£35. Prepared with L^{A}TEX, and Microsoft Word converted to L^{A}TEX.

Przechlewski:2004:MTR

[Prz04] Tomasz Przechlewski. Managing TEX resources with XML topic maps. In Syropoulos et al. [SBH^{+}04], pages 216– 228. CODEN LNCSD9. ISBN 3-540-22801-2 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN Z253.3 I58 2004. URL

http://link.springer-
ny.com/link/service/
series/0558/tocs/t3130.
htm;   http://www.springerlink.
com/openurl.asp?genre=
issue&issn=0302-9743&
volume=3130;   http:
//www.springerlink.
com/openurl.asp?genre=
volume&id=doi:10.1007/
b99374.

Partl:1993:DL

[PSH93] Hubert Partl, Elis. Schlegl, and Irene Hyna. Una descripción de L^{A}TEX, 1993. Spanish translation by Tomas Bautista of German introduction to L^{A}TEX, L^{A}TEX-Kurzbeschreibung. Includes material on the spanish style option. Available in Spain from

ftp.eunet.es (also known
as goya.uu.es) in the
directory info/UNIX/
text/TeX/spanish/tutorial,

as well as from the main TEX archives.

Rahtz:1987:ITH

[Rah87a] Sebastian Rahtz, editor.

Information technology in the humanities: tools, techniques and applications. Ellis Horwood series in computers and their applications. Halsted Press, New York, USA, 1987. ISBN 0-470-20852-X, 0-7458-0148-X. 184 pp. LCCN AZ105 .I54 1987. UK£25.00, US$39.95.

Rahtz:1987:PW

[Rah87b] Sebastian Rahtz. The processing of words. In Information technology in the humanities: tools, techniques and applications [Rah87a], pages 69–79. ISBN 0-470-20852-X, 0-7458-0148-X. LCCN AZ105 .I54 1987. UK£25.00, US$39.95.

Rahtz:1989:STG

[Rah89] Sebastian Rahtz. A survey of TEX and graphics. Technical Report CSTR 89-7, University of Southampton, Department of Electronics and Computer Science, Southampton SO9 5NH, UK, October 1989.

Reid:1980:SDS

[Rei80] Brian Keith Reid. Scribe: A Document Specification Language and its Compiler. Ph.D. thesis, Department of Computer Science, Carnegie-Mellon University, Pittsburgh, PA, USA, October 1980. x + 148

                                        pp.  URL http://
                                        search.proquest.com/
                                        docview/303093737;
                                        http://reports-archive.
                                        adm.cs.cmu.edu/anon/
                                        scan/CMU-CS-81-100.

pdf. This influential work, completely independent of TEX, produced a commercial product (no longer marketed) that was one of the first document formatters to emphasize logical markup over physical markup. Its markup system strongly influenced the design of both L^{A}TEX and BibTEX. Some formatting variations of a Scribe bibliography file can be compatible with BibTEX.

Reid:1988:PLP

[Rei88] Glenn C. Reid. PostScript Language Program Design. Addison-Wes-ley, Reading, MA, USA, 1988. ISBN 0-201-14396-8. xii + 224 pp. LCCN QA76.73.P67 R45 1988. US$22.95. URL http:/

/www.rightbrain.com/
rightbrain.shtml.

Revesz:1991:TOM

[Rev91] G. E. Revesz. On translating ordinary mathematical notation. Structured programming, 12 (3):115–122, 1991. CO-DEN STPGEM. ISSN 0935-1183.

Rice:1978:BDT

[Ric78]  Stanley Rice.  Book

Design—Text Format Models. R. R. Bowker Company, 1180 Avenue of the Americas, New York, NY 10036, USA, 1978. ISBN 0-8352-1045-6. xiv + 215 pp. LCCN Z245 R5 1978. US$17.50.

Rosson:1988:DUB

[RKM88] Mary B. Rosson, Wendy Kellogg, and Susanne Maass. The designer as user: building requirements for design tools from design practice. Communications of the Association for Computing Machinery, 31 (11):1288–1298, November 1988. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). From Computing Reviews: “Inter-Views objects are combined through a model similar to the boxes and glues in Knuth’s TEX, which allows interface objects to be combined with little concern for undesired interactions.”.

Rolland:1993:LGP

[Rol93] Christian Rolland. L^{A}TEX guide pratique. Addison-Wesley, Reading, MA, USA, July 1993. ISBN 2-87908-025-8. 280 pp. FF 190. Distributed by Bourdas, code 009808.

Rolland:1995:LGP

[Rol95] Christian Rolland. L^{A}TEX: guide pratique. Addison-Wesley France, Paris, France, deuxième edition, 1995. ISBN 2-87908-104-1. xix + 345 pp. LCCN ????

Rolland:1999:LPP

[Rol99] Christian Rolland. L^{A}TEX par la pratique. O’Reilly &Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, 1999. ISBN 2-84177-073-

7. xx + 559 pp. LCCN ????

Rosset:1984:SSA

[Ros84] J.-L. Rosset, editor.

S.A.PR.I.S.T.I. (système assisté de PRoduction intégrant simplement textes et images). Supplément au No. 65, CIRCE, Orsay, France, 1984.

Rose:1992:HTP

[Ros92a] Kristoffer H. Rose. How to typeset pretty diagram arrows with TEX

— design decisions used in Xy-pic. In Zlatuška [Zla92], pages 183–210. ISBN 80-210-0480-0.

Rose:1992:TDX

[Ros92b] Kristoffer H. Rose. Typesetting diagrams with XY-pc: user’s manual. In Zlatuška [Zla92], pages 273–292. ISBN 80-210-0480-0.

Roth:1988:RWP

[Rot88] Stephen E. Roth, editor. Real World Post-Script. Addison-Wes-ley, Reading, MA, USA, 1988. ISBN 0-201-06663-7. xiv + 383 pp. LCCN Z286.D47 R4 1988.

Rowley:2001:LLA

[Row01] Chris Rowley. The L^{A}TEX legacy: 2.09 and all that. In ACM, editor, Proceedings of the Twentieth Annual ACM Symposium on Principles of Distributed Computing 2001, Newport, Rhode Island, United States, pages 17–25. ACM Press, New York, NY, USA, 2001. ISBN 1-58113-383-9. LCCN ????

Roy:1984:TWT

[Roy84] Yves Roy. TEX/WEB, et le traitement de textes mathématiques. Études et Recherches en Informatique. [Studies and Research in Computer Science]; Monographs in Visual Communication. Masson Editeur, Paris, France, 1984. ISBN 2-225-80249-1. vii + 102 pp. LCCN Z253.4.T47 R69 1984. Preface de D. Foata.

Romberger:1985:ATL

[RS85] Staffan Romberger and Yngve Sundblad. Adapting TEX to Languages that use Latin Alphabetic Characters. In Lucarella [Luc85b], pages 27–40. ISBN 0-201-13399-7. LCCN Q224.E97 1985.

Risueno:2016:IPC

[RSV16] Pablo García Risueño, Apostolos Syropoulos, and Natàlia Vergés. Ideograms for physics and chemistry. Foundations of Physics, 46 (12):1713–1721, December 2016. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:

//link.springer.com/
article/10.1007/s10701-
016-0044-5.

Rubinstein:1988:DTI

[Rub88] Richard Rubinstein. Digital Typography—An Introduction to Type and Composition for Computer System Design. Addison-Wesley, Reading, MA, USA, November 1988. ISBN 0-201-17633-5. xi + 340 pp. LCCN Z253.3 .R8 1988. US$32.25. Reprinted with corrections.

Ruder:1988:TG

[Rud88] Emil Ruder. Typographie; Ein Gestaltungsbuch. Niggli/Hatje, Heiden, Stuttgart, Germany, fifth revised edition, 1988. Contains complete English and French translation.

Ruggles:1983:LDS

[Rug83]  Lynn Ruggles.  Let-

terform design systems. Technical Report STAN-CS-83-971, Stanford University, Stanford, CA, USA, April 1983.

Ruggles:1987:PIE

[Rug87] Lynn Elizabeth Ruggles. Paragon, an interactive, extensible, environment for typeface design. Ph.D. dissertation, University of Massachusetts Amherst, Amherst, MA, USA, 1987. 192 pp.

                                        URL http://wwwlib.
                                        umi.com/dissertations/
                                        fullcit/8805968.

Rybicka:1995:PZ

[Ryb95]  Jiři Rybička.  L^{A}TEX

pro začátečníky. Konvoy, Brno, Czech Republic, 1995. ISBN 80-85615-42-8. 169 pp.

Rycko:1993:CNA

[Ryć93] Marek Ryćko. Co nieco o akapicie. broszura z konferencji członków GUSTu, Bachotek, 1993. In Polish.

Salomon:1995:AT

[Sal95] David Salomon. The Advanced TEXbook. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1995. ISBN 0-387-94556-3. xx + 490 pp. LCCN Z253.4.T47 S25 1995. US$39.95.

URL http://www.ecs.
csun.edu/~dxs/tatb.
advertis/tatbAd.html.

Samuel:1983:FGT

[Sam83] Arthur Samuel. First grade TEX. A beginner’s TEX manual. Technical Report STAN-CS-83-985, Stanford University, Stanford, CA, USA, November 1983. Also published by TUG as TEXniques No. 11.

Sankar:1989:ASG

[San89] S. Sankar. APE—a set of TEX macros to format Ada programs. ACM SIGADA Ada Letters, 9 (7):114–128, November/ December 1989. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Sutton:1990:TB

[SB90] James Sutton and Alan Bartram. Typefaces for Books. The British Library and New Amsterdam Books, London, UK and New York, NY, USA, 1990. ISBN 1-56131-016-6, 1-56131-017-4 (paperback). 288 pp. LCCN Z250 .S896 1990. Contains full-page type samples in a variety of point sizes for a large number of typefaces.

Syropoulos:2004:TXD

[SBH^{+}04] Apostolos Syropoulos, Karl Berry, Yannis Haralambous, Baden Hughes, Steven Peter, and John Plaice, editors. TEX, XML, and Digital Typography: International Conference on TEX, XML, and Digital Typography, held jointly with the 25th Annual Meeting of the TEX Users Group, TUG 2004, Xanthi, Greece, August 30–September 3, 2004: Proceedings, volume 3130 of Lecture Notes in Computer Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2004. CODEN LNCSD9. ISBN 3-540-22801-2 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN Z253.3 I58 2004. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t3130.htm;
 http://www.springerlink.
com/openurl.asp?genre=
issue&issn=0302-9743&
volume=3130;   http:
//www.springerlink.
com/openurl.asp?genre=
volume&id=doi:10.1007/
b99374.

Schwartz:1988:WYA

[Sch88a] Laurens R. Schwartz.

What You Aren’t Supposed To Know About Writing And Publishing: An Expos/’e of Editors, Agents, Publishing Houses, and More —An Insider’s Report. Shapolsky Publishers, New York, NY, USA, 1988. ISBN 0-944007-02-3. viii + 173 pp. LCCN KF300.Z9 S35 1988.

Schwarz:1988:ET

[Sch88b] Norbert Schwarz. Einführung in TEX. Addison-Wes-ley Verlag, Bonn, Germany, second edition, 1988. ISBN 3-925118-97-7. 272 pp. Available in English translation, [KS90], and in Dutch translation, [Sch90b].

Schumann:1989:PBT

[Sch89] Lothar Schumann. Professioneller Buchsatz mit TEX. R. Oldenbourg Verlag, Munich, Germany and Vienna, Austria, 1989. ISBN 3-486-21173-0. viii + 336 pp.

Schwarz:1990:ET

[Sch90a] Norbert Schwarz. Einführung in TEX. Addison-Wesley Verlag, Bonn, Germany, second extended edition, 1990. ISBN 3-925118-97-7. 272 pp. Available in English translation, [KS90], and in Dutch translation, [Sch90b].

Schwarz:1990:IT

[Sch90b] Norbert Schwarz. Inleiding TEX. Addison-Wes-ley Europe, Amsterdam, The Netherlands, 1990. ISBN 90-6789-151-7. 282 pp. Dutch translation of [Sch88b]. See also the English translation, [KS90].

Schrod:1991:IVM

[Sch91a] Joachim Schrod. An international version of makeindex. Cahiers GUTenberg, 10–11:81– 90, September 1991. ISSN 1140-9304. Schumann:1991:PBT

[Sch91b] Lothar Schumann. Professioneller Buchsatz mit TEX: Lehrbuch für Anwender. R. Oldenbourg Verlag, Munich, Germany and Vienna, Austria, 1991. ISBN 3-486-21736-4. viii + 336 pp.

Schwarz:1991:ET

[Sch91c] Norbert Schwarz. Einführung in TEX. Addison-Wes-ley Verlag, Bonn, Germany, third edition, 1991. ISBN 3-925118-97-7. 336 pp. Available in English translation, [KS90], and in Dutch translation, [Sch90b].

Schrell:1993:DTS

[Sch93] DANTE-TEX-stammtisch in wuppertal: Jahrbuch 1992/93, 1993.

Scherber:1994:OST

[Sch94]  Peter Scherber.  Of-

fizin—Schriftenreihe zu TEX, L^{A}TEX, und Metafont (Band 1, Hrsg. von DANTE e.V). Addison-Wesley, Reading, MA, USA, 1994. ISBN 3-89319-396-0. x + 159 pp. DM 49,00.

Schaefer:2004:SSC

[Sch04] Frank-Rene Schaefer.

ŞäferTEX: Source code esthetics for automated typesetting. In Syropoulos et al. [SBH^{+}04], pages 229–239. CO-DEN LNCSD9. ISBN 3-540-22801-2 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN Z253.3 I58 2004. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t3130.htm;
 http://www.springerlink.
com/openurl.asp?genre=
issue&issn=0302-9743&
volume=3130;   http:
//www.springerlink.
com/openurl.asp?genre=
volume&id=doi:10.1007/
b99374.

Scheinerman:2011:MN

[Sch11] Edward R. Scheinerman.

Mathematical notation. CreateSpace, USA, 2011. ISBN 1-4662-3052-5. viii + 84 pp. LCCN QA41 .S34 2011.

Sojka:1992:TDP

[SČS92] Petr Sojka, Rudolf Červenka, and Martin Svoboda. TEX for database publishing. In Zlatuška [Zla92], pages 53–58. ISBN 80-210-0480-0.

Seroul:1989:PLT

[Ser89] Raymond Seroul. Le petit Livre de TEX. InterEditions, Paris, France, 1989. ISBN 2-7296-0233-X. 317 pp.

Sewell:1989:WPL

[Sew89] E. Wayne Sewell. Weaving a Program: Literate Programming in WEB. Van Nostrand Reinhold, New York, NY, USA, 1989. ISBN 0-442-31946-0. xx + 556 pp. LCCN QA76.73.W24 S491 1989. US$34.95.

Seybold:1982:DPS

[Sey82] J. W. Seybold. Document preparation systems and commercial typesetting. In Nievergelt et al. [NCNS82], pages 243–264. ISBN 0-444-86493-8. LCCN Z244 .D63 1982. US$46.50.

Skillin:1974:WT

[SG74] Marjorie E. Skillin and Robert Malcolm Gay. Words into Type. Prentice-Hall, Upper Saddle River, NJ 07458, USA, third edition, 1974. ISBN 0-13-964262-5. xx + 583 pp. LCCN PN160.S52 1974. US$39.95.

Shaw:2001:FJT

[Sha01] J. G. Shaw. Full journal titles to be given for citations. The EASE Journal: European Science Editing, 27(1):3, ???? 2001. ISSN 0950-1835.

Shirokov:1994:PPS

[Shi94] Boris Shirokov. PCTEX.

Prostoi sposob iziaschnogo oformleniia mathematicheskikh tekstov: uchebnoe posobie [English: PCTEX. Simple method of elegant markup of mathematical texts: textbook]. Petrozavodsk State University, Petrozavodsk, Russia, 1994. ISBN 5-230-08971-7. 80 pp. This small manual talks about working with TEX on IBM PC compatible computers (not about PCTEX!).

Scavo:1993:TM

[SHS93] Tom Scavo, Yannis Haralambous, and Werenfried Spit. TEX on the macintosh. Notices of the American Mathematical Society, 40 (10):1353–1360, December 1993. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Shultis:1994:LNP

[Shu94] J. Kenneth Shultis. L^{A}TEX Notes: Practical Tips for Preparing Technical Documents. Prentice-Hall PTR, Upper Saddle River, NJ 07458, USA, 1994. ISBN 0-13-120973-6. xi + 180 pp. LCCN Z253.4.L38 S48 1994. US$32.00.

Siegel:1985:EPA

[Sie85] David R. Siegel. The Euler Project at Stanford. Technical report, Stanford University, Stanford, CA, USA, 1985.

Siebenmann:1986:TWU

[Sie86] L. C. Siebenmann. Toward Wider Use of TEXTypesetting. A Pre– TEXManifesto. Notices of the American Mathematical Society, 33(4):597–607, August 1986. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Siemoneit:1989:TG

[Sie89] Manfred Siemoneit. Typographisches Gestalten. Polygraph-Verlag, Frankfurt am Main, Germany, second edition, 1989.

Siebenmann:1992:CGR

[Sie92a] Laurent Siebenmann. CaesarCM — A gentle road to perfect hyphenation in modest TEX environments. In Zlatuška [Zla92], pages 97–106. ISBN 80-210-0480-0.

Siebenmann:1992:MAD

[Sie92b] Laurent Siebenmann.

La mise en application définitive de la norme de Cork. In Zlatuška [Zla92], pages 107–118. ISBN 80-210-0480-0.

Siebenmann:1992:LM

[Sie92c] Laurent Siebenmann.

The lion and the mouse. In Zlatuška [Zla92], pages 43–52. ISBN 80-210-0480-0.

Simon:1987:SCO

[Sim87] John J. Simon, Jr. From sand to circuits and other inquiries. Harvard Univ. Of. for Info. and Tech., Cambridge, MA, USA, 1987. ISBN 0-674-32575-3. xi + 271 pp. LCCN QA76.24 .F76 1986, T185.F76 1986, TK7828 .S48 1986. US$24.95. Includes chapter on computer typesetting.

Sawyer:1995:TPS

[SK95] Stanley A. Sawyer and Steven Krantz. A TEX Primer for Scientists. CRC Press, 2000 N.W. Corporate Blvd., Boca Raton, FL 33431-9868, USA, 1995. ISBN 0-8493-7159-7. viii + 399 pp. US$29.95.

Shin:1997:TMF

[SKS97] K.-H. Shin, K. Kobayashi, and A. Suzuki. Tafel Musik: formatting algorithm of tables. Mathematical and Computer Modelling, 26(1):97–112, 1997. CODEN MC-MOEG. ISSN 0895-7177.

Seroul:1991:BBT

[SL91] Raymond Seroul and Silvio Levy. A Beginner’s Book of TEX. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1991. ISBN 0-387-97562-4, 3-540-97562-4. xii + 282 pp. LCCN Z253.4.T47 S47 1991. This is a translation and adaption by Silvio Levy of [Ser89].

Smith:1989:CLF

[Smi89] N. E. Smith. Coping with a large Forth system. In Forsley [For89], pages 113–115. ISBN 0-914593-09-9. LCCN TJ212.2.A15 1989.

Smith:1990:PCS

[Smi90]  N. E. Smith.  Pars-

ing character strings in Forth. In 1990 Rochester Forth Conference: Embedded Systems, pages 133–134. Institute for Applied FORTH Research, Rochester, NY, USA, 1990. ISBN 0-914593-10-2. LCCN QA76.73.F24 R59 1990.

Snow:1992:TB

[Sno92] Wynter Snow. TEX for the Beginner. Addison-Wesley, Reading, MA, USA, 1992. ISBN 0-201-54799-6. xii + 377 pp. LCCN Z253.4.T47 S64 1992.

Soiffer:1995:MTM

[Soi95]  N. Soiffer.  Mathe-

matical typesetting in Mathematica. In Levelt [Lev95], pages 140– 149. ISBN 0-89791-699-9. LCCN QA 76.95 I59 1995. ACM order number: 505950.

Southall:1985:MPD

[Sou85a]  Richard Southall.  -

and the problem of type design. In André [And85], pages 133–145.

Southall:1985:DNT

[Sou85b] Richard Southall. Designing new typefaces with Metafont. Technical Report STAN-CS-85-1074, Stanford University, Stanford, CA, USA, September 1985.

Southall:1992:PRR

[Sou92] Richard Southall. Presentation rules and rules of composition in the formatting of complex text. In Vanoirbeek and Coray [VC92], pages 275–290. ISBN 0-521-43277-4. LCCN Z 286 E43 I58 1992. UK£35. Prepared with L^{A}TEX, and Microsoft Word converted to L^{A}TEX.

Sowa:1994:TLG

[Sow94] Friedhelm Sowa. TEX/ L^{A}TEX und Graphik: Eine Überblick über die Verfahren [English: TEX/L^{A}TEX and Graphics: A Survey of the Methods]. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1994. ISBN 3-540-56468-3. xv + 115 pp. DM39.00.

Schwartz:1992:LSS

[SP92] Stefan Schwartz and Rudolf Potuček. L^{A}TEX —Satzkunst statt DTP (L^{A}TEX—The Art of Typography rather than DTP). Vogel-Verlag, Postfach 67 40, D-8700 Würzburg, Germany, 1992. ISBN 3-8023-1178-7. 106 pp. DM34.00.

Schwarz:1997:TRF

[SP97] Stefan Schwarz and Rudolf Potuček. Das TEXikon: Referenzhandbuch für TEX und L^{A}TEX. Addison-Wesley Longman, Reading, MA, USA, 1997. ISBN 3-89319-690-0. 774 pp. DM 69,90.

Spivak:1980:JTG

[Spi80] M. Spivak. The Joy of TEX: a Gourmet Guide to Typesetting Technical Text by Computer: With Numerous Explicit Illustrations (Version 1). American Mathematical Society, Providence, RI, USA, 1980. 134 pp. US$9.00.

Spivak:1985:PM

[Spi85] Michael D. Spivak. PCTEX Manual. 12 Madrona Street, Mill Valley, CA 94941, USA, 1985.

Spivak:1986:JTG

[Spi86a] Michael D. Spivak. The Joy of TEX—A Gourmet Guide to Typesetting with the AMS-TEX macro package. American Mathematical Society, Providence, RI, USA, 1986. ISBN 0-8218-2999-8. xviii + 290 pp. LCCN Z253.4.T47 S6731 1986.

Spivak:1986:MPT

[Spi86b] Michael D. Spivak. Manual “PCTEX” ou TEX en 9 leçons. Orsay, France, 1986. French translation of [Spi85].

Spivak:1989:LS

[Spi89] Michael D. Spivak. LAMS-TEX, The Synthesis. The TEXplorators Corporation, 3701 W. Alabama, Suite 450-273, Houston, TX 77027, USA, 1989. viii + 289 pp.

Spivak:1990:JTG

[Spi90] Michael D. Spivak. The Joy of TEX: a Gourmet Guide to Typesetting with the AMS-TEX macro package. American Mathematical Society, Providence, RI, USA, second revised edition, 1990. ISBN 0-8218-2997-1. xxii + 309 pp. LCCN Z253.4.T47 S6731 1990. US$40.00. URL https:

//ctan.org/pkg/joy-
of-tex.

Spivak:1991:LWM

[Spi91] Michael D. Spivak. The L^{A}MS-TEX Wizard’s Manual. The TEXplorators Corporation, 3701 W. Alabama, Suite 450-273, Houston, TX 77027, USA, 1991.

Spivak:1993:VT

[Spi93] Michael D. Spivak. Voskhititel’nii TEX [English: Joy of TEX]. Mir Publishers, Moscow, Russia, 1993. ISBN 5-03-001942-1. 285 pp. Russian translation of The Joy of TEX. Translated by Irina Makhovaya and edited by Andrei Khodulev. First book about TEX in Russian. First press run: 7000 copies.

Srinivasan:1993:UDP

[Sri93] Balasubramaniam Srinivasan. UNIX Document Processing and Typesetting. World Scientific, Singapore, 1993. ISBN 981-02-0605-4. xiv + 445 pp. LCCN Z52.5.U54S65 1993.

Sherborne:1990:RGC

[SS90] Brad S. Sherborne and G. E. Stedman. Recursive generation of Cartesian angular momentum coupling trees for SO(3). Computer Physics Communications, 59(2):417– 428, June 1990. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/0010465590901887.

Smith:1991:ABL

[SS91] Lisa M. C. Smith and Mansur H. Samadzadeh. An annotated bibliography of literate programming. ACM SIGPLAN Notices, 26:14–20, January 1991. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Tang:1981:PIC

[sT81] Chih sung Tang. On the problem of inputting Chinese characters. Report STAN-CS-81-848, Stanford University, Department of Computer Science, Stanford, CA, USA, 1981.

Schlager:2005:WLA

[ST05] Petra Schlager and Manfred Thibud. Wissenschaftlich mit L^{A}TEX arbeiten. (German) [Working scientifically with L^{A}TEX]. Pearson Studium, Munich, Germany, 2005. ISBN 3-8273-7078-7. 218 pp. LCCN ???? Includes CD-ROM.

Steen:1981:TTE

[Ste81] Lynn Arthur Steen. Technical text editing. The UMAP Journal, 2 (3):17–35, 1981. ISSN 0197-3622 (print), 1938-338X (electronic).

Stone:1992:CAR

[Sto92] Harold S. Stone. Copyrights and author responsibilities. Computer, 25(12):46–51, December 1992. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). This paper discusses U.S. copyright law and its implications for the reuse of published text. Guidelines are given for the citation and quoting of such material.

Syropoulos:2003:DTU

[STS03] Apostolos Syropoulos, Antonis Tsolomitis, and Nick Sofroniou. Digital typography using L^{A}TEX. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2003. ISBN 0-387-95217-9 (paperback), 0-387-22436-X (e-book). xxix + 510 pp. LCCN Z253.4.L38 S97 2003.

Steemers:1989:JSG

[SvdL89] L. Steemers and C. G.

van der Laan. Journal style guidelines. RC-Rapport 26, Rekencentrum der Rijksuniversiteit te Groningen, Groningen, The Netherlands, 1989.

Stoffels:1992:APS

[SvR92] J. Stoffels and C. van Reeuwijk. Ampdes: a program for the synthesis of high-performance amplifiers. In IEEE ECDA ’92 [IEE92], pages 474– 479. ISBN 0-8186-2645-3. LCCN TK7867 .E93 1992. Semenzato:1992:AAL

[SW92] Luigi Semenzato and Edward Wang. Aleph: a language for typesetting. In Vanoirbeek and Coray [VC92], pages 65– 75. ISBN 0-521-43277-4. LCCN Z 286 E43 I58 1992. UK£35. Prepared with L^{A}TEX, and Microsoft Word converted to L^{A}TEX.

Swanson:1986:MTC

[Swa86] Ellen Swanson. Mathematics into Type; Copy Editing and Proofreading of Mathematics for Editorial Assistants and Authors. American Mathematical Society, Providence, RI, USA, revised edition, 1986. ISBN 0-8218-0053-1. x + 90 pp. LCCN T11.4.S96 1979.

Talbot:2012:LCN

[Tal12] Nicola L. C. Talbot. L^{A}TEX for Complete Novices, volume 1 of Dickimaw LaTeX Series. Dickimaw Books, Saxlingham Nethergate, UK, 2012. ISBN 1-909440-00-0 (paperback). ix + 277 pp. LCCN Z253.4.L38 T35 2012.

Talbot:2015:LAW

[Tal15] Nicola L. C. Talbot. L^{A}TEX for Administrative Work, volume 3 of Dickimaw LaTeX Series. Dickimaw Books, Norfolk, UK, 2015. ISBN 1-909440-07-8 (paperback). xiv + 656 pp. LCCN ????

Tatarkiewicz:1988:TLD

[Tat88] Jakub Tatarkiewicz. TEXnika ładnego druku. Computer, 8:36–37, 1988. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). In Polish.

Taupin:1992:MUT

[Tau92]  Daniel Taupin.  Mu-

sicTEX: using TEX to write polyphonic or instrumental music. In Zlatuška [Zla92], pages 257–272. ISBN 80-210-0480-0. Invited talk.

Taylor:1992:FT

[Tay92] Phillip Taylor. The future of TEX. In Zlatuška [Zla92], pages 235–254. ISBN 80-210-0480-0.

TEXEURO:1989:TE

[TEX a] TEX-euro, 1989–. This is an unmoderated discussion list for TEX with emphasis on European issues. To subscribe, send a request with the text subscribe

tex-euro to listserv@
dhdurz1.bitnet.

TEXHAX:1987:T

[TEX b] TEXhax, 1987–. This is a TUG-supported moderated electronic mailing list. To subscribe, send a request to

texhax-request@june.
cs.washington.edu.

Clark:1987:T

[TEX c] Malcolm Clark, editor. TEXline, 1987–. ISSN 0961-3978. Warwick, UK. This is an informal newsletter of the TEX community.

TEXMAG:1988:T

[TEX d] TEXmag, 1988–. This is an electronic magazine with articles about TEX. To subscribe, send a request with the text

                                        subscribe texmag-l to
                                        listserv@vm.byu.edu.

ITUG:2003:LRM

[The03] The Indian TEX Users Group. The L^{A}TEX Reference Manual. GNU Press, Boston, MA, USA, 2003. ISBN 1-882114-70-1. 250 (est.) pp. LCCN ???? US$40.00 (est.). URL http:

                                        //www.gnupress.org/
                                        book15.html.

Thimbleby:1991:LTT

[Thi91] H. Thimbleby. Low-tech TEX? (typesetting program). In Proceedings of the Fourth Annual Conference on Computers and the Writing Process, pages 124–130. Univ. Sussex, Brighton, UK, 1991.

Teo:1988:UVS

[TM88] G. S. Teo and M. Mac an Airchinnigh. The use of VDM in the specification of Chinese characters. In Bloomfield et al. [BMJ88], pages 476– 499. ISBN 0-387-50214-9 (New York), 3-540-50214-9 (Berlin). LCCN QA76.76.D47 V8541 1988.

Tobin:1985:EMS

[Tob85] Georgia K. M. Tobin.

The elements of Metafont style, 1985.

Trickey:1988:UL

[Tri88] Howard Trickey. Using L^{A}TEX. ACM SIGACT News, 19(2):55–57, Summer 1988. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Trzeciak:1995:WMP

[Trz95] Jerzy Trzeciak. Writing Mathematical Papers in English: a Practical Guide. European Mathematical Society Publishing House, Seminar for Applied Mathematics, Zürich, Switzerland, revised edition, 1995. ISBN 3-03719-014-0, 83-85694-02-1. 49 pp. LCCN QA42 .T79 1995.

URL http://www.ems-
ph.org/books/book.php?
proj_nr=34.

Tschichold:1928:NTH

[Tsc28]  Jan Tschichold.  Die

neue Typographie: ein Handbuch für zeitgemäss Schaffende. Verl. das Bildungsverbandes der Deutschen Buchdrucker, Berlin, Germany, 1928. 240 pp.

Tschichold:1986:TMS

[Tsc86]  Jan Tschichold.  Ty-

pographische Mitteilungen, Sonderheft Elementare Typographie. Universitätsdruckerei und Verlag Hermann Schmidt, Mainz, Germany, 1986. ISBN 3-87439-233-3. 36 pp. DM 24.

Tschichold:1987:AAF

[Tsc87] Jan Tschichold. Ausgewählte Aufsätze über Fragen der Gestalt des Buches. Birkhäuser, Cambridge, MA, USA; Berlin, Germany; Basel, Switzerland, 1987. ISBN 3-7643-1946-1. Second printing.

Tschichold:1991:NT

[Tsc91a]  Jan Tschichold.  Die

Neue Typographie. Universitätsdruckerei und Verlag Hermann Schmidt, Mainz, Germany, 1991. 240 pp. DM60.

Tschichold:1991:FB

[Tsc91b] Jan Tschichold. The Form of the Book. Lund Humphries, London, UK, 1991. ISBN 0-85331-623-6. xviii + 180 pp. LCCN Z116.A3 T7613 1991b. UK£18.95. Originally published in German as Ausgewählte Ausfsätze über Fragen der Gestalt des Buches und der Typographie by Birkhäuser Verlag, Basel, 1975.

Tschichold:1995:NTHa

[Tsc95a]  Jan Tschichold.  Die

neue Typographie: ein Handbuch für zeitgemäss Schaffende. Brinkmann & Bose, Berlin, Germany, second edition, 1995. ISBN 3-922660-23-1. 240 pp. DM 80,00.

Tschichold:1995:NTHb

[Tsc95b] Jan Tschichold. The new typography: a handbook for modern designers, volume 8 of Weimar and now. University of California Press, Berkeley, CA, USA, 1995. ISBN 0-520-07146-8. xliv + 236 pp. LCCN Z116.T7513 1995. Translated from the original German editions [Tsc28, Tsc95a] by Ruari McLean. With an introduction by Robin Kinross.

Tschichold:1995:TAL

[Tsc95c] Jan Tschichold. Treasury of Alphabets and Lettering: a Source Book of the Best Letter Forms of Past and Present for Sign Painters, Graphic Artists, Commercial Artists, Typographers, Printers, Sculptors, Architects, and Schools of Art and Design. W. W. Norton & Co., New York, NY, USA, 1995. ISBN 0-393-70197-2. 236 pp. LCCN Z250 .T883 1995. US$35.00. This is identical to the 1966 edition, with a new introduction by Ben Rosen.

Tufte:1983:VDQ

[Tuf83] Edward R. Tufte. The Visual Display of Quantitative Information. Graphics Press, Box 430, Cheshire, CT 06410, USA, 1983. ISBN 0-9613921-0-X, 0-9613921-4-2. 197 pp. LCCN K27.S8 T84. US$34.00.

Tufte:1990:EI

[Tuf90] Edward R. Tufte. Envisioning Information. Graphics Press, Box 430, Cheshire, CT 06410, USA, 1990. 126 pp. LCCN QA90 .T82 1990. US$48.00.

Tufte:1997:VEI

[Tuf97] Edward R. Tufte. Visual Explanations: Images and Quantities, Evidence and Narrative. Graphics Press, Box 430, Cheshire, CT 06410, USA, 1997. ISBN 0-9613921-2-6. 156 pp. LCCN P93.5.T846 1997. US$45.00.

Mei:1980:LLC

[Tun80] Tung Yun Mei. LCCD, A Language for Chinese Character Design. Report STAN-CS-80-824, Stanford University, Department of Computer Science, Stanford, CA, USA, 1980. See also Software—Practice and ExperienceVol. 11 (1981), No. 12, pp. 1273–1292.

Tuthill:1983:TUS

[Tut83]  B. Tuthill.  Typeset-

ting on the UNIX system. Byte Magazine, 8(19):253–265, October 1983. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Tymms:1987:AI

[TW87] W. R. Tymms and M. D.

Wyatt. The Art of Illuminating. Chartwell Books, Inc., 110 Enterprise Avenue, Secaucus, New Jersey 07094, USA, 1987. ISBN 1-55521-198-4. 104 pp. LCCN ND3310 .T95 1987.

Utting:1983:TSC

[UCBB83] I. A. Utting, D. J.

Caul, H. Brown, and S. E. Binns. A typesetting server for the Cambridge ring. In Dallas and Spratt [DS83], pages 243–248. ISBN 0-444-86852-6. LCCN TK5105.5 .I3427 1983.

Udell:1989:PPI

[Ude89] J. Udell. The power of the press (interleaf’s TEchnical publishing software). Byte Magazine, 14(12):271–276, November 1989. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Uehara:1990:CTP

[Ueh90] T. Uehara. Current technology and problems in computer fonts. Journal of Information Processing, 31(11):1570–1580, 1990. CODEN JOSHA4. ISSN 0447-8053.

UKTEX:1987:U

[UKT ] UKTEX, 1987–. This is an electronic discussion list for TEX issues in the United Kingdom.

Ulichney:1987:DH

[Uli87] Robert Ulichney. Digital Halftoning. MIT Press, Cambridge, MA, USA, 1987. ISBN 0-262-21009-6. xiv + 362 pp. LCCN T385 .U451 1987.

Urban:1986:IL

[Urb86]  Michael Urban.  An

introduction to L^{A}TEX. Prepared for the TRW Software Productivity Project; reprinted with permission and distributed by TUG, 1986.

Urban:1991:PPL

[Urb91] Michael Urban (adaptation française de Éric Cornelis). Premiers pas en L^{A}TEX. Cahiers GUTenberg, May 1991. French translation of [Urb86].

Utting:1985:RDI

[Utt85] Ian Utting. The Rôle of device independent output. In Lucarella [Luc85b], pages 183– 191. ISBN 0-201-13399-7. LCCN Q224.E97 1985.

Vacha:1992:TTP

[Vác92] Ondrřej Vácha. TEX for typesetting in a publishing house. In Zlatuška [Zla92], pages 59–65. ISBN 80-210-0480-0.

ValienteFeruglio:1997:CTC

[Val97] Gabriel Valiente Feruglio. Composicion de textos cientificos con L^{A}TEX. Aula Teorica; 61. Universitat Politecnica de Catalunya, Barcelona, Spain, 1997. ISBN 84-8301-204-9. 259 pp. LCCN ???? US$22.78.

Wyk:1980:LTG

[Van80] C. J. Van Wyk. A language for typesetting graphics. Report STAN-CS-80-803, Stanford University, Department of Computer Science, Stanford, CA, USA, 1980.

VanWyde:1981:GTL

[Van81] C. T. Van Wyde. A graphics typesetting language. In ACM-SIGOA’81 [ACM81], pages 97–107. ISBN 0-89791-043-5. LCCN QA76.7 .S54 v.16:6.

vanHerwijnen:1990:PS

[van90] Eric van Herwijnen. Practical SGML. Kluwer Academic Publishers Group, Norwell, MA, USA, 1990. ISBN 0-7923-0635-X, 0-7923-9434-8. xviii + 307 pp. LCCN QA76.73.S44 V36 1990. UK£24.90, US$49.00.

VanLoon:1994:NDW

[Van94] A. J. Van Loon. No, I don’t want a numerical reference list! The EASE Journal: European Science Editing, (51):15–16, January 1994. ISSN 0950-1835.

vonBechtolsheim:1992:TP

[vB92a] Stephan von Bechtolsheim. TEX in Practice. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1992. ISBN 0-387-97296-X. 1587, with 31 illustrations pp. LCCN Z253.4.T47 B4 1993. US$169.00. Four volume set.

vonBechtolsheim:1992:TPB

[vB92b] Stephan von Bechtolsheim. TEX in Practice: Basics, volume 1. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1992. ISBN 0-387-97595-0. xl + 396 pp. LCCN Z253.4.T47 B4 1993. US$49.00. With a foreword by David F. Rogers.

vonBechtolsheim:1992:TPP

[vB92c] Stephan von Bechtolsheim. TEX in Practice: Paragraphs, Math, and Fonts, volume 2. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1992. ISBN 0-387-97596-9. xl + 368 pp. LCCN Z253.4.T47 B4 1993. US$49.00. With a foreword by David F. Rogers.

vonBechtolsheim:1992:TPT

[vB92d] Stephan von Bechtolsheim. TEX in Practice: Tokens, Macros, volume 3. Springer-Ver-lag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1992. ISBN 0-387-97597-7. xlvi + 656 pp. LCCN Z253.4.T47 B4 1993. US$49.00. With a foreword by David F. Rogers.

vonBechtolsheim:1993:TPO

[vB93a] Stephan von Bechtolsheim. TEX in Practice: Output Routines, volume 4. Springer-Ver-lag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1993. ISBN 0-387-97598-5. xxxvi + 422 pp. LCCN Z253.4.T47 B4 1993. US$49.00. With a foreword by David F. Rogers.

vonBechtolsheim:1993:TPP

[vB93b] Stephan von Bechtolsheim. TEX in Practice: Paragraphs, Math, and Fonts, volume 2. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1993. ISBN 0-387-97596-9, 3-540-97596-9. LCCN Z253.4.T47 B4 1993.

vonBechtolsheim:1993:TPT

[vB93c] Stephan von Bechtolsheim. TEX in Practice: Tokens, Macros, volume 3. Springer-Ver-lag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1993. ISBN 0-387-97597-7, 3-540-97597-7. LCCN Z253.4.T47 B4 1993.

Vanoirbeek:1992:EPE

[VC92] C. Vanoirbeek and G. Coray, editors. EP92—Proceedings of Electronic Publishing, ’92, International Conference on Electronic Publishing, Document Manipulation, and Typography, Swiss Federal Institute of Technology, Lausanne, Switzerland, April 7-10, 1992. Cambridge University Press, New York, 1992. ISBN 0-521-43277-4. LCCN Z 286 E43 I58 1992. UK£35. Prepared with L^{A}TEX, and Microsoft Word converted to L^{A}TEX.

vanDongen:2012:LFH

[vD12] Marc R. C. van Dongen. L^{A}TEX and Friends. X.media.publishing. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2012. ISBN 3-642-23815-7 (hardcover), 3-642-23816-5 (e-book). x + 330 pp. LCCN ????

vanderLaan:1992:FFI

[vdL92a] Kees van der Laan. FIFO and FIFO incognito. In Zlatuška [Zla92], pages 225–234. ISBN 80-210-0480-0.

vanderLaan:1992:TD

[vdL92b] Kees van der Laan. Table diversions. In Zlatuška [Zla92], page 211. ISBN 80-210-0480-0.

vanderLaan:1992:TCT

[vdL92c] Kees van der Laan. Typesetting crosswords via TEX. In Zlatuška [Zla92], pages 217–224. ISBN 80-210-0480-0.

vanderLaan:1989:SLM

[vdLCL89] C. G. van der Laan, D. C. Coleman, and J. R. Luyten. SGML–L^{A}TEX 1. Mathematical formulas. RC-Rapport 24, Rekencentrum der Rijksuniversiteit te Groningen, Groningen, The Netherlands, 1989. (English version).

vanderLaan:1989:LTL

[vdLL89] C. G. van der Laan and J. R. Luyten. L^{A}TEX templates: Letter and article. RC-Rapport 27, Rekencentrum der Rijksuniversiteit te Groningen, Groningen, The Netherlands, November 1989.

Veer:1988:HCI

[vdVM88] Gerrit C. van der Veer and Gijsbertus Mulder, editors. Humancomputer Interaction: Psychonomic Aspects. Conference of the Dutch Psychonomic Society. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1988. ISBN 0-387-18901-7.

Vens:1992:IPF

[Ven92] Erik-Jan Vens. Incorporating PostScript fonts in TEX. In Zlatuška [Zla92], pages 173–182. ISBN 80-210-0480-0.

Verhoeven:2000:DMP

[Ver00] Richard Verhoeven. The design of the Math∫ pad editor. Dissertation, Technische Universiteit Eindhoven, Eindhoven, The Netherlands, 2000. ii + 139 pp. Dissertation, Technische Universiteit Eindhoven, Eindhoven, 2000.

Vesilo:1992:TBO

[Ves92] R. A. Vesilo. Translating between ODA and multimedia DVI(TEX) documents. In IREECON ’91, Australia’s Electronics Convention Proceedings, pages 646–649 (vol. 2). IREE, Edgecliff, NSW, Australia, September 1992. ISBN 0-909394-26-1. LCCN ???? 2 vol.

vanHuu:1985:TIS

[vH85] Le van Huu. TEX and ISO/STPL standard. In Lucarella [Luc85b], pages 67–78. ISBN 0-201-13399-7. LCCN Q224.E97 1985.

vanHerwijnen:1989:UTI

[vH89] E. van Herwijnen. The use of text interchange standards for submitting physics articles to journals. In International Conference on Computing in High-Energy Physics, volume 57(1–

3) of Computer Physics Communications, pages 244–250. North-Holland Publishing Company, Amsterdam, The Netherlands, December 1989. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

Vinogradov:1992:RTN

[Vin92] Michael Vinogradov. Russian TEX: new eight bit fonts and IBM PC equipment. In Zlatuška [Zla92], pages 149–152. ISBN 80-210-0480-0.

Vipond:1996:PMA

[Vip96] Douglas Vipond. Problems with a monolithic APA style. American Psychologist, 51(6): 653, ???? 1996. CO-DEN AMPSAB. ISSN 0003-066X (print), 1935-990X (electronic). URL

http://psycnet.apa.
org/journals/amp/51/
6/653b.pdf;   https:
//www.researchgate.
net/publication/232601714_
Problems_with_a_monolithic_
APA_style.

Voss:2012:LQR

[Vos12] Herbert Voss. L^{A}TEX quick reference: a guide to L^{A}TEX commands, symbols, fonts, packages, and creating bibliographies. UIT Cambridge, Cambridge, UK, 2012. ISBN 1-906860-21-1 (paperback). 228 + 6 pp. LCCN ????

Voss:2004:PGT

[Voß04] Herbert Voß. PSTricks:

Grafik für TEX und L^{A}TEX. Lehmanns, Berlin, Germany, 2004. ISBN 3-86541-053-7. 486 pp. LCCN ???? 34.95 EUR.

Voss:2005:PGT

[Voß05] Herbert Voß. PSTricks:

Grafik für TEX und L^{A}TEX. Lehmanns, Berlin, Germany, second edition, 2005. ISBN 3-86541-073-1. xxiv + 486 pp. LCCN ???? 19.95 EUR. URL http:

//www.dante.de/help/
literatur/pstricks.
shtml.

Voss:2011:PGP

[Voß11a] Herbert Voß. PSTricks:

graphics and PostScript for TEX und L^{A}TEX. UIT Cambridge, Cambridge, UK, 2011. ISBN 1-906860-13-0. 912 pp. LCCN ????

Voss:2011:TML

[Voß11b] Herbert Voß. Typesetting mathematics with L^{A}TEX. UIT Cambridge, Cambridge CB5 8AF, UK, 2011. ISBN 1-906860-17-3 (paperback). vi + 282 pp. LCCN Z253.4.L38 V6713 2011.

                                        URL http://www.uit.
                                        co.uk/BK-TMWL/.

Voss:2011:TTL

[Voß11c] Herbert Voß. Typesetting tables with L^{A}TEX. UIT Cambridge, Cambridge CB5 8AF, UK, 2011. ISBN 1-906860-25-4. LCCN ????

                                        URL http://www.uit.
                                        co.uk/BK-TTWL/.

Voss:2012:PL

[Voß12]  Herbert Voß.  Pre-

sentations with L^{A}TEX: which package, which command, which syntax? Dante. Lehmanns Media, Berlin, Germany, 2012. ISBN 3-86541-496-6. vi + 205 pp. LCCN ????

Vulis:1992:MTA

[Vul92] Michael Vulis. Modern TEX and its Applications. CRC Press, 2000 N.W. Corporate Blvd., Boca Raton, FL 33431-9868, USA, 1992. ISBN 0-8493-4431-X. vii + 294 pp. LCCN Z253.4.T47V84 1993. US$32.95.

vanVliet:1986:TPD

[vV86] J. C. van Vliet, editor. Text Processing and Document Manipulation: Proceedings of the International Conference, University of Nottingham, 14–16 April 1986. Cambridge University Press, New York, April 1986. ISBN 0-521-32592-7. LCCN QA76.9.T48 T481 1986.

vanVliet:1988:DMT

[vV88] J. C. van Vliet, editor. Document Manipulation and Typography. Proceedings of the International Conference on Electronic Publishing, Document Manipulation and Typography, Nice (France), April 20–22, 1988. Cambridge University Press, New York, 1988. ISBN 0-521-36294-6. LCCN Z286.E43 I571 1988. US$47.50.

Waldock:1982:TAE

[Wal82a] W. C. Waldock. Typesetting in the academic environment. In User services, proceedings of the Tenth ACM SIGUCCS conference (Chicago, IL, Nov. 10–12, 1982), pages 25–28. ACM Press, New York, NY, USA, 1982.

Waller:1981:JTT

[Wal82b] R. H. W. Waller. Journal typography in transition. Journal of Research Communication Studies, 3(??):335–350, ???? 1981–1982. CO-DEN JRCSDF. ISSN 0378-5939.

Walden:1987:MFF

[Wal87a] Jeff Walden. More File Formats for Popular PC software: a Programmer’s Reference. Wiley, Chichester, UK, 1987. ISBN 0-471-85077-2. 369 pp. LCCN QA76.9.F5 W35 1987. US$24.95. From Computing Reviews: “The volume under review deals with (among others) Word-Perfect, SuperCalc 4, and the unusual Microsoft Rich Text Format (RTF). RTF has the distinction of being entirely readable ASCII and is reminiscent of TEX command format.” Conversion software from RTF to L^{A}TEX written by Daniel Taupin is available from major TEX archive sites.

Waldschmidt:1987:ECG

[Wal87b] H. Waldschmidt. An editor for constructing graphics with TEX. Computer Graphics Forum, 6(4):359–364, December 1987. CODEN CG-FODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Walsh:1994:MTW

[Wal94] Norman Walsh. Making TEX Work. O’Reilly &Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, 1994. ISBN 1-56592-051-1. xxxvi + 483 pp. LCCN TK5105.875.I57 M36 1994; Z253.4.T47 W34 1994. US$29.95.

URL ftp://ftp.ora.
com/pub/products/catalogs/
book.catalog;   http:
//www.oreilly.com/catalog/
9781565920514. Written

specifically for database maintainers and system administrators, this Unix-based technical guide covers installing, setting up, and running Internet applications such as gopher holes, FTP and telnet sites, mailing lists, WAIS, and World Wide Web pages, and other sites, as well as keeping them secure.

Wang:1996:TAE

[Wan96] Xinxin Wang. Tabular abstraction, editing, and formatting. Ph.D. dissertation, University of Waterloo, Waterloo, ON, Canada, 1996. 205

pp.  URL http://
search.proquest.com/
docview/304354192.

Witten:1982:PTD

[WBS82] I. H. Witten, M. Bonham, and E. Strong. On the power of traps and diversions in a document preparation language. Software—Practice and Experience, 12 (12):1119–1131, December 1982. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Werntges:1992:GIL

[Wer92] Heinz Werntges. Grafikimport in L^{A}TEX [English: Importing graphics in L^{A}TEX]. c’t, 12:252– 258, 1992. ISSN 0724-8679.

Wilberg:1997:L

[WF97] Hans Peter Wilberg and Friedrich Forssman. Lesetypographie. Verlang Hermann Schmidt, Mainz, Germany, 1997. ISBN 3-87439-375-5. 332 pp.

Wolcott:1976:CCT

[WH76] N. M. Wolcott and J. Hilsenrath. A contribution to computer typesetting techniques: Tables of coordinates for Hershey’s repertory of occidental type fonts and graphic symbols. Special Publication 424, U.S. National Bureau of Standards, Gaithersburg, MD, USA, April 1976.

Wezler:1986:PXI

[WH86] Albrecht Wezler and Ernst Hammerschmidt, editors. Proceedings of the XXXII International Congress for Asian and North African Studies, Hamburg, 25th– 30th August 1986, number IX in Zeitschrift der Deutschen Morgenländischen Gesellschaft, Supplement. Steiner, Stuttgart, Germany, 1986. ISBN 3-515-04808-1. LCCN DS1.5 .I565 1986.

White:1983:MG

[Whi83] Jan V. White. Mastering Graphics. R. R. Bowker Company, 1180 Avenue of the Americas, New York, NY 10036, USA, 1983. ISBN 0-8352-1704-3. xiii + 180 pp. LCCN Z253 .W47 1983.

White:1987:HST

[Whi87] Alex White. How to SPEC Type. Watson-Guptill Publications, 1515 Broadway, New York, NY 10036, USA, 1987. ISBN 0-8230-2473-3. 128 pp. LCCN Z253.W46 1987. US$14.95.

White:1988:GDE

[Whi88] Jan V. White. Graphic Design for the Electronic Age. Watson-Guptill Publications, 1515 Broadway, New York, NY 10036, USA, 1988. ISBN 0-8230-2121-1, 0-8230-2122-X (paperback). xi + 211 pp. LCCN Z286.D47 W5 1988.

White:1990:GP

[Whi90] Jan V. White. Great Pages. Serif Publishing, A subsidiary of Xerox Corporation, 701 South Aviation Boulevard, El Segundo, CA 90245, USA, 1990. ISBN 1-878567-01-2. 128 pp. LCCN Z286.D47 W5 1990.

Wikeckowski:1992:TPZ

[Wie92] Z. Wieckowski. TEX — polskie znaki inaczej. PCkurier, 15:89, lipiec 1992. In Polish.

Williamson:1983:MBD

[Wil83] Hugh Williamson. Methods of Book Design: The Practice of an Industrial Craft. Yale University Press, New Haven, CT, USA, third edition, 1983. ISBN 0-300-02663-3 (hardcover), 0-300-03035-5 (paperback). xiv + 391 pp. LCCN Z116.A3 W5 1983. UK£13. Bibliography, many examples and illustrations, index.

Williams:2004:BGA

[Wil04] George Williams. Beyond glyphs, advanced typographic features of fonts. In Syropoulos et al. [SBH^{+}04], pages 257–263. CO-DEN LNCSD9. ISBN 3-540-22801-2 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN Z253.3 I58 2004. URL http:

//link.springer-ny.
com/link/service/series/
0558/tocs/t3130.htm;
http://www.springerlink.
com/openurl.asp?genre=
issue&issn=0302-9743&
volume=3130;   http:
//www.springerlink.
com/openurl.asp?genre=
volume&id=doi:10.1007/
b99374.

Winkel:1988:TSU

[Win88] Brian J. Winkel. TEX:

The software used to produce this journal. Collegiate Microcomputer, 5(1):93–96, April/May 1988. CODEN CMICDL. ISSN 0731-4213.

Wise:2001:RR

[Wis01] Steve Wise. Revolution in references: Give readers a chance by putting page numbers. Social Studies of Science, 31(2):309–310, April 1, 2001. CODEN SSS-CDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https:

//journals.sagepub.
com/doi/pdf/10.1177/
0306312701031002009.

Witten:1985:ECT

[Wit85] Ian H. Witten. Elements of computer typography. International Journal of Man-Machine Studies, 23(6):623–687, December 1985. CO-DEN IJMMBC. ISSN 0020-7373. URL http:

//www.sciencedirect.
com/science/article/
pii/S0020737385800626.

Wittbecker:1986:GTB

[Wit86] Alan E. Wittbecker. Genesis of a TEX-based markup metalanguage. In Conference proceedings: setting the direction / ACM SIGUCCS User Services Conference XIV, Le Centre Sheraton, Montreal, Quebec, September 27th – October 1st, 1986, pages 31–33. ACM Press, New York, NY, USA, 1986. ISBN 0-89791-208-X. LCCN QA74.A15 1986. URL http:

//www.acm.org/pubs/
contents/proceedings/
userservices/324239/

. ACM order number 459860.

Weyerich:1993:LLLa

[WKS^{+}93a] Astrid Weyerich, Rainer Klötzer, Sabine Stromp, Kerstin Rieder, and Rainer Oesterreich. L^{A}TEX Lernen leicht gemacht— Eine aufgabenorientierte Anleitung—Teil A: Arbeitsheft. Technische Universität Berlin, Institut für Humanwissenschaft in Arbeit und Ausbildung, 10623 Berlin, Germany, 1993. ISBN 3-7983-1546-9. iv + 100 pp.

Weyerich:1993:LLLb

[WKS^{+}93b] Astrid Weyerich, Rainer Klötzer, Sabine Stromp, Kerstin Rieder, and Rainer Oesterreich. L^{A}TEX Lernen leicht gemacht— Eine aufgabenorientierte Anleitung—Teil B: Manual. Technische Universität Berlin, Institut für Humanwissenschaft in Arbeit und Ausbildung, 10623 Berlin, Germany, 1993. ISBN 3-7983-1546-9. 176 pp.

Wettl:2004:LKH

[WMS04] Ferenc Wettl, Gyula Mayer, and Péter Szabó. L^{A}TEX Kézikönyv. (Hungarian) [L^{A}TEX Handbook]. Könnyen is lehet. Panem, Budapest, Öv u. 146. 1146-Hungary, 2004. ISBN 963-545-398-1. ???? pp.

Wohlfeil:1998:PCB

[Woh98]  Stefan Wohlfeil.  On

the Pagination of Complex, Book-Like Documents. Shaker Verlag, Aachen and Maastricht, The Netherlands, 1998. ISBN 3-8265-3304-6. 224 pp. LCCN ???? DM 98.00. URL http://

www.shaker.de/Online-
Gesamtkatalog/Details.
idc?ID=24201&CC=59311&
IDSRC=1&ISBN=3-8265-
3304-6&Reihe=15;   http:
//wwwpi6.fernuni-hagen.
de/Forschung/Pagination/

.

Wonneberger:1986:VVT

[Won86a] R. Wonneberger. “verheissung und versprechen”, a third generation approach to theological typesetting. In Désarménien [Dés86], pages 180– 198. ISBN 3-540-16807-9, 0-387-16807-9 (paperback). LCCN Z253.4.T47 T49 1986.

Wonneberger:1986:TPT

[Won86b] Reinhard Wonneberger. TEX for philological typesetting. In Wezler and Hammerschmidt [WH86], pages 595– 596. ISBN 3-515-04808-1. LCCN DS1.5 .I565 1986.

Wonneberger:1987:L

[Won87] Reinhard Wonneberger. L^{A}TEX. Addison-Wesley Kompaktführer. Addison-Wesley Verlag, Bonn, Germany, 1987. ISBN 3-925118-46-2. xiii + 141 pp.

Wonneberger:1988:L

[Won88] Reinhard Wonneberger.

L^{A}TEX. Addison-Wesley Kompaktführer. Addison-Wesley Verlag, Bonn, Germany, second edition, 1988. ISBN 3-89319-152-6. xiv + 141 pp. Reprinted in 1990 and 1991.

Wonneberger:1990:SDP

[Won90] Reinhard Wonneberger.

Structured document processing: the L^{A}TEX approach. In Nadrchal [Nad90], pages 177–189. CODEN CPHCBZ. Invited paper.

Wonneberger:1993:L

[Won93] Reinhard Wonneberger.

L^{A}TEX. Addison-Wesley Kompaktführer. Addison-Wesley Verlag, Bonn, Germany, third edition, April 1993. ISBN 3-89319-589-0. xvi + 166 pp. DM29.80.

Wood:1987:TWG

[Woo87] Patrick Wood. Typesetting: when gray isn’t. The PostScript Language Journal, 1(2):15– 19, June 1987. ISSN 0891-5873.

Wu:2008:ITX

[Wu08] Qingyu Wu. Image of transfer from GL(2) × GL(3) to GL(6) done with L^{A}TEX. PhD thesis, Purdue University, West Lafayette, IN, USA, 2008. 87

pp.  URL http://
search.proquest.com/
docview/288334139. The-

sis (Ph.D.)–Purdue University.

Wyrostek:1992:LSP

[Wyr92] Piotr Wyrostek, translator. L^{A}TEX: System przygotowywania dokumentów. Przewodnik użytkownika i podrecznik. Ariel Sp. z o.o, Kraków, Poland, 1992. ISBN 83-900460-1-6. 91000 zl. Polish translation of [?].

Yang:2004:EME

[YF04] Michael Yang and Richard Fateman. Extracting mathematical expressions from PostScript documents. In Gutierrez [Gut04], pages 305–311. ISBN 1-58113-827-X. LCCN ???? URL http:

//www.ocf.berkeley.
edu/~mlyang/papers/
MichaelYangPsmath.pdf.

Yamakawa:1990:DPV

[YKT90] T. Yamakawa, Y. Kawabata, and H. Tamura. Desktop publishing viewed from the office automation industries. Journal of Information Processing, 31(11):1508–1517, 1990. CODEN JOSHA4. ISSN 0447-8053.

Youngen:1997:TMM

[You97] Ralph Youngen. Toward a mathematical markup language. Notices of the American Mathematical Society, 44 (9):1107–1109, October 1997. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic). URL http://

www.ams.org/notices/
199709/comm-youngen.
pdf.

Yuan:2001:LPS

[Yua01]  Michael Yuan.  La-

TeX2HTML: Publish science to the Web. Linux Journal, 92:93–94, December 2001. CODEN LIJOFX. ISSN 1075-3583 (print), 1938-3827 (electronic).

Zagretdinov:1994:ISL

[Z^{+}94] Renat Zagretdinov et al.

Izdatel’skaiia systema L^{A}TEX. Kratkoe rukovodstvo [English: The L^{A}TEX editing system. Brief guide]. Kazan State University, Kazan, Russia, 1994. 94 pp. This manual is a short introduction to L^{A}TEX 2.09, with some material added about the Russian version CyrTUG-emTEX; it also contains a glossary of L^{A}TEX control sequences, etc. First press run: 1000 copies.

ZabalaSalelles:1982:IGO

[Zab82] Ignacio Andres Zabala Salelles. Interfacing with graphics objects. PhD thesis, Department of Computer Science, Stanford University, Stanford, CA, USA, December 1982. 146 pp.

                                        URL http://wwwlib.
                                        umi.com/dissertations/
                                        fullcit/8314505.

Zapf:1987:HZH

[Zap87] Hermann Zapf. Hermann Zapf and His Design Philosophy: Selected Articles and Lectures on Calligraphy and Contemporary Developments in Type Design, with Illustrations and Bibliographical Notes, and a Complete List of his Typefaces. Society of Typographic Arts, Chicago, IL, USA, 1987. ISBN 0-941447-00-6. 254 pp. LCCN Z250 .Z323 1987. US$45.00.

Zapf:2007:ASC

[Zap07a] Hermann Zapf. Alphabet Stories: a Chronicle of Technical Developments. Mergenthaler Edition, Linotype GmbH, Bad Homburg, Germany, 2007. ISBN 3-9810319-6-2. 149 pp. LCCN ????

Zapf:2007:ACT

[Zap07b]  Hermann Zapf.  Al-

phabetgeschichten: Eine Chronik technischer Entwicklungen. Mergenthaler Edition, Linotype GmbH, Bad Homburg, Germany, 2007. ISBN 3-9810319-5-4. 156 pp. LCCN ???? 29.00 EUR.

Zlatuska:1992:EPE

[Zla92] Jiří Zlatuška, editor. EuroTEX ’92: Proceedings of the 7th European TEX Conference, Prague, Czechoslovakia, September 14–18, 1992, Proceedings of the European TEX Conference. Masarykova Universita, Brno, Czechoslovakia, September 1992. ISBN 80-210-0480-0.